### v1

# The Cohen–Lenstra moments over function fields via the stable homology of non-splitting Hurwitz spaces

Abstract.
We compute the average number of surjections from class groups of quadratic function fields over $\mathbb{F}_{q}(t)$ onto finite odd order groups $H$, once $q$ is sufficiently large. These yield the first known moments of these class groups, as predicted by the Cohen–Lenstra heuristics, apart from the case $H=\mathbb{Z}/3\mathbb{Z}$. The key input to this result is a topological one, where we compute the stable rational homology groups of Hurwitz spaces associated to non-splitting conjugacy classes.
Key words and phrases:
The Cohen–Lenstra heuristics, Hurwitz spaces, homological stability, higher algebra
2020 Mathematics Subject Classification:
Primary 11R29; Secondary 11R11, 11R58, 55P43
1. Introduction
The Cohen–Lenstra heuristics, introduced by Cohen and Lenstra in [CL84], predict the distribution of the odd part of class groups of quadratic number fields, and have been one of the driving conjectures in arithmetic statistics over the last four decades. If $K$ is a quadratic extension of $\mathbb{Q}$, and $\mathscr{O}_{K}$ is its ring of integers, the Cohen–Lenstra heuristics predict the distribution of the odd part of $\operatorname{Cl}(\mathscr{O}_{K})$. Next, let $\mathbb{F}_{q}$ be a finite field with odd characteristic, $K$ be a quadratic extension of $\mathbb{F}_{q}(t)$, and $\mathscr{O}_{K}$ denote the normalization of $\mathbb{F}_{q}[t]$ in $K$. The natural extension of the Cohen–Lenstra heuristics to global function fields similarly predicts the distribution of the odd part of $\operatorname{Cl}(\mathscr{O}_{K})$.
One of the primary approaches to determining this distribution is to compute its moments, by which we mean the average number of surjections $\left|\operatorname{Surj}(\operatorname{Cl}(\mathscr{O}_{K}),H)\right|$, for $H$ a finite abelian group of odd order, as $K$ ranges over imaginary quadratic fields or real quadratic fields. As further motivation for considering these moments, it was shown in [WW21, Theorem 1.3] that the Cohen–Lenstra distribution for imaginary quadratic fields or real quadratic fields is determined by its associated moments. Hence, if one were able to compute all conjectured moments, one would prove the Cohen–Lenstra heuristics.
Although the problem of computing the average size of the $\ell$-torsion (corresponding to the $\mathbb{Z}/\ell\mathbb{Z}$ moment) in the class groups of quadratic fields may appear substantially more tractable than determining the entire distribution of these class groups, the only $\ell$ for which this has been carried out is $\ell=3$. This was computed over $\mathbb{Q}$ by Davenport and Heilbronn [DH71] and the analog over function fields has been verified by Datskovsky and Wright in [DW88]. To the best of our knowledge, since 1988, no additional odd order moments of class groups of quadratic fields have been computed. We note that the original Cohen–Lenstra conjectures have been extended from odd order abelian groups to all finite abelian groups by Gerth, and the limiting distribution associated to $2^{\infty}$-part of the class group in this sense has been computed by Smith [Smi22, Theorem 1.9].
Over function fields of the form $\mathbb{F}_{q}(t)$, significant progress was made when a weaker version of these moments was computed. Namely, in [Ach08, Theorem 3.1], building on [Ach06], Achter computes these moments where one first takes a large $q$ limit. Using techniques from homological stability, Ellenberg–Venkatesh–Westerland were able to strengthen Achter’s results by computing a version of these moments where one first takes a large height limit, and only after takes a large $q$ limit [EVW16]. But even via this approach, no moment (other than the case $H=\mathbb{Z}/3\mathbb{Z}$) over a fixed $\mathbb{F}_{q}(t)$ has been computed, or even shown to exist.
In this paper, we compute the moment associated to any odd order abelian group associated to class groups of quadratic extensions of $\mathbb{F}_{q}(t)$ over suitably large finite fields $\mathbb{F}_{q}$ of suitable characteristic. Our main new input which lets us accomplish this is a computation of the stable rational homology of certain Hurwitz spaces.
1.1. Results toward the Cohen–Lenstra heuristics
Let $q$ be an odd prime power and let $\mathcal{MH}_{n,q}$ denote the set of function fields $K$ of monic smooth hyperelliptic curves of the form $y^{2}=f(x)$, where $f(x)$ is a monic squarefree degree $n$ polynomial with coefficients in $\mathbb{F}_{q}$. (We consider monic smooth hyperelliptic curves to be the analog of quadratic number fields in the function field case.) Let $\mathscr{O}_{K}$ denote the normalization of $\mathbb{F}_{q}[t]$ in the quadratic extension $K$ and let $\operatorname{Cl}(\mathscr{O}_{K})$ denote its class group.
Theorem 1.1.1.
Suppose $H$ is a finite abelian group of odd order. Let $q$ be an odd prime power with $\gcd(|H|,q(q-1))=1$. There is an integer $C$, depending only on $H$, so that if $q>C$ and $i\in\{0,1\}$,
(1.1) $\lim_{\begin{subarray}{c}n\to\infty\\ n\equiv i\bmod 2\end{subarray}}\frac{\sum_{K\in\mathcal{MH}_{n,q}}\left|\operatorname{Surj}(\operatorname{Cl}(\mathscr{O}_{K}),H)\right|}{\sum_{K\in\mathcal{MH}_{n,q}}1}=\begin{cases}1&\text{ if }i=1\\ \frac{1}{|H|}&\text{ if }i=0.\end{cases}$
Theorem 1.1.1 is the special case of Theorem 1.2.1, stated below, where we take the number $h$, defined there, to be $1$.
Remark 1.1.1.
Even though for any given $H$, we compute the $H$-moment of the class group of quadratic fields over $\mathbb{F}_{q}(t)$ for sufficiently large $q$, we do not prove the Cohen–Lenstra heuristics. Although it is true that knowledge of all moments do determine the Cohen–Lenstra distribution, for a fixed value of $q$, we are only able to compute moments associated to sufficiently small groups $H$ relative to $q$. We do so by computing the stable homology of related Hurwitz spaces. If one were also able to obtain a better understanding of the unstable homology of these Hurwitz spaces, together with the eigenvalues of the Frobenius action, one might be able to verify the entire predictions of the Cohen–Lenstra heuristics over $\mathbb{F}_{q}(t)$.
Remark 1.1.1.
It is possible to make the constant $C$ in Theorem 1.1.1 explicit: see the discussion in § 5.3.
Remark 1.1.1.
Since it was geometrically more natural, we took a limit in (1.1) of the $H$-moments of the class group as $K\in\mathcal{MH}_{n,q}$, where $K$ has fixed log discriminant degree $n$. However, from this we can easily also deduce
(1.2) $\lim_{\begin{subarray}{c}n\to\infty\\ n\equiv i\bmod 2\end{subarray}}\frac{\sum_{K\in\cup_{j\leq n,j\equiv n\bmod 2}\mathcal{MH}_{j,q}}\left|\operatorname{Surj}(\operatorname{Cl}(\mathscr{O}_{K}),H)\right|}{\sum_{K\in\cup_{j\leq n,j\equiv n\bmod 2}\mathcal{MH}_{j,q}}1}=\begin{cases}1&\text{ if }i=1\\ \frac{1}{|H|}&\text{ if }i=0\end{cases}$
by summing the terms in (1.1) over all integers up to $n$ with the same parity as $n$.
Remark 1.1.1.
One may wonder whether the analog of Theorem 1.1.1 would hold if one averaged over all smooth hyperelliptic curves instead of just monic ones. We explain in § 2.1 why the resulting averages will be the same in the case $n$ is odd if one takes non-monic smooth hyperelliptic curves ramified over $\infty$ in place of monic ones ramified over $\infty$. In the case that $n$ is even, the statistics of non-monic smooth hyperelliptic curves seem somewhat more subtle. We believe it would be quite interesting and doable to work this case out in detail.
1.2. Results toward the Cohen–Lenstra heuristics with roots of unity
We note that Theorem 1.1.1 assumes $\gcd(q-1,|H|)=1$, so that there are no roots of unity in the base field of order dividing $|H|$. There has been a significant amount of work toward stating and proving the correct version of the Cohen–Lenstra heuristics in the presence of such roots of unity, see [Mal08, Mal10, AM15, Gar15, LST20, SW23]. To this end, we prove the following generalization, which removes the condition that $\gcd(q-1,|H|)=1$. For $G$ a finite abelian group and $h\in\mathbb{Z}_{>0}$, we use $G[h]$ to denote the $h$-torsion subgroup of $G$.
Theorem 1.2.1.
Suppose $H$ is a finite abelian group of odd order. Fix $q$ an odd prime power with $\gcd(|H|,q)=1$ and let $h:=\gcd(|H|,q-1)$. There is an integer $C$ depending only on $H$ so that, if $q>C$ and $i\in\{0,1\}$,
(1.3) $\lim_{\begin{subarray}{c}n\to\infty\\ n\equiv i\bmod 2\end{subarray}}\frac{\sum_{K\in\mathcal{MH}_{n,q}}\left|\operatorname{Surj}(\operatorname{Cl}(\mathscr{O}_{K}),H)\right|}{\sum_{K\in\mathcal{MH}_{n,q}}1}=\begin{cases}|\wedge^{2}H[h]|&\text{ if }i=1\\ \frac{|\wedge^{2}H[h]|}{|H|}&\text{ if }i=0.\end{cases}$
Theorem 1.2.1 is an immediate consequence of Theorem 5.4.1, which proves a refined statement about the rate of convergence of the limit.
1.3. Results toward the non-abelian Cohen–Lenstra heuristics
Having focused on the abelian Cohen–Lenstra heuristics above, in recent years, there has been significant progress toward generalizing these heuristics to the non-abelian case. This involves replacing the class group (the maximal abelian unramified extension) by the maximal unramified extension.
For quadratic extensions, these conjectures for pro-$p$ groups were formulated in [BBH17] and [BBH21], and function field versions for pro-odd groups were proven in a suitable large $q$ limit in [BW17] when there are no additional roots of unity in the base field. This conjecture was further generalized to certain even order groups and the case when there are roots of unity in the base field (with an additional data of a lifting invariant to keep track of the the fact that the moments may no longer equal $1$) in [Woo19]. We will next prove many cases of [Woo19, Conjecture 5.1], the nonabelian Cohen–Lenstra heuristics when there are roots of unity in the base field, keeping track of the additional data of a lifting invariant.
We note that the above statements were all about quadratic extensions, and they have since been generalized to conjectures about the maximal unramified extension of a $\Gamma$ extension, for $\Gamma$ an arbitrary finite group. Namely, this generalization was made in [LWZB24] for the case that there are few roots of unity in the base field and it was made in [Liu22] when there are additional roots of unity in the base field, where the author also keeps track of the additional data of a lifting invariant.
To state our results proving many cases of [Woo19, Conjecture 5.1], we will introduce some notation. These cases we prove are also cases of [Liu22, Conjecture 1.2]. Because the relevant notation in [Liu22] appears to be a bit simpler for the cases we will prove, we follow the notation there.
Notation 1.3.0.
Let $K$ be a quadratic extension of $\mathbb{F}_{q}(t)$, let $K^{\sharp}_{\emptyset}/K$ denote the maximal unramified extension of $K$ that is split completely over $\infty$, in the sense of § 2.3, and has degree prime to $2q$. Define $G^{\sharp}_{\emptyset}(K):=\operatorname{Gal}(K^{\sharp}_{\emptyset}/K)$. Then, $G^{\sharp}_{\emptyset}(K)$ has an action of $\mathbb{Z}/2\mathbb{Z}$ coming from the $\mathbb{Z}/2\mathbb{Z}=\operatorname{Gal}(K/\mathbb{F}_{q}(t))$ action on $K$. This action is well-defined using the Schur–Zassenhaus lemma; for further explanation, see the paragraph prior to [LWZB24, Definition 2.1]. For $H$ a group with a $\mathbb{Z}/2\mathbb{Z}$ action, we use $\operatorname{Surj}_{\mathbb{Z}/2\mathbb{Z}}(G^{\sharp}_{\emptyset}(K),H)$ to denote the set of $\mathbb{Z}/2\mathbb{Z}$ equivariant surjections $G^{\sharp}_{\emptyset}(K)\to H$. We use ^$\hat{\mathbb{Z}}(1)_{(2q)^{\prime}}$ to denote the pro-prime to $2q$ completion of for ^$\widehat{\mathbb{Z}}(1):=\lim_{n}\mu_{n}(\overline{\mathbb{F}}_{q})$. For $L/K$ an extension of fields, there is a certain map $\omega_{L/K}:\hat{\mathbb{Z}}(1)_{(2q)^{\prime}}\to H_{2}(\operatorname{Gal}(L/\mathbb{F}_{q}(t)),\mathbb{Z})_{(2q)^{\prime}}$ defined in [Liu22, Definition 2.13]. For $\pi\in\operatorname{Surj}_{\mathbb{Z}/2\mathbb{Z}}(G^{\sharp}_{\emptyset}(K),H)$, we use $\pi_{*}:H_{2}(\operatorname{Gal}(K^{\sharp}_{\emptyset}/\mathbb{F}_{q}(t)),\mathbb{Z})_{(2q)^{\prime}}\to H_{2}(H\rtimes\mathbb{Z}/2\mathbb{Z},\mathbb{Z})_{(2q)^{\prime}}$ to denote the corresponding map induced by $\pi$. We say that $H$ is admissible for the action of $\mathbb{Z}/2\mathbb{Z}$ if $H$ has odd order and is generated by elements of the form $h^{-1}\cdot\iota(h)$ for $h\in H$ and $\iota$ the nontrivial element of $\mathbb{Z}/2\mathbb{Z}$.
Theorem 1.3.1.
Using notation from § 1.3, suppose $H$ is an admissible group with a $\mathbb{Z}/2\mathbb{Z}$ action. Fix $q$ an odd prime power with $\gcd(|H|,q)=1$. Let $\delta:\hat{\mathbb{Z}}(1)_{(2q)^{\prime}}\to H_{2}(H\rtimes\mathbb{Z}/2\mathbb{Z},\mathbb{Z})_{(2q)^{\prime}}$ be a group homomorphism with $\operatorname{ord}(\operatorname{im}\delta)\mid q-1.$ There is an integer $C$, depending only on $H$ so that, if $q>C$,
(1.4) $\lim_{\begin{subarray}{c}n\to\infty\\ n\equiv 0\bmod 2\end{subarray}}\frac{\sum_{K\in\mathcal{MH}_{n,q}}\left|\{\pi\in\operatorname{Surj}_{\mathbb{Z}/2\mathbb{Z}}\left(G^{\sharp}_{\emptyset}(K),H\right):\pi_{*}\circ\omega_{K^{\sharp}/K}=\delta\}\right|}{\sum_{K\in\mathcal{MH}_{n,q}}1}=\frac{1}{[H:H^{\mathbb{Z}/2\mathbb{Z}}]},$
where $[H:H^{\mathbb{Z}/2\mathbb{Z}}]$ denotes the index of the $\mathbb{Z}/2\mathbb{Z}$ invariants of $H$ in $H$.
Theorem 1.3.1 is an immediate consequence of Theorem 5.5.1, which proves a refined statement about the rate of convergence of the limit.
1.3.2.
In order to prove Theorem 1.1.1, Theorem 1.2.1, and Theorem 1.3.1, we adapt the general strategy outlined in [EVW12]. [^1] Before Ellenberg, Venkatesh, and Westerland published their preprint [EVW12], Oscar Randal-Williams discovered a serious error in [EVW12], which the authors were unable to address. See [ell] for further details. The main result of this paper can be viewed as a resolution of that error. Namely, one constructs certain Hurwitz spaces parameterizing covers of $\mathbb{P}^{1}$ branched at $n$ points over $\mathbb{A}^{1}$. Via the Grothendieck–Lefschetz trace formula, computing the above moments amounts to obtaining a precise count of the number of $\mathbb{F}_{q}$ points of these Hurwitz spaces. It was shown in [EVW16] that the homology groups of these spaces stabilize in a linear range. The authors in loc. cit. used this stability to prove a weaker version of our results, which included a large $q$ limit. In order to compute the desired moments, it suffices to compute the stable value of these homology groups. Therefore, after the reductions mentioned above, the key new result of this paper is to compute the stable homology groups of certain Hurwitz spaces, which we discuss next.
1.4. Stable homology of Hurwitz spaces
There are a number of situations in algebraic geometry where one may be interested in determining the stable homology or cohomology groups of a sequence of spaces. One natural example is the sequence of spaces $\{\mathscr{M}_{g}\}_{g\geq 2}$. In [Har85], Harer proved that the $i$th cohomology of $\mathscr{M}_{g}$ stabilizes as $g$ grows, see also [Wah08] and [Wah13]. Later, Madsen and Weiss [MW07] computed the stable values of these cohomology groups.
As a companion to the moduli spaces of curves, Hurwitz stacks also form prominent objects of study in algebraic geometry. There are a number of equivalent viewpoints on how to think about Hurwitz stacks. If one fixes a finite group $G$, a conjugacy class $c$, and an integer $n$, the associated Hurwitz stack parameterizes $G$ covers of $\mathbb{A}^{1}$ branched at a degree $n$ divisor with inertia at each point of the divisor lying in $c$.
It is natural to ask whether the $i$th homology of these Hurwitz stacks also stabilizes as $n$ grows. Ellenberg, Venkatesh, and Westerland [EVW16] showed that this is indeed the case for certain non-splitting $(G,c)$, namely those $(G,c)$ so that $c\cap G^{\prime}$ consists of at most one conjugacy class for every subgroup $G^{\prime}\subset G$. An important example of non-splitting $(G,c)$ is the following, which is relevant for the Cohen–Lenstra heuristics: $G=H\rtimes\mathbb{Z}/2\mathbb{Z}$, for $H$ an odd order abelian group, and $c$ the conjugacy class of order $2$ elements. However, the value of these stable homology groups remained open. In this paper, we will compute this stable value of the homology of Hurwitz spaces associated to non-splitting $(G,c)$.
We will consider a slight variant of the above notion of Hurwitz stack, which we refer to as the pointed Hurwitz space $\operatorname{Hur}^{G,c}_{n},$ which we define precisely in § 2.1 and § 2.1. This variant is a finite étale cover of the Hurwitz stack described above, and involves marking a point over $\infty$. We briefly mention four equivalent descriptions of this space. The equivalence between the descriptions given below in § 1.4.1, § 1.4.2, and § 1.4.4 is shown in [EVW16, §2.3], while the equivalence between these and the description in § 1.4.3 is given in [RW20, Remark 6.4].
1.4.1. In topology as covers of the disc
One description of our pointed Hurwitz spaces, $\operatorname{Hur}^{G,c}_{n},$ is that they parameterize pointed branched covers of the disc, branched at $n$ points, where the inertia type of each branch point lies in a conjugacy class $c\subset G$.
1.4.2. In homotopy theory as an orbit space
The pointed Hurwitz space $\operatorname{Hur}^{G,c}_{n}$ can also be described as the homotopy quotient $(c^{n})_{hB_{n}}$, where $B_{n}$ is the braid group on $n$ strands.
1.4.3. In homotopy theory as a free $\mathbb{E}_{2}$ algebra
Conjugation by $G$ gives a $G$-action on $\operatorname{Hur}_{n}^{G,c}$, and $\coprod_{n\geq 0}\operatorname{Hur}_{n}^{G,c}$ can also be described as the free $\mathbb{E}_{2}$-algebra generated by $c$ in the category of $G$-crossed spaces, see [RW20, Remark 6.4].
1.4.4. In algebraic geometry as a moduli space
Algebro-geometrically, we can also think of complex valued points of $\operatorname{Hur}^{G,c}_{n}$ as parameterizing $G$-covers $X\to\mathbb{P}^{1}_{\mathbb{C}}$, branched at a degree $n$ divisor in $\mathbb{A}^{1}_{\mathbb{C}}$ with inertia at these branch points lying in $c\subset G$ and with a marked point over $\infty\in\mathbb{P}^{1}_{\mathbb{C}}$. [^2] When the monodromy at $\infty$ is nontrivial of order $r$, we need to replace $\mathbb{P}^{1}_{\mathbb{C}}$ with a root stack of order $r$ at $\infty$ to make this technically correct. See § 2.1 for a formal algebro-geometric definition of these pointed Hurwitz spaces. We call them pointed because they involve making a choice of marked point of the cover $X$ over $\infty$.
Notation 1.4.4.
Let $\operatorname{Conf}_{n}$ denote the configuration space of $n$ unordered, distinct points in $\mathbb{A}^{1}_{\mathbb{C}}$. There is a natural map $\operatorname{Hur}^{G,c}_{n}\to\operatorname{Conf}_{n}$ which sends a branched cover $X\to\mathbb{P}^{1}$ to the intersection of its branch locus with $\mathbb{A}^{1}$.
As setup for our main result computing the stable homology of Hurwitz spaces, following [EVW16, Definition 3.1], we recall the notion of a non-splitting conjugacy class.
Definition 1.4.4.
Let $G$ be a finite group and $c\subset G$ be a conjugacy class. We say $(G,c)$ is non-splitting if $c$ generates $G$ and, for every subgroup $G^{\prime}\subset G$, $c\cap G$ is either empty or consists of a single conjugacy class.
That is, $(G,c)$ is non-splitting if $c$ does not split into multiple conjugacy classes upon intersecting with subgroups. See § 5.1 for some examples relevant to the Cohen–Lenstra heuristics. Our main result on the stable homology of Hurwitz spaces is the following.
Theorem 1.4.5.
Let $G$ be a finite group, $c\subset G$ a conjugacy class, and suppose $(G,c)$ is non-splitting. There are constants $I$ and $J$ depending only on $G$ so that for any $i\geq 0$ and $n>iI+J$ and any connected component $Z\subset\operatorname{Hur}^{G,c}_{n,\mathbb{C}}$, the map $H_{i}(Z;\mathbb{Q})\to H_{i}(\operatorname{Conf}_{n};\mathbb{Q})$ is an isomorphism.
We prove Theorem 1.4.5 in § 4.2.2.
Remark 1.4.5.
Theorem 1.4.5 is known to hold for certain special components $Z$ by work of Bianchi and Miller [BM23, Corollary C’] and the text following it. Via personal communication, we learned this argument was also known to Ellenberg–Venkatesh–Westerland. In particular, this applies when $G$ is the dihedral group of order $2n$, for $n$ odd, to components $Z$ whose boundary monodromy (see § 2.2) is a generator of the index $2$ abelian subgroup of $G$. However, it does not apply, for example, when $G$ is the dihedral group of order $2n$ and the boundary monodromy of $Z$ is trivial. See also [Tie16, §6] for additional computations of the homology of Hurwitz spaces in other special cases similar to the above, via essentially the same strategy.
Additionally, we mention that the results of [BM23, Corollary C’] do not require a non-splitting hypothesis on $(G,c)$. As a trade off, without the non-splitting hypothesis, they obtain bounds on the stability range which are much worse than linear in the homological degree; see also [BM23, §1.2].
Remark 1.4.5.
The subtlety in understanding homological stability for $\operatorname{Hur}_{n}^{G,c}$ (which includes points corresponding to disconnected covers) is that it does not stabilize with respect to the operator $V\in H_{0}(\operatorname{Hur}^{G,c}_{\operatorname{ord}(c)\cdot|c|,\mathbb{C}};\mathbb{Q})$ corresponding to $\prod_{g\in c}[g]^{\operatorname{ord}(g)}$. Rather, [EVW16] show that the operator $\sum_{g\in c}[g]^{\operatorname{ord}(g)}$ stabilizes the homology, which is not an operation that exists at the level of spaces. Oscar Randal-Williams mentions in his Bourbaki survey article that “homological stability of the spaces $\mathrm{CHur}_{n}^{G,c}$ with respect to the maps $V$ should serve as a guiding problem for mathematicians working in this subject.” [RW20, p. 24].
On the subspace $\mathrm{CHur}_{n}^{G,c}\subset\operatorname{Hur}_{n,\mathbb{C}}^{G,c}$ corresponding to $G$-covers that are connected, it follows from Theorem 1.4.5 that for $(G,c)$ non-splitting, $V$ does stabilize the homology. Moreover, we show the stable value on each component agrees with that of $\operatorname{Conf}_{n}$.
1.5. The stable homology for $S_{3}$ and degree $3$ covers
Before proceeding further, we pause to highlight a seemingly elementary case of Theorem 1.4.5 which, surprisingly, was previously unknown. Consider the special case that $G=S_{3}$, and the conjugacy class $c\subset G$ consists of transpositions. The pointed Hurwitz space $\operatorname{CHur}^{G,c}_{n}$ associated to this parameterizes connected degree $3$ simply branched covers of $\mathbb{P}^{1}$, branched over a degree $n$ divisor in $\mathbb{A}^{1}$, see § 2.1 for a more formal definition. Simply branched triple covers have been studied extensively in algebraic geometry, and it is natural to ask about the homology of the space of such covers as the number of branch points grows.
Corollary 1.5.0.
There are constants $I$ and $J$ so that for any $i\geq 0$ and $n>Ii+J$
$h^{i}([\operatorname{CHur}^{S_{3},\{(12),(13),(23)\}}_{n}/S_{3}];\mathbb{Q})=\begin{cases}1&\text{ if }i\in\{0,1\}\\ 0&\text{ otherwise. }\\ \end{cases}$
Remark 1.5.0.
If one considers the space of all smooth trigonal curves, instead of just simply branched ones, the stable cohomology of such spaces has been computed in [Zhe24] using Vassiliev’s method. See [Tom05], [Gor05], and [Vas99] for references on Vassiliev’s method. However, via personal communication, we learned that it is unclear whether a similar method can be used to compute the stable cohomology of the space of simply branched trigonal curves.
Remark 1.5.0.
There is a cycle class map from the Chow ring of these Hurwitz spaces to the cohomology of these Hurwitz spaces. The same explicit parameterizations which allowed Davenport and Heilbronn [DH71] to compute the number of cubic fields allowed Patel and Vakil to compute the stable Chow rings of these Hurwitz spaces [PV15, Theorem C], see also [CL22].
1.6. Conjectures on stable homology of Hurwitz spaces
Before proceeding to describe the new ideas in the proof of Theorem 1.4.5, we give a conjecture for what we think the stable homology of Hurwitz spaces should look like in general.
Notation 1.6.0.
To set up notation, let $G$ be a finite group and $c\subset G$ denote a union of conjugacy classes. That is, $c=c_{1}\cup\cdots c_{k}$, where each $c_{i}\subset G$ is a conjugacy class. Let $\operatorname{CHur}^{G,c_{1},\ldots,c_{k}}_{n_{1},\ldots,n_{k}}$ denote the Hurwitz space parameterizing connected $G$ covers of $\mathbb{A}^{1}_{\mathbb{C}}$ with a trivialization at $\infty$ and $n_{i}$ branch points over $\mathbb{A}^{1}_{\mathbb{C}}$ with inertia in the conjugacy class of any element of $c_{i}$. We also let $\operatorname{Conf}_{n_{1},\ldots,n_{k}}$ denote the multi-colored configuration space in $\mathbb{A}^{1}_{\mathbb{C}}$ parameterizing $n_{i}$ points of color $i$ so that all $n_{1}+\cdots+n_{k}$ points in $\mathbb{A}^{1}_{\mathbb{C}}$ are distinct. That is, $\operatorname{Conf}_{n_{1},\ldots,n_{k}}$ is the quotient of ordered configuration space of $n_{1}+\cdots+n_{k}$ points by the action of $S_{n_{1}}\times\cdots\times S_{n_{k}}$, where $S_{n_{i}}$ permutes points $n_{1}+\cdots+n_{i-1}+1,\ldots,n_{1}+\cdots+n_{i}$.
Remark 1.6.0.
One might require a weaker notion that $c$ is only conjugation invariant, in the sense that for any $x,y\in c$, we also have $x^{-1}yx\in c$. However, if $c$ generates $G$, such a subset must actually be a union of conjugacy classes. In general, without this assumption, such a conjugation invariant subset can be viewed as a union of conjugacy classes in the subgroup it generates.
Remark 1.6.0.
Because there will not be any connected $G$ covers with inertia in $c$ unless $c$ generates $G$, using notation from § 1.6, $\operatorname{CHur}^{G,c_{1},\ldots,c_{k}}_{n_{1},\ldots,n_{k}}$ will be the empty set unless the conjugacy classes $c_{1},\ldots,c_{k}$ jointly generate $G$.
Our main conjecture on the stable homology of Hurwitz spaces is the following, which predicts that the only stable homology is the “obvious” stable homology.
Conjecture 1.6.0.
Fix an integer $i$. With notation as in § 1.6, suppose all $n_{1},\ldots,n_{k}$ are sufficiently large (with how large they must be depending on $i$) and let $Z\subset\operatorname{CHur}^{G,c_{1},\ldots,c_{k}}_{n_{1},\ldots,n_{k}}$ denote a connected component. Then, the natural map sending a cover to its branch locus $Z\to\operatorname{Conf}_{n_{1},\ldots,n_{k}}$ induces an isomorphism $H_{i}(Z;\mathbb{Q})\to H_{i}(\operatorname{Conf}_{n_{1},\ldots,n_{k}};\mathbb{Q}).$
The special case of the above conjecture where $G=S_{d}$, $k=1$, and $c_{1}$ is the conjugacy class of transpositions is given in [EVW16, Conjecture 1.5]. The above conjecture was suggested to us by Melanie Wood, who in turn pointed out to us that a statement very closely related to this conjecture was formulated in [EVW12, Corollary 5.8].
Remark 1.6.0.
Admittedly, the evidence for § 1.6 in the literature is somewhat limited. The case that $G$ is abelian is fairly trivial, since in that case, the relevant Hurwitz space of $G$-covers is precisely identified with the corresponding configuration space: indeed, in the case $G$ is abelian, all elements correspond to distinct conjugacy classes, so one can read off the total monodromy of the cover from the conjugacy classes of the inertia and the locations of the branch points, and therefore recover the cover. To our knowledge, the only nonabelian cases of this conjecture which are known are those in the present paper, namely those verified in Theorem 1.4.5. However, this conjecture is largely motivated by Malle’s conjecture on counting $G$ extensions of number fields. See [VE10] for a detailed description of this connection.
Remark 1.6.0.
The authors are currently working on proving additional cases of § 1.6.
Remark 1.6.0.
Note that we make § 1.6 for arbitrary $G,c_{1},\ldots,c_{k}$ even though it is not known whether the homology of Hurwitz spaces stabilize at all, except in the case that $k=1$ and $c_{1}$ is non-splitting.
Perhaps more ambitiously, We ask whether the analogous statement could possibly also hold for Chow groups of Hurwitz spaces.
Conjecture 1.6.0.
Fix an integer $i$. With notation as in § 1.6, suppose all $n_{1},\ldots,n_{k}$ are sufficiently large (with how large they must be depending on $i$). Then, the $i$th rational Chow groups of $\operatorname{CHur}^{G,c_{1},\ldots,c_{k}}_{n_{1},\ldots,n_{k}}$ stabilize in the sense that they are independent of $n_{1},\ldots,n_{k}$.
Even more ambitiously, we conjecture that the Chow rings stabilize to those of configuration space.
Conjecture 1.6.0.
In the setting of § 1.6, let $Z\subset\operatorname{CHur}^{G,c_{1},\ldots,c_{k}}_{n_{1},\ldots,n_{k}}$ denote a connected component. Then, the natural map $Z\to\operatorname{Conf}_{n_{1},\ldots,n_{k}}$ sending a cover to its branch locus induces an isomorphism on $i$th rational Chow groups.
Remark 1.6.0.
As pointed out to us by Dan Petersen, the rational Chow group of $\operatorname{Conf}_{n_{1},\ldots,n_{k}}$ is 00 in positive degrees and $\mathbb{Q}$ in degree 00, so if § 1.6 were true, this would give an especially simple description of the stable Chow groups of Hurwitz spaces.
1.7. Idea of the Proof
We now describe the new ideas going into our proof of Theorem 1.1.1, and more generally Theorem 1.2.1, the computation of the moments predicted by the Cohen–Lenstra conjectures over function fields. A standard reduction also outlined in § 1.3.2, which we carry out in § 5, reduces us to proving Theorem 1.4.5.
We next explain the idea of the proof of Theorem 1.4.5, which computes the stable value of the homology for certain Hurwitz spaces. We fix a finite group $G$ and a conjugacy class $c\subset G$ so that $(G,c)$ non-splitting. We use $\operatorname{Hur}^{G,c}$ to denote the union over $n$ of the Hurwitz spaces $\operatorname{Hur}^{G,c}_{n}$ and $\operatorname{Conf}$ to denote the union over $n$ of $\operatorname{Conf}_{n}$. We use $\operatorname{Conf}^{G,c}$ to denote $\operatorname{Conf}\times_{\pi_{0}\operatorname{Conf}}\pi_{0}\operatorname{Hur}^{G,c}$; in other words, it has the same set of components as $\operatorname{Hur}^{G,c}$, but we replace each component of $\operatorname{Hur}^{G,c}_{n}$ with $\operatorname{Conf}_{n}$. There is a projection $v:\operatorname{Hur}^{G,c}\to\operatorname{Conf}^{G,c}$, which can be thought of as sending a cover to its branch locus, together with the data of which component the cover came from. If we use $U$ to denote the stabilization map $\operatorname{Hur}^{G,c}$ (the map under which the homology of Hurwitz space was shown to stabilize in [EVW16, Theorem 6.1]) we can rephrase our goal as showing $v[U^{-1}]:H_{*}(\operatorname{Hur}^{G,c};\mathbb{Q})[U^{-1}]\to H_{*}(\operatorname{Conf}^{G,c};\mathbb{Q})[U^{-1}]$ is an equivalence.
Considering $\operatorname{Hur}^{G,c}$ and $\operatorname{Conf}^{G,c}$ as associative monoids (or $\mathbb{E}_{1}$-algebras), our goal is equivalent to showing that the map $v[U^{-1}]:C_{*}(\operatorname{Hur}^{G,c};\mathbb{Q})[U^{-1}]\to C_{*}(\operatorname{Conf}^{G,c};\mathbb{Q})[U^{-1}]$ is an equivalence of associative ring spectra [^3] Since we are working over $\mathbb{Q}$, these associative ring spectra can be viewed as the dga computing the homology with its multiplication given via the Pontryagin product., where $C_{*}(X;\mathbb{Q})$ denotes the chains of a space $X$ with $\mathbb{Q}$-coefficients, viewed as either a spectrum, or in the derived category of $\mathbb{Q}$-vector spaces.
We can view the center $Z(R)$ of $R:=H_{0}(\operatorname{Hur}^{G,c};\mathbb{Q})$ as a commutative ring acting on the source and target of the map $v$, and we study the map $v[U^{-1}]$ via decomposing it geometrically along open subsets covering $\operatorname{Spec}(Z(R))$. It then suffices to show that locally on each of these subsets the map $v[U^{-1}]$ becomes an equivalence. Algebraically, the way we implement restricting to an open subset of $\operatorname{Spec}(Z(R))$ is by using localizations in the setting of ring spectra (such as in [Lur17, Section 7.2.3]), and completions along a collection of elements, which we show in our case happens to also be a localization. Combining this strategy with a group theoretic argument shows that $v[U^{-1}]$ is an equivalence if $v$ is an equivalence after inverting $\langle c^{\prime}\rangle$, and then completing at $c-c^{\prime}$, for every subset $c^{\prime}\subset c$ closed under conjugation (i.e., $x,y\in c^{\prime}\implies x^{-1}yx\in c^{\prime}$).
When $c^{\prime}=c$, this localized map was already known to be an equivalence. This is because inverting all elements of $c$ gives the group completion of $\operatorname{Hur}^{G,c}$, whose homology can be understood via its classifying space $B\operatorname{Hur}^{G,c}$. The homology of the classifying space of is the rack homology of the rack $c$, which was computed in [EG03], see [RW20, Corollary 5.4].
When $c^{\prime}$ is a proper subset of $c$, the key notion helping us show the map becomes an equivalence is the notion of a homological epimorphism. A map $R\to S$ of $\mathbb{E}_{1}$-rings is called a homological epimorphism if the multiplication map $S\otimes_{R}S\to S$ is an equivalence in the $\infty$-category of spectra. A key result we use is a rigidity property of homological epimorphisms, which is that a homological epimorphism of connective associative ring spectra that is an isomorphism on $\pi_{0}$ is an equivalence. Applying the above results to the subgroup $\langle c^{\prime}\rangle\subset G$ generated by $c^{\prime}$, we can reduce to showing that a certain ‘restriction map’ between the localized homology of Hurwitz spaces is a homological epimorphism. This fact is something that can be verified at the level of pointed spaces, by showing that a certain map between two sided bar constructions is a homotopy equivalence. We use a convenient topological model for these two sided bar constructions, the details of which are carried out in Appendix A. Using this topological model, we write down an explicit homotopy to prove this homotopy equivalence; see Figure 1 for a pictorial depiction.
Remark 1.7.0.
It is also possible to prove Theorem 1.4.5 without the machinery of higher algebra. We originally came up with a more elementary argument but found that by translating things to higher algebra, the argument became significantly cleaner and more conceptual. A version of our elementary argument will appear in [LL24].
1.8. Outline
The sections of this paper are organized as follows. First, we set up our notation for Hurwitz spaces in § 2. In § 3, we introduce notation from higher algebra relevant to our paper, and prove some facts about homological epimorphisms, which will be crucial to our main results. In § 4, we prove our main topological result, Theorem 1.4.5, computing the stable homology of non-splitting Hurwitz spaces. We in fact compute the stable homology more generally of Hurwitz spaces associated to non-splitting racks. In § 5, we prove our main result toward the Cohen–Lenstra heuristics, Theorem 1.2.1, by deducing it from our topological results about the stable homology of Hurwitz spaces. We conclude with Appendix A, which uses various scanning arguments to construct convenient topological models for certain bar constructions on Hurwitz spaces.
1.9. Acknowledgements
We thank Jordan Ellenberg and Craig Westerland for sharing a number of the ideas they tried while working on this problem. The ideas in paper build heavily upon ideas in the retracted preprint [EVW12], and we were able to draw much inspiration from the ideas presented there. We thank Melanie Wood for suggesting § 1.6, listening to a detailed explanation of the argument, and for helpful comments on the paper. We thank Andrea Bianchi for an extremely helpful suggestion which allowed us to remove an extraneous condition we had on Theorem 4.2.1. We thank Yuan Liu for explaining to us the correct version of the Cohen–Lenstra heuristics in the function field case, and for help understanding the non-abelian Cohen–Lenstra heuristics with roots of unity. We also thank Samir Canning, Kevin Chang, Ben Knudsen, Jef Laga, Jeremy Miller, Sun Woo Park, Anand Patel, Sam Payne, Dan Petersen, Eric Rains, Oscar Randal-Williams, Will Sawin, Mark Shusterman, Nathalie Wahl, and Angelina Zheng for helpful conversations. Landesman was supported by the National Science Foundation under Award No. DMS 2102955. Levy was supported by the NSF Graduate Research Fellowship under Grant No. 1745302, and by the Clay Research Fellowship.
2. Notation for Hurwitz spaces
2.1. The definition of Hurwitz spaces
We begin by carefully defining the Hurwitz spaces we will work with. We start by defining the Hurwitz stack $[\operatorname{Hur}^{G,c}_{n,B}/G]$. This will be a quotient by an action of $G$ of the pointed Hurwitz space $\operatorname{Hur}^{G,c}_{n,B}$, which we will define next, and this explains why there is a quotient by $G$ in the notation for $[\operatorname{Hur}^{G,c}_{n,B}/G]$.
Definition 2.1.0.
Let $G$ be a finite group, $c\subset G$ be a union of conjugacy classes, and $B$ be a scheme with $|G|$ invertible on $B$. Let $[\operatorname{Hur}^{G,c}_{n,B}/G]$ denote the Hurwitz stack whose $T$ points $[\operatorname{Hur}^{G,c}_{n,B}/G](T)$ is the groupoid
$\left(D,i:D\to\mathbb{P}^{1}_{T},X,h:X\to\mathbb{P}^{1}_{T}\right)$
satisfying the following conditions:
  1. (1)
    $D$ is a finite étale cover of $T$ of degree $n$.
  2. (2)
    $i$ is a closed immersion $i:D\subset\mathbb{A}^{1}_{T}\subset\mathbb{P}^{1}_{T}$.
  3. (3)
    $X$ is a smooth proper relative curve over $T,$ not necessarily having geometrically connected fibers.
  4. (4)
    $h:X\to\mathbb{P}^{1}_{T}$ is a finite locally free Galois $G$-cover, (meaning that $G$ acts simply transitively on the geometric generic fiber of $h$,) which is étale away from $\infty_{T}\cup i(D)\subset\mathbb{P}^{1}_{T}$.
  5. (5)
    The inertia of $X\to\mathbb{P}^{1}_{T}$ over any geometric point of $i(D)$ lies in $c\subset G$.
  6. (6)
    Two such covers are considered equivalent if they are related by the $G$-conjugation action.
  7. (7)
    The morphisms between two points $(D_{i},i_{i},X_{i},h_{i})$ for $i\in\{1,2\}$ are given by $(\phi_{D},\psi_{X})$ where $\phi_{D}:D_{1}\simeq D_{2}$ is an isomorphism so that $i_{2}\circ\phi_{D}=i_{2}$ and $\psi_{X}:X_{1}\simeq X_{2}$ is an isomorphism such $h_{2}\circ\psi_{X}=h_{1}$ and $\psi_{X}=g^{-1}\psi_{X}g$ for every $g\in G$.
Remark 2.1.0.
The definition of Hurwitz spaces we give here is a special case of the definition in [EL23, Definition 2.4.2]. See [EL23, Remark 2.4.3] for an explanation of why these Hurwitz stacks are indeed stacks. It essentially follows from [ACV03, §1.3.2 and Appendix B].
One important idea in this paper is to work with pointed Hurwitz spaces, which deal with the case of “real quadratic” or “ramified at $\infty$” quadratic fields. Variants of these were also used in [EL23], but prior work in the context of the Cohen–Lenstra heuristics appears to primarily focus on the case over covers unramified at $\infty$. These Hurwitz spaces parameterize covers of a stacky $\mathbb{P}^{1}$, with a root stack at $\infty$ of order $2$. We now define these pointed Hurwitz spaces.
Definition 2.1.0.
Fix an integer $w$ and define $\mathscr{P}^{w}$ to be the root stack of order $w$ along $\infty$ of $\mathbb{P}^{1}$. (See [Cad07, Definition 2.2.4], where this is notated as (,$\mathbb{P}^{1}_{(\infty,w)}$.) The fiber of this root stack over $\infty$ is the stack quotient $[\left(\operatorname{Spec}_{B}\mathscr{O}_{B}[x]/(x^{r})\right)/\mu_{r}]$ of the relative spectrum $\operatorname{Spec}_{B}\mathscr{O}_{B}[x]/(x^{r})$ by $\mu_{r}$. Let ~$\widetilde{\infty}:B\to\mathscr{P}^{w}$ denote the section over $\sigma$ corresponding to map $B\to[\left(\operatorname{Spec}_{B}\mathscr{O}_{B}[x]/(x^{r})\right)/\mu_{r}]$ given by the trivial $\mu_{r}$ torsor over $B$, $\mu_{r}\to B$, and the $\mu_{r}$ equivariant map $\mu_{r}\to B\to\operatorname{Spec}_{B}\mathscr{O}_{B}[x]/(x^{r})$. We use notation as in § 2.1,
Define the $w$-pointed Hurwitz space, $\left(\operatorname{Hur}^{G,c}_{n,B}\right)^{w}$, to be the algebraic space whose $T$ points are the set parameterizing data of the form
$\left(D,h^{\prime}:X\to\mathscr{P}^{w}_{T},t:T\to X\times_{h^{\prime},\mathscr{P}^{w}_{T},\widetilde{\infty}_{T}}T,i:D\to\mathbb{P}^{1}_{T},X,h:X\to\mathbb{P}^{1}_{T}\right),$
where $D,i,X,$ and $h$ satisfy the properties listed in § 2.1. We also assume the order of inertia of $h$ along $\infty$ is $w$ and define ~$\widetilde{\infty}_{T}$ to be the base change of the section $\widetilde{\infty}$ defined above to $T$. We additionally impose the condition that $h^{\prime}$ is a finite locally free $G$-cover, étale over $\widetilde{\infty}$, such that the composition of $h^{\prime}:X\to\mathscr{P}^{w}_{T}$ with the coarse space map $\mathscr{P}^{w}_{T}\to\mathbb{P}^{1}_{T}$ is $h$, and $t:T\to X\times_{h^{\prime},\mathscr{P}^{w}_{T},\widetilde{\infty}_{T}}T$ is a section of $h^{\prime}$ over $\widetilde{\infty}$.
In general, we define the pointed Hurwitz space as $\operatorname{Hur}^{G,c}_{n,B}:=\coprod_{w\geq 1}\left(\operatorname{Hur}^{G,c}_{n,B}\right)^{w}.$
Remark 2.1.0.
We note that $\operatorname{Hur}^{G,c}_{n,B}$, defined as an algebraic space, is in fact a scheme. Indeed, $\operatorname{Hur}^{G,c}_{n,B}$ is a finite étale cover of the configuration space of $n$ unordered points in $\mathbb{A}^{1}_{B}$, as can be verified in an analogous fashion to [LWZB24, Proposition 11.4].
Remark 2.1.0.
In the definition of pointed Hurwitz space, § 2.1, we choose a particular section ~$\widetilde{\infty}:B\to\mathscr{P}^{w}$ over $\infty:B\to\mathscr{P}^{1}_{B}$. However, one can show that if we chose a different section ~$\widetilde{\infty}^{\prime}:B\to\mathscr{P}^{w}$, the resulting pointed Hurwitz space would be isomorphic. The idea is that the sections $\widetilde{\infty}$ and $\widetilde{\infty}^{\prime}$ factor respectively through maps $\alpha:B\to[B/\mu_{w}]$ and $\alpha^{\prime}:B\to[B/\mu_{w}]$, and so it suffices to produce a map $\beta:[B/\mu_{w}]\to[B/\mu_{w}]$ so that $\alpha\circ\beta=\alpha^{\prime}$, as this map will induce a map taking $\widetilde{\infty}$ to $\widetilde{\infty}^{\prime}$, which then induces an isomorphism of these two Hurwitz spaces. The desired map $\beta$ is simply given as the composition of the structure map $[B/\mu_{w}]\to B$ with $\alpha^{\prime}$.
Remark 2.1.0.
Note that monic smooth hyperelliptic curves $y^{2}=f(x)$ with $f$ of odd degree correspond to the choice of marked section given in § 2.1 while non-monic smooth hyperelliptic curves $y^{2}=f(x)$ with $f$ of with the same odd degree correspond to a different choice of marked section, see § 2.3. Hence, the two relevant Hurwitz spaces will be isomorphic by § 2.1. From this, one can obtain the fact that the distributions of class groups of monic smooth hyperelliptic curves of odd degree are isomorphic to the distribution of class groups of non-monic smooth hyperelliptic curves of odd degree. This was mentioned in § 1.1.
There is also a more elementary way to see this equivalence, as we learned from [BW17, Corollary 6.5]. This can be viewed as a reformulation of the argument in the previous paragraph. Namely, there is a bijection between monic smooth hyperelliptic curves and non-monic smooth hyperelliptic curves which additionally preserves their isomorphism type, and hence their class groups. For any $\alpha\in\mathbb{F}_{q}-(\mathbb{F}_{q})^{2}$ every non-monic smooth hyperelliptic curve can be written as $y^{2}=f(x)$, where $f(x)$ has leading coefficient $\alpha$. Using this representation of such non-monic smooth hyperelliptic curves, the map is given by sending a quadratic extension corresponding to a monic smooth hyperelliptic curve $K$ over $\mathbb{F}_{q}(t)$ to the quadratic extension corresponding to a non-monic smooth hyperelliptic curve $K\otimes_{\mathbb{F}_{q}(t)}\mathbb{F}_{q}(s)$ where $t\mapsto\alpha s$.
Notation 2.1.0.
We use $\operatorname{CHur}^{G,c}_{n,B}$ to denote the open and closed subscheme of $\operatorname{Hur}^{G,c}_{n,B}$ parameterizing covers $X\to\mathbb{P}^{1}_{T}$ so that the geometric fibers of $X$ over $T$ are connected. We use $[\operatorname{CHur}^{G,c}_{n,B}/G]$ to denote the open and closed substack of $[\operatorname{Hur}^{G,c}_{n,B}/G]$ parameterizing covers $X\to\mathbb{P}^{1}_{T}$ so that the geometric fibers of $X$ over $T$ are connected.
Notation 2.1.0.
We will often work over the complex numbers, and hence it will be convenient to assume $B=\operatorname{Spec}\mathbb{C}$. Therefore, we define $[\operatorname{Hur}^{G,c}_{n}/G]:=[\operatorname{Hur}^{G,c}_{n,\operatorname{Spec}\mathbb{C}}/G],\operatorname{Hur}^{G,c}_{n}:=\operatorname{Hur}^{G,c}_{n,\operatorname{Spec}\mathbb{C}},[\operatorname{CHur}^{G,c}_{n}/G]:=[\operatorname{CHur}^{G,c}_{n,\operatorname{Spec}\mathbb{C}}/G],\operatorname{CHur}^{G,c}_{n}:=\operatorname{CHur}^{G,c}_{n,\operatorname{Spec}\mathbb{C}}$. When $R$ is a ring, we often use use $\operatorname{Hur}^{G,c}_{n,R}$ in place of $\operatorname{Hur}^{G,c}_{n,\operatorname{Spec}R},$ and use analogous notation for other variants of this Hurwitz space.
2.2. Notation for components of Hurwitz spaces
We now introduce some notation for the Hurwitz spaces we will work with. First, we describe the notion of boundary monodromy of a component. If we think of the component as parameterizing certain covers $X\to\mathbb{P}^{1}$, the boundary monodromy corresponds to the inertia of the cover over $\infty$.
Definition 2.2.0.
With notation as in § 2.1, there is a map from the set of components of $\operatorname{Hur}^{G,c}_{n,B}$ to $G$ defined as follows. We can represent connected components of $\operatorname{Hur}^{G,c}_{n,B}$ as tuples $(g_{1},\ldots,g_{n})$ modulo the action of the braid group $B_{n}$, and we associate to this the element $g_{1}\cdots g_{n}\in G$. For such a component mapping to $g\in G$, we say the boundary monodromy of this component is $g$. We define $\operatorname{CHur}^{G,c,g}_{n,B}$ denote the open and closed subscheme of $\operatorname{CHur}^{G,c}_{n,B}$ consisting of those components whose boundary monodromy is $g$.
2.3. Covers split completely at infinity
Definition 2.3.0.
Let $X\to\mathbb{P}^{1}_{\mathbb{F}_{q}}$ be a Galois $G$ cover with branch divisor of degree $n$ and inertia over $\mathbb{A}^{1}_{\mathbb{F}_{q}}$ all lying in a union of conjugacy classes $c\subset G$. Hence, this cover corresponds to a point $\operatorname{Spec}\mathbb{F}_{q}\to[\operatorname{Hur}^{G,c}_{n,\mathbb{F}_{q}}/G]$. We say this cover is completely split over $\infty$ if it is in the image of the map $\operatorname{Hur}^{G,c}_{n,\mathbb{F}_{q}}(\mathbb{F}_{q})\to[\operatorname{Hur}^{G,c}_{n,\mathbb{F}_{q}}/G](\mathbb{F}_{q})$.
Next, we wish to verify that monic smooth hyperelliptic curves precisely correspond to smooth covers split over infinity. This is fairly straightforward in the case $n$ is even. It is a bit more involved when $n$ is odd, but we are able to verify this via an explicit blow-up computation. For the next lemma, recall that a monic smooth hyperelliptic curve over $\mathbb{F}_{q}$ is by definition a hyperelliptic curve defined affine locally by an equation of the form $y^{2}=f(x)$ with $f(x)=x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0}$ for $a_{i}\in\mathbb{F}_{q}$. Explicitly, we view the hyperelliptic curve as a cover of $\mathbb{P}^{1}_{\mathbb{F}_{q}}$ with $x$ as above the coordinate on $\mathbb{A}^{1}_{\mathbb{F}_{q}}$, viewed as a subscheme of $\mathbb{P}^{1}=\mathbb{P}(\mathbb{F}_{q}[X,Y])$ with $x=X/Y$.
Lemma 2.3.0.
Take $G=\mathbb{Z}/2\mathbb{Z}$ and $c$ to be the nontrivial element of $\mathbb{Z}/2\mathbb{Z}$ in § 2.3. The $\mathbb{Z}/2\mathbb{Z}$ smooth covers of $\mathbb{P}^{1}_{\mathbb{F}_{q}}$ split completely over the point $\infty\in\mathbb{P}^{1}_{\mathbb{F}_{q}}$ precisely correspond to monic smooth hyperelliptic curves $y^{2}=f(x)$ over $\mathbb{F}_{q}$.
Proof.
First, we describe the case that $n$ is even. Any hyperelliptic curve in characteristic not $2$ which is unramified over $\infty$ can be uniquely written in the form $f(x)=\alpha x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0}$ where either $\alpha=1$ or $\alpha\in\mathbb{F}_{q}^{\times}$ is a fixed non-square. Such a curve is split completely over infinity if and only if its fiber over $\infty\in\mathbb{P}^{1}_{\mathbb{F}_{q}}(\mathbb{F}_{q})$ consists of two $\operatorname{Spec}\mathbb{F}_{q}$ points. The fiber over $\infty$ of this curve can be identified with the fiber over 00 of the curve $z^{2}=\alpha+a_{n-1}u+\cdots+a_{1}u^{n-1}+a_{0}u^{n}$, where we have set $u=1/x$ and then $z=y/x^{n/2}$. This is the same as $\operatorname{Spec}\mathbb{F}_{q}[y]/(y^{2}-\alpha)$, which will be two copies of $\mathbb{F}_{q}$ when $\alpha$ is a square and a single copy of $\mathbb{F}_{q}^{2}$ when $\alpha$ is a non-square. Therefore, the curves where $\alpha$ is a square (which are all isomorphic to such curves with $\alpha=1$) precisely correspond to those completely split over $\infty$.
We next consider the case that $n$ is odd. As above, we can still write our hyperelliptic curve in the form $f(x)=\alpha x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0}$ with $\alpha$ either $1$ or a fixed non-square. Changing coordinates to $u=1/x$, we can write the curve in the form $z^{2}=u(\alpha+a_{n-1}u+\cdots+a_{1}u^{n-1}+a_{0}u^{n})$ where $u=1/x$ and $z=y/x^{n+1/2}$. Let $X$ denote the proper regular curve corresponding to the above equation. We can identify the stack quotient $[X/(\mathbb{Z}/2\mathbb{Z})]$ Zariski locally around the residual gerbe at $\infty$ with a Zariski open of the root stack $\mathscr{P}$, which is a root stack of $\mathbb{P}^{1}$ over the divisor $\infty$, so $\mathscr{P}$ has a single stacky point of order $2$.
Let ~$\widetilde{\infty}:\operatorname{Spec}\mathbb{F}_{q}\to\mathscr{P}$ denote the map which factors through the residual gerbe $B\mu_{2}$ at infinity and corresponds to the map $\operatorname{Spec}\mathbb{F}_{q}\to B\mu_{2}$ associated to the trivial double cover $\operatorname{Spec}\mathbb{F}_{q}\coprod\operatorname{Spec}\mathbb{F}_{q}\to\operatorname{Spec}\mathbb{F}_{q}$. We wish to show that the fiber product $X\times_{\mathscr{P},\infty}\operatorname{Spec}\mathbb{F}_{q}$ consists of two copies of $\operatorname{Spec}\mathbb{F}_{q}$. The fiber product above is identified with the fiber product $X\times_{\mathscr{P}}\mathbb{P}^{1}\times_{\mathbb{P}^{1},\infty}\operatorname{Spec}\mathbb{F}_{q}$, where the map $\mathbb{P}^{1}\to\mathscr{P}$ is the double cover branched over 00. (In particular, on coarse spaces $\mathbb{P}^{1}\to\mathscr{P}$ corresponds to the map given in local coordinates by $x\mapsto x^{2}$.) The above fiber product $X\times_{\mathscr{P}}\mathbb{P}^{1}$ is a normal curve in a neighborhood over $\infty$, and hence it is, locally around $\infty$, the normalization of $X\times_{\mathbb{P}^{1}_{u}}\mathbb{P}^{1}_{t}$, where again the map $\mathbb{P}^{1}_{t}\to\mathbb{P}^{1}_{u}$ is induced by the double cover $u\mapsto t^{2}$. In a neighborhood of $\infty$, we can express the projection map $X\times_{\mathbb{P}^{1}_{u}}\mathbb{P}^{1}_{t}\to X$ as
$\operatorname{Spec}k[u,z,t]/(t^{2}-u,z^{2}-u(\alpha+a_{n-1}u+\cdots+a_{1}u^{n-1}+a_{0}u^{n})$
$\to\operatorname{Spec}k[u,z]/(z^{2}-u(\alpha+a_{n-1}u+\cdots+a_{1}u^{n-1}+a_{0}u^{n}).$
The source is then isomorphic to
$\operatorname{Spec}k[t,z]/(z^{2}-t^{2}(\alpha+a_{n-1}t^{2}+\cdots+a_{1}t^{2(n-1)}+a_{0}t^{2n}).$
This is not normal, but its blow up at the origin is normal in a neighborhood of the origin, and its blow up is isomorphic to
$\operatorname{Spec}k[Z/U,t]/((Z/U)^{2}-(\alpha+a_{n-1}t^{2}+\cdots+a_{1}t^{2(n-1)}+a_{0}t^{2n}).$
The fiber of this over the point $z=t=0$ is then simply $k[Z/U]/((Z/U)^{2}-\alpha),$ which is a copy of $\mathbb{F}_{q^{2}}$ if $\alpha$ is a non-square, while it is two copies of $\mathbb{F}_{q}$ if $\alpha$ is a square. This fiber product is precisely the fiber product $X\times_{\mathscr{P},\infty}\operatorname{Spec}\mathbb{F}_{q}$ we wished to compute, and so we are done. ∎
3. Algebra setup
The goal of this section is to recall and set up some of the algebraic machinery that we use. After introducing some of our notation that we use from higher category theory in § 3.1, we study the notion of a homological epimorphism in § 3.2, which is key to our study of stable homology. Throughout the rest of the section, we collect a number of basic facts in higher algebra, primarily related to homological epimorphisms. In § 3.3 we check localization and certain completions are homological epimorphisms. In § 3.4 we record a rigidity property for homological epimorphisms, which shows in many cases that they are actually equivalences. Finally, in § 3.5 we prove some well known lemmas about nilpotence of endomorphisms.
An alternate, more elementary perspective on the calculations in this section and the next section is given in [LL24] in the case that $G$ is a dihedral group of order twice an odd prime, and $c$ is the conjugacy class of order $2$ elements. This may be helpful for the reader to consult, especially if they are less familiar with higher algebra.
3.1. Algebra notations and recollections
Here, we recall some basic notions from higher algebra that we use here to study the rational homology of Hurwitz spaces. A convenient foundation for studying objects in homotopy theory is that of $\infty$-categories, as developed by Joyal and Lurie (see [Lur09], [Lur17] or [Lur18]). If the reader prefers, they may also interpret the results below as about differential graded algebras: the use of spectra is not necessary to what we are doing here, and working with the derived category of the rational numbers suffices.
3.1.1. The infinity category of spaces
We use $\operatorname{Spc}$ to denote the $\infty$-category of spaces, which we view as symmetric monoidal with respect to the product. We use $\operatorname{Spc}_{*}$ to denote the $\infty$-category of pointed spaces, which is symmetric monoidal with respect to the smash product $\wedge$. Recall that for $(X,x)$ and $(Y,y)$ two pointed spaces, $X\wedge Y$ is defined as the quotient of $X\times Y$ by collapsing the subspace $x\times Y\cup y\times X$. We have a symmetric monoidal functor $(-)_{+}:\operatorname{Spc}\to\operatorname{Spc}_{*}$ given by adding a disjoint base point to a space.
There is a functor $L:\mathrm{Top}\to\operatorname{Spc}$ from the category of topological spaces to the $\infty$-category of spaces, obtained by universally inverting all morphisms that are weak homotopy equivalences.
3.1.2. $\mathbb{E}_{1}$ algebras
Given a symmetric monoidal $\infty$-category ${\mathcal{C}}$, we can form the symmetric monoidal $\infty$-category $\operatorname{Alg}({\mathcal{C}})$ of associative (or $\mathbb{E}_{1}$) algebras in ${\mathcal{C}}$.
3.1.3. Stable $\infty$ categories
We recall that an $\infty$-category ${\mathcal{C}}$ is stable if it admits finite limits and finite colimits, a zero object, and pullback squares and pushout squares coincide. We use $\operatorname{Sp}$ to denote the symmetric monoidal stable $\infty$-category of spectra. Given a stable $\infty$-category ${\mathcal{C}}$ and objects $X,Y\in{\mathcal{C}}$, we use $\operatorname{map}_{{\mathcal{C}}}(X,Y)$ to denote the spectrum of maps between $X$ and $Y$.
3.1.4. $\mathbb{E}_{1}$ rings
We refer to an $\mathbb{E}_{1}$-algebra in spectra as an $\mathbb{E}_{1}$-ring. An ordinary associative ring $R$ can be viewed as an $\mathbb{E}_{1}$-ring via the lax symmetric monoidal inclusion of abelian groups into spectra. All of the $\mathbb{E}_{1}$-rings that we will be interested in are actually $\mathbb{E}_{1}$$-\mathbb{Q}$-algebras, so the reader may interpret the phrase $\mathbb{E}_{1}$-ring to mean differential graded algebra up to quasi-isomorphism.
Given an $\mathbb{E}_{1}$-ring, we can form the stable $\infty$-categories $\operatorname{Mod}_{\operatorname{Sp}}(R)$ and $\operatorname{Mod}_{\operatorname{Sp}}(R^{\operatorname{op}})$ of left and right $R$-modules in spectra respectively. For example, when $R$ is an ordinary associative ring, then $\operatorname{Mod}_{\operatorname{Sp}}(R)$ agrees with the derived $\infty$-category of $R$-modules, so that objects can be presented as (unbounded) chain complexes of $R$-modules. More generally given a DGA $R$, objects of $\operatorname{Mod}_{\operatorname{Sp}}(R)$ can be presented as differential graded modules over $R$ up to quasi-isomorphism.
Given $M\in\operatorname{Mod}_{\operatorname{Sp}}(R)$, we use either $\pi_{i}M$ or $H_{i}M$ depending on the context to refer to $\pi_{i}$ of the underlying spectrum of $M$. In the case $R$ is an ordinary ring and $M$ is presented as a chain complex of $R$-modules, this agrees with the $i^{\mathrm{th}}$ homology of the chain complex.
3.1.5. $\mathbb{E}_{2}$-central elements
Let $R$ be an $\mathbb{E}_{1}$ ring. We say $r\in\pi_{i}R$ is $\mathbb{E}_{2}$-central if there is an $R$-bimodule map $\Sigma^{i}R\to R$ sending $1$ to $r$. In particular, if $r$ is $\mathbb{E}_{2}$-central, it is also graded commutative in the homotopy ring $\pi_{*}R$ of $R$. Given an $\mathbb{E}_{2}$-central element $r$, we implicitly associate to it a choice of $R$-bimodule map refining $r$.
3.1.6. Chains functor
We have a symmetric monoidal functor $\tilde{C}_{*}(-;\mathbb{Q}):\operatorname{Spc}_{+}\to\operatorname{Sp}$, taking $A$ to its reduced chains with coefficients in $\mathbb{Q}$, with the comultiplication coming from the diagonal map, which is dual to the cup product. Composing $\tilde{C}_{*}(-;\mathbb{Q})$ with the symmetric monoidal functor $(-)_{+}$, we get the symmetric monoidal functor $C_{*}(-;\mathbb{Q}):\operatorname{Spc}\to\operatorname{Sp}$. For $X$ a space, $H_{i}(X;\mathbb{Q})=\pi_{i}C_{*}(X;\mathbb{Q})$.
3.2. Homological epimorphisms
Our goal is to study the algebra of chains on Hurwitz spaces, which forms an $\mathbb{E}_{1}$-ring. We will study this $\mathbb{E}_{1}$-ring by forming certain completions and localizations of it. An important property of these operations that makes them well behaved in this setting is that they are homological epimorphisms. To define this notion, we recall the following lemma:
Lemma 3.2.0 ([NK24, Lemma 2.9.8]).
For a map $R\to S$ of $\mathbb{E}_{1}$-rings with fiber $I$, the following conditions are equivalent:
  1. (1)
    The forgetful functor $\operatorname{Mod}_{\operatorname{Sp}}(S)\to\operatorname{Mod}_{\operatorname{Sp}}(R)$ is fully faithful.
  2. (2)
    The multiplication map $S\otimes_{R}S\to S$ is an equivalence.
  3. (3)
    $I\otimes_{R}S=0$
  4. (4)
    The multiplication map induces an equivalence $I\otimes_{R}I\cong I$.
Definition 3.2.0.
A map $R\to S$ of $\mathbb{E}_{1}$-rings is called a homological epimorphism (or hom. epi) if any of the equivalent conditions of § 3.2 are satisfied.
3.3. Localization and completion
Two sources of hom. epis for us will be certain localization and completion maps. We recall below that given an $\mathbb{E}_{1}$-ring $R$, some $i\in\mathbb{Z}$, and $r\in\pi_{i}R$, we can form the completion $M^{\wedge}_{r}$, and localization $M[r^{-1}]$, of an $M\in\operatorname{Mod}_{\operatorname{Sp}}R$, and that this forms an $\mathbb{E}_{1}$-ring when $M=R$.
[…]
Example 3.3.0.
The map $R\to R[r^{-1}]$ is always a hom. epi [NK24, Example 2.9.14]. More generally for a set of elements $S$ in the homotopy groups of $R$, we can form $M[S^{-1}]$, which can be described as the localization away from $R/r,r\in S$, and when $S$ is a finite set, agrees with iteratively localizing away from each element of $S$ in any order.
Remark 3.3.0.
When $r\in\pi_{i}R$ satisfies the left Ore condition, as defined in [Lur17, Definition 7.2.3.1], the underlying spectrum of $M[r^{-1}]$ can be computed as the colimit of $M$ along the left multiplication by $r$ [Lur17, Proposition 7.2.3.20].
We will also want to show certain completions are homological epimorphisms. We next define completions, and later, in § 3.3, we will show certain completions are homological epimorphisms by showing they are localizations.
Definition 3.3.0.
For $r\in\pi_{i}R$, the $r$-completion of the $\infty$-category of $R$-modules is defined as the presentable localization of $\operatorname{Mod}_{\operatorname{Sp}}R$ away from $R[r^{-1}]$. In other words, $\operatorname{Mod}_{\operatorname{Sp}}(R)^{\wedge}_{r}$ is the full subcategory of $R$-modules such that $\operatorname{map}(R[r^{-1}],-)$ vanishes. The inclusion of this subcategory $\operatorname{Mod}_{\operatorname{Sp}}(R)^{\wedge}_{r}\subset\operatorname{Mod}_{\operatorname{Sp}}R$ has a left adjoint sending $M$ to $M^{\wedge}_{r}$.
There are natural $\mathbb{E}_{1}$-ring maps $R\to R^{\wedge}_{r}$, $R\to R[r^{-1}]$ induced by taking the map induced on endomorphism rings of the $R$-module $R$ via the completion and localization functor. Note also that $R[r^{-1}]=R[(r^{n})^{-1}]$ and $R^{\wedge}_{r}=R^{\wedge}_{r^{n}}$ for each $n\geq 1$.
The following lemma is a standard manipulation of Bousfield localizations. See [Rav84, Lemma 1.34] for a closely related statement.
Lemma 3.3.0.
Let $R$ be an $\mathbb{E}_{1}$-ring. Suppose that $f:M\to N$ is a map of $R$-modules and $r\in\pi_{i}R$ is an element. The following are equivalent:
  1. (1)
    $f$ is an equivalence.
  2. (2)
    $f[\frac{1}{r}]$ and $f^{\wedge}_{r}$ are equivalences.
  3. (3)
    $f[\frac{1}{r}]$ and $R/r\otimes_{R}f$ are equivalences.
Proof.
We have $(1)\implies(2)$ because localizations and completions preserve equivalences. Next, $(2)\iff(3)$ follows since $M$ is an $R$ module with $r$ acting invertibly if and only if $(R/r)\otimes_{R}M=0$.
Lastly, we show $(3)\implies(1)$. If $(3)$ is true, then $\operatorname{map}_{\operatorname{Mod}_{\operatorname{Sp}}(R)}(R/r,M)\to\operatorname{map}_{\operatorname{Mod}_{\operatorname{Sp}}(R)}(R/r,N)$ is an equivalence, as is $\operatorname{map}_{\operatorname{Mod}_{\operatorname{Sp}}(R)}(R[r^{-1}],M)\to\operatorname{map}_{\operatorname{Mod}_{\operatorname{Sp}}(R)}(R[r^{-1}],N)$. Hence, $R$ is an extension of $R[r^{-1}]$ by $\operatorname{fib}(R\to R[r^{-1}])$, which is generated under colimits by $R/r$, by definition of the Bousfield localization $(-)[r^{-1}]$. So, it follows that $M\cong\operatorname{map}_{\operatorname{Mod}_{\operatorname{Sp}}(R)}(R,M)\cong\operatorname{map}_{\operatorname{Mod}_{\operatorname{Sp}}(R)}(R,N)\cong N$ is an equivalence. ∎
The following lemma gives a situation under which completion happens to invert a central idempotent.
Lemma 3.3.0.
Let $R$ be an $\mathbb{E}_{1}$-ring, and $x\in\pi_{0}R$. Suppose there is some integers $n,m$ with $n>m$ and some $z\in\pi_{0}R$ so that $zx^{n}=x^{m}$. Also assume $x^{m^{\prime}}$ is central for some $m^{\prime}>0$.
Then, $e:=z^{nm^{\prime}}x^{n(n-m)m^{\prime}}$ is a central idempotent in $\pi_{*}R$ so that $R\cong R[e^{-1}]\times R[(1-e)^{-1}]$ as an $\mathbb{E}_{1}$-ring. We also have $R^{\wedge}_{x}\cong R[(1-e)^{-1}],R[e^{-1}]\cong R[x^{-1}]$, so that $R\to R^{\wedge}_{x}$ is in particular a localization. Moreover, $\pi_{*}R^{\wedge}_{x}$ is the quotient of $\pi_{*}R$ by the two sided ideal generated by $x^{m}$.
Proof.
To see that $e=z^{nm^{\prime}}x^{n(n-m)m^{\prime}}$ is an idempotent, we have
$(z^{nm^{\prime}}x^{n(n-m)m^{\prime}})^{2}=z^{2nm^{\prime}}x^{2n(n-m)m^{\prime}}=z^{nm^{\prime}}x^{n(n-m)m^{\prime}},$
where, the first equality uses that $x^{m^{\prime}}$ is central and the second equality uses the equation $zx^{n}=x^{m}$ a total of $nm^{\prime}$-times and the assumption that $n(n-m)m^{\prime}\geq n$.
We next show that $e$ is central. Indeed, for an arbitrary element $y\in\pi_{*}R$, we have
$ey$ $=z^{nm^{\prime}}x^{n(n-m)m^{\prime}}y$
$=z^{nm^{\prime}}yx^{n(n-m)m^{\prime}}$
$=z^{nm^{\prime}}yz^{nm^{\prime}}x^{2n(n-m)m^{\prime}}$
$=z^{nm^{\prime}}x^{2n(n-m)m^{\prime}}yz^{nm^{\prime}}$
$=x^{n(n-m)m^{\prime}}yz^{n}$
$=ye.$
It follows that the localizations $R[e^{-1}],R[(1-e)^{-1}]$ are Ore localizations, so localizations can be computed as in § 3.3.
Since $R$ is the pullback $R[e^{-1}]\times_{R[e^{-1},(1-e)^{-1}]}R[1-e^{-1}]$, and $R[e^{-1},(1-e)^{-1}]=0$, we have $R\cong R[e^{-1}]\times R[(1-e)^{-1}]$. Since $e$ is sent to $1$ in $R[x^{-1}]$, and $x$ is invertible in $R[e^{-1}]$, we have $R[x^{-1}]=R[x^{-1},e^{-1}]=R[e^{-1}]$. Since the completion is the localization away from $R[x^{-1}]$, it follows that the completion is the other summand $R^{\wedge}_{x}\cong R[(1-e)^{-1}]$. To see that $x^{m}=0$ in the completion, we have
$(1-e)x^{m}=x^{m}-z^{nm^{\prime}}x^{n(n-m)m^{\prime}+m}=z^{nm^{\prime}}x^{n(n-m)m^{\prime}+m}-z^{nm^{\prime}}x^{n(n-m)m^{\prime}+m}=0.$
Because $x^{m}$ maps to 00 in the component $R[(1-e)^{-1}]\simeq R^{\wedge}_{x}$, the kernel of surjective map $\pi_{*}R\to\pi_{*}R^{\wedge}_{x}$ contains two sided ideal generated by $x^{m}$. From the above, this kernel is equal to $\pi_{*}R[x^{-1}]$. Since $\pi_{*}R[x^{-1}]$ is contained in the two sided ideal generated by $x^{m}$ since any $y\in\pi_{*}R[x^{-1}]$ can be written as $yx^{-m}x^{m}$, we obtain that the kernel is equal to the two sided ideal generated by $x^{m}$. ∎
Notation 3.3.0.
Suppose that $R$ is an $\mathbb{E}_{1}$-ring and $X:=\{x_{1},\ldots,x_{|X|}\}$ is a finite set of elements in $\pi_{0}R$ each independently satisfying the conditions of § 3.3. That is, for each $i$ with $1\leq i\leq|X|$, there are some $n_{i},m_{i}$ and $m_{i}^{\prime}$ with $n_{i}>m_{i}$ and $z\in\pi_{0}R$ so that $zx^{n_{i}}=x_{m_{i}}$ and $x^{m_{i}^{\prime}}$ is central. Then we use $R^{\wedge}_{X}$ to denote the iterated completion $(\cdots(R^{\wedge}_{x_{1}})\cdots)^{\wedge}_{x_{|X|}}$. By § 3.3, this doesn’t depend on the choice of ordering, since each completion is also a localization (see § 3.3).
3.4. Tools for homological epimorphisms
Lemma 3.4.0.
Suppose we have maps $R\to S\to S^{\prime}$ such that $R\to S$ is a hom. epi. Then $R\to S^{\prime}$ is a hom. epi iff $S\to S^{\prime}$ is a hom. epi.
Proof.
We use the description in § 3.2(1) of hom. epi. We have the composite of forgetful functors
$\operatorname{Mod}(S^{\prime})\to\operatorname{Mod}(S)\to\operatorname{Mod}(R),$
where the second is fully faithful. It follows that the composite is fully faithful iff the first functor is. ∎
The following proposition is the key to relate the notion of a hom. epi to homological stability. Heuristically, it says that hom. epis to connective rings are determined by their $\pi_{0}$. See [HS24, Theorem B] for a generalization of this.
Proposition 3.4.0 (Rigidity of homological epimorphisms).
Consider a diagram of $\mathbb{E}_{1}$-rings
[…]
where $S\to S^{\prime}$ is a map of connective $\mathbb{E}_{1}$-rings that is an isomorphism on $\pi_{0}$, and where $R\to S$ and $R\to S^{\prime}$ are homological epimorphisms. Then the map $S\to S^{\prime}$ is an equivalence.
Proof.
Using § 3.4, we learn that $S\to S^{\prime}$ is a homological epimorphism. Letting $I$ be the fiber of $S\to S^{\prime}$, we have $I\otimes_{S}S^{\prime}=0$. $I$ is the fiber of a map of connective $\mathbb{E}_{1}$-rings, so it is $-1$-connective. We will inductively show on $i\geq-1$ that $\pi_{i}I$ is 00. The inductive hypothesis implies that $I$ is $i$-connective, so and since $S\to S^{\prime}$ is an equivalence on $\pi_{0}$, we claim
$\pi_{i}(I)\cong\pi_{i}(I\otimes_{S}S)\cong\pi_{i}(I\otimes_{S}\pi_{0}(S))\cong\pi_{i}(I\otimes_{S}\pi_{0}(S^{\prime}))\cong\pi_{i}(I\otimes_{S}S^{\prime})\cong 0.$
Indeed, the second and fourth equivalences come from the fact that $I$ is $i$-connective and the maps $S^{\prime}\to\pi_{0}S$ and $S\to\pi_{0}S$ are $1$-connective, so the map on tensor products is $i+1$-connective (see [Lev22, Lemma 3.1]). The first equivalence is tautological, and the third and fifth equivalences follow by our assumptions. Thus $\pi_{i}I=0$ for all $I$, so $I=0$, and hence $S\to S^{\prime}$ is an isomorphism. ∎
The next lemma gives a criterion for being a hom. epi which is often easier to check in practice.
Lemma 3.4.0.
Let $f:R\to S$ be a map of connective $\mathbb{E}_{1}$-rings, and suppose that the map $\pi_{0}S\otimes_{R}\pi_{0}S\to\pi_{0}S\otimes_{S}\pi_{0}S$ is an equivalence. Then $f$ is a hom. epi.
Proof.
Consider the subcategory $C$ of pairs $(N,M)\in\operatorname{Mod}_{\operatorname{Sp}}(S^{\operatorname{op}})\times\operatorname{Mod}_{\operatorname{Sp}}(S)$ such that $N\otimes_{R}M\cong N\otimes_{S}M$. Observe that $C$ is closed under colimits and desuspensions in each variable.
Note that if $N,M$ are connective, then the pair $(N,M)$ is contained in $C$ if $\tau_{\leq n}N,\tau_{\leq n}M$ is contained in $C$ for all $n\in\mathbb{N}$. This is because $\lim_{n}\tau_{\leq n}N\otimes_{R}\tau_{\leq n}M\cong N\otimes_{R}M$, and similarly for $S$.
Thus to show that $f$ is a hom. epi., i.e the pair $(S,S)$ is in $C$, it suffices to show that $(\tau_{\leq n}S,\tau_{\leq n}S)$ is for each $n\in\mathbb{N}$. However, $\tau_{\leq n}S$ is bounded in the the $t$-structure on $\operatorname{Mod}_{\operatorname{Sp}}(S)$ and $\operatorname{Mod}_{\operatorname{Sp}}(S^{\operatorname{op}})$, and $\pi_{0}S$ generates the hearts of these categories, so the result follows since we know by assumption that $(\pi_{0}S,\pi_{0}S)\in C$. ∎
The following lemma allows us to commute quotients by $\mathbb{E}_{2}$-central elements with homological epimorphisms.
Lemma 3.4.0.
If $f:R\to S$ is a hom. epi, and $x:R\to R$ is a bimodule map, then there is a natural isomorphism of $R$-modules $\operatorname{cof}(x)\otimes_{R}(S\otimes_{R}M)\cong S\otimes_{R}\operatorname{cof}(x)\otimes_{R}M$
Proof.
Consider the full subcategory of $R$-bimodules generated under colimits by $R$. If $N$ is such a bimodule, then the functor $M\otimes_{R}(-)$ preserves the subcategory of $R$-modules that are $S$-modules, because it is true for the $R$-bimodule $R$ itself, and the subcategory is closed under colimits. Thus there is a natural transformation from $S\otimes_{R}N\otimes_{R}M\to N\otimes_{R}(S\otimes_{R}M)$ defined when $N$ is in this subcategory. To see that this map is an equivalence, we observe it is an equivalence for $N=R$, and the condition is closed under colimits. ∎
3.5. Nilpotence of actions
Here we record a few well known lemmas we use to understand the nilpotence of actions of $\mathbb{E}_{2}$-central elements:
Lemma 3.5.0.
Let ${\mathcal{C}}$ be a stable $\infty$-category, and consider a diagram as below:
[…]
Then the composite of the associated maps on vertical cofibers $\operatorname{cof}(f)\to\operatorname{cof}(f^{\prime})\to\operatorname{cof}(f^{\prime\prime})$ is nullhomotopic.
In particular, if a natural endomorphism $g$ of the identity functor acts by 00 on $X$ and $Y$, then $g^{2}$ acts by 00 on $\operatorname{cof}(f)$ for any map $X\to Y$.
Proof.
For the first claim, the composite $\operatorname{cof}(f)\to\operatorname{cof}(f^{\prime})\to\Sigma X^{\prime}$ is null since $X\to X^{\prime}$ is null, so we obtain a factorization of $\operatorname{cof}(f)\to\operatorname{cof}(f^{\prime})$ through $Y^{\prime}$. Composing with the second horizontal map, since $Y^{\prime}\to Y^{\prime\prime}$ is 00, we obtain a nullhomotopy of the desired map.
For the second claim, we apply the first claim when $X=X^{\prime}=X^{\prime\prime},Y=Y^{\prime}=Y^{\prime\prime}$ and $f=f^{\prime}=f^{\prime\prime}=g$. ∎
For the next lemma, we say that a monoidal stable $\infty$-category is exactly monoidal if the tensor product is exact in each variable.
Lemma 3.5.0.
Let ${\mathcal{C}}$ be an exactly monoidal stable $\infty$-category, let $f:\operatorname{\mathbb{1}}\to\operatorname{\mathbb{1}}$ be a map in ${\mathcal{C}}$, where $\operatorname{\mathbb{1}}$ is the monoidal unit. Then the map $\operatorname{cof}(f)\otimes f^{\otimes 2}$ is null.
In particular, if $R$ is an $\mathbb{E}_{1}$-algebra, $f:R\to\Sigma^{i}R$ a bimodule map, and $M\in\operatorname{Mod}_{\operatorname{Sp}}(R)$, then $f^{2}$ acts by 00 on $\operatorname{cof}f\otimes_{R}M$ as a map of $R$-modules.
Proof.
By identifying the category of $R$-bimodules with the category of colimit preserving endofunctors of $\operatorname{Mod}_{\operatorname{Sp}}(R)$, which is a monoidal stable $\infty$-category, the second statement follows from the first.
[…]
4. Computing stable homology
The goal of this section is to compute the stable homology of the Hurwitz spaces of interest. First, in § 4.1, we give some background on racks, which is the natural setting for our main stability result. Next, in § 4.2 we state the main result, Theorem 4.2.1, and explain how this implies our main results on stable homology. Then, in § 4.3 we prove that multiplication by certain components of Hurwitz spaces act centrally on homology. In § 4.4 we reduce the proof of Theorem 4.2.1 to showing that a certain map of pointed bar constructions is a homotopy equivalence. We prove this equivalence in § 4.5, relying on a topological model for the relevant spaces provided by Appendix A.
4.1. Background on racks
Although we will ultimately only be interested in working with Hurwitz spaces where the monodromy lives in a certain conjugacy class in a group, all of the arguments go through just as easily when working with racks. Hence, for this section, we will work with racks. One concise reference for the basics of racks is [EG03].
Definition 4.1.0.
A rack is a set $c$ with a binary operation $\triangleright:c\times c\to c$ such that the following two properties hold:
  1. (1)
    For every $x\in c$, the map $\phi_{x}:c\to c,y\mapsto x\triangleright y$ is a bijection.
  2. (2)
    For any $x,y,z\in c$, $x\triangleright(y\triangleright z)=(x\triangleright y)\triangleright(x\triangleright z)$.
Example 4.1.0.
The only type of rack we will need for our application to the Cohen–Lenstra heuristics and to prove Theorem 1.4.5 is a rack which is a conjugacy class $c$ in a group $G$, in which case we take $x\triangleright y:=x^{-1}yx$, where, on the right hand side, $x^{-1}yx$ is viewed as multiplication in the group.
We now introduce some further notation we will use for racks.
Notation 4.1.0.
For $c$ a rack and $g,g_{1},\ldots,g_{n}\in c$, we will use the notation
$(g_{1}\cdots g_{n})\triangleright g:=(g_{n}\triangleright(\cdots(g_{2}\triangleright(g_{1}\triangleright g))\cdots)).$
Note the reversal of order in the iteration of the $g_{i}$, which is compatible with the convention from § 4.1.
Notation 4.1.0.
For $c$ a rack, we use $\operatorname{Hur}^{c}_{n}$ to denote the pointed Hurwitz space of degree $n$ for $c$, which is by definition the homotopy quotient $(c^{n})_{hB_{n}}$ of $c^{n}$ by the action of $B_{n}$, where the positive generator of $B_{2}\simeq\mathbb{Z}$ acts via $(g,h)\mapsto(h,h\triangleright g)$. We use $\operatorname{Hur}^{c}$ to refer to the union of the Hurwitz spaces $\coprod_{n}\operatorname{Hur}_{n}^{c}$, and let $\operatorname{Conf}$ refer to the union of the configuration spaces $\coprod_{n}\operatorname{Conf}_{n}$, for $\operatorname{Conf}_{n}$ denoting the configuration space parameterizing unordered sets of $n$ distinct points in $\mathbb{C}$.
Notation 4.1.0.
We can identify the connected components of the pointed Hurwitz space $\operatorname{Hur}^{c}_{n}$ with orbits of the $B_{n}$ action on $c^{n}$. Under this identification, for $x_{1},\ldots,x_{n}\in c$ we use $[x_{1}]\cdots[x_{n}]$ to denote the connected component of $\operatorname{Hur}^{G,c}_{n,\mathbb{C}}$ corresponding to the $B_{n}$ orbit of the tuple $(x_{1},\ldots,x_{n})$.
The following definition is a generalization of the non-splitting condition to the setting of racks.
Definition 4.1.0.
We say a rack $c$ is connected if there is a single orbit under the operations $x\triangleright(-)$ for each $x\in c$. We say that a rack $c$ is non-splitting if every nonempty subrack $c^{\prime}\subset c$ is connected.
Remark 4.1.0.
It follows from the definitions that if $c$ comes from a conjugacy class in a group $G$ in the sense of § 4.1, then $(G,c)$ is non-splitting in the sense of § 1.4.4 if $c$ is a non-splitting rack.
Our main result computes the stable homology of Hurwitz spaces associated to non-splitting racks $c$. See § 5.1 for some examples of non-splitting racks $c$ relevant for the Cohen–Lenstra heuristics.
4.2. The main stability result
We now set up some notation we will use in this section and state our main result. The map $\operatorname{Hur}^{c}\to\operatorname{Conf}$ is a map in $\operatorname{Alg}(\operatorname{Spc})$. To do this, we introduce a suitable comparison map:
Notation 4.2.0.
For a rack $c$, let $\operatorname{Conf}^{c}$ be the pullback $\operatorname{Conf}\times_{\pi_{0}\operatorname{Conf}}\pi_{0}\operatorname{Hur}^{c}$, so that there is a map $\operatorname{Hur}^{c}\to\operatorname{Conf}^{c}$ in $\operatorname{Alg}(\operatorname{Spc})$. Let $A:=C_{*}(\operatorname{Hur}^{c};\mathbb{Q})$ and $A^{\prime}:=C_{*}(\operatorname{Conf}^{c};\mathbb{Q})$. Applying $C_{*}(-;\mathbb{Q})$ to the map $\operatorname{Hur}^{c}\to\operatorname{Conf}^{c}$, we obtain a map $v:A\to A^{\prime}$ of $\mathbb{E}_{1}$-algebras in spectra, i.e in the $\infty$-category $\operatorname{Alg}(\operatorname{Sp})$.
The next result implies our main theorem on stable homology. After stating it, we will explain why it implies Theorem 1.4.5 from the introduction.
Theorem 4.2.1.
With notation as in § 4.2, suppose $c$ is a finite non-splitting rack. Let $U\in\pi_{0}A$ be $\sum\alpha_{i}$, where each $\alpha_{i}$ comes from a class in $\pi_{0}\operatorname{Hur}_{n}^{c}$ for $n>0$ which is central. Then, the map $v[U^{-1}]:A[U^{-1}]\to A^{\prime}[U^{-1}]$ is an equivalence.
We prove this in § 4.5.1. From this, we easily deduce Theorem 1.4.5 from the introduction, computing the stable homology of our Hurwitz spaces.
4.2.2. Proof of Theorem 1.4.5 assuming Theorem 4.2.1
First, note that there is an identification between the pointed Hurwitz space $\operatorname{Hur}_{n}^{G,c}$ and $\operatorname{Hur}_{n}^{c}$ as both can be described as the homotopy quotient $(c^{n})_{hB_{n}}$, see § 1.4.2. Let us fix a value of $i\geq 0$. By [EVW16, Theorem 6.1], there exist constants $I,J$, and $D$ [^4] In fact, one may take $D=1$ by [EL23, Proposition A.3.1]. so that for any $i\geq 0$ and $n>Ii+J$ and $U:=\sum_{h\in c}\alpha_{h}^{D\operatorname{ord}(h)}$, $\pi_{i}A_{n}$ agrees with $\pi_{i}A[U^{-1}]_{n}$.
Because of the definition of $\operatorname{Conf}^{c}$, it follows that $\pi_{i}A^{\prime}_{n}$ also agrees with $\pi_{i}A^{\prime}[U^{-1}]_{n}$ as soon as $\pi_{0}A$ has stabilized. Thus it follows from Theorem 4.2.1 that for any $i\geq 0$ and $n>Ii+J$, the map $\pi_{i}A_{n}\to\pi_{i}A^{\prime}_{n}$ is an isomorphism, since it agrees with the map induced by $v[U^{-1}]$ on $\pi_{i}$. But this is exactly saying that $H_{i}(\operatorname{Hur}_{n}^{c};\mathbb{Q})\cong H_{i}(\operatorname{Conf}_{n}^{c};\mathbb{Q})$ for $n\gg 0$, which is a reformulation of the statement that for each component $Z$ of $\operatorname{Hur}_{n}^{c}$, we have $H_{i}(Z;\mathbb{Q})\cong H_{i}(\operatorname{Conf}_{n};\mathbb{Q})$, proving Theorem 1.4.5. ∎
4.3. Centrality
In this subsection, we let $c$ be an arbitrary rack. The goal of this subsection is to show that if $x\in\pi_{0}\operatorname{Hur}^{c}$ is central, then $x$ is actually central in a much stronger sense. Namely, it is $\mathbb{E}_{2}$-central in $\operatorname{Hur}^{c}$, meaning that multiplication by $x$ can be refined to a $\operatorname{Hur}^{c}$-bimodule map.
Informally, in the case of Hurwitz spaces associated to groups, the reason for this stronger form of centrality is that there is a homotopy $[\alpha][\beta]\simeq[\beta][\beta^{-1}\alpha\beta]$ obtained by changing the direction from which one concatenates $G$-covers. Thus if $\beta$ centralizes $c$, this identifies left and right multiplication by $\beta$. The above argument is not enough to make multiplication by $\beta$ into a bimodule map, but it can be made as such. We now argue this below and further generalize from groups to racks:
Lemma 4.3.0.
Using notation as in § 3.3, we let $c$ be a rack. Let $\gamma\in\pi_{0}\operatorname{Hur}^{c}$ be central. If we view $\operatorname{Hur}^{c}$ as an $\mathbb{E}_{1}$-algebra in the $\infty$-category $\operatorname{Spc}_{/\operatorname{Conf}}$, then multiplication by $\gamma$ in $\operatorname{Hur}^{c}$ can be refined to the structure of an $\operatorname{Hur}^{c}$-bimodule map.
Proof.
We can consider the $\mathbb{E}_{2}$-monoidal $\infty$-category of local systems on $\operatorname{Conf}$, $\operatorname{Spc}^{\operatorname{Conf}}$. This can be identified with the $\infty$-category of tuples $(X_{i})_{i\in\mathbb{N}}$, where $X_{i}\in\operatorname{Spc}^{BB_{i}}$ is a space with an action of the braid group on $i$-letters. Indeed this follows because $\operatorname{Conf}$ is the free $\mathbb{E}_{2}$-algebras in spaces on a point, and is $1$-truncated, and $\coprod_{n}BB_{n}$ is the free $\mathbb{E}_{2}$-monoidal category on an object [JS86, Theorem 4], and is a groupoid, so they agree.
We can view $\operatorname{Hur}^{c}$ as the colimit of the multiplicative local system given by the tuple $(c^{i})_{i\in\mathbb{N}}$, where the action of the braid group is determined by the fact that the twist sends a pair $(x,y)\in c\times c$ to $(y,y\triangleright x)$, and the multiplication is given by concatenation. Note that since this is a local system of sets, we may view $\operatorname{Hur}^{c}$ as an object in the (ordinary) braided monoidal category $\operatorname{Fun}(\coprod_{n}BB_{n},\operatorname{Set})$. Here the tensor product of $(X_{i})_{i\in\mathbb{N}},(Y_{i})_{i\in\mathbb{N}}$ is given by $(\coprod_{i+j=n}\operatorname{Ind}^{B_{n}}_{B_{i}\times B_{j}}(X_{i}\times Y_{j}))_{n\in\mathbb{N}}$, and the braiding is determined by the formula given in [JS86, p7].
Let $m$ be such that $\gamma\in\operatorname{Hur}_{m}^{c}$. Consider $m_{!}(*)$, where $m$ is considered as the inclusion of the base point $*\to BB_{m}$. Explicitly, $m_{!}(*)$ is the local system which is $B_{m}$ in degree $m$ and empty otherwise. There is a left module map $(c^{i})_{i\in\mathbb{N}}\otimes m_{!}(*)\to(c^{i})_{i\in\mathbb{N}}$, given by right multiplying by the class $\gamma$ on the identity element of $B_{m}$. Here we realize $(c^{i})_{i\in\mathbb{N}}\otimes m_{!}(*)$ as a bimodule via the $\mathbb{E}_{2}$-monoidal structure, i.e using the twist map $(c^{i})_{i\in\mathbb{N}}\otimes m_{!}(*)\to m_{!}(*)\otimes(c^{i})_{i\in\mathbb{N}}$. The fact that this multiplication map is a bimodule map follows if the diagram
${(c^{i})_{i\in\mathbb{N}}\otimes m_{!}(*)}$${(c^{i})_{i\in\mathbb{N}}}$${m_{!}(*)\otimes(c^{i})_{i\in\mathbb{N}}}$$\scriptstyle{\bullet_{r}}$$\scriptstyle{\operatorname{tw}}$$\scriptstyle{\bullet_{l}}$
commutes, where $\bullet_{l}$ is left multiplication, $\bullet_{r}$ is right multiplication, and $\operatorname{tw}$ is the twist map. Unraveling the definitions, this holds because the twist sends a tuple $(x,\gamma),x\in c^{i}$ to $(\gamma,x)$, which uses that $\gamma$ is central in $\pi_{0}\operatorname{Hur}^{c}$, so that $(\gamma\triangleright)$ acts trivially on $c$.
Thus we have constructed a bimodule map of multiplicative local systems on $\operatorname{Conf}$ induced by multiplication by $\gamma$. By applying the $\mathbb{E}_{2}$-monoidal functor $\operatorname{colim}:\operatorname{Spc}^{\operatorname{Conf}}\to\operatorname{Spc}_{/\operatorname{Conf}}$, (where the map to $\operatorname{Conf}$ is obtained via the colimit of the terminal local system) we see that this bimodule map refines the multiplication by $\gamma$ map. ∎
Corollary 4.3.0.
Let $\gamma\in\pi_{0}\operatorname{Hur}^{c}$ be central. Then the associated class $[\gamma]\in\pi_{0}A$ is $\mathbb{E}_{2}$-central.
Proof.
This follows by applying the functor $C_{*}(-;\mathbb{Q})$ to the bimodule map produced in § 4.3. ∎
4.4. Decomposition of the comparison map
To check that the comparison map $v$ is an equivalence, we geometrically decompose the algebra $A$ based on central elements acting on it. We accomplish this decomposition in § 4.4, which roughly says it suffices to check $v[U^{-1}]$ is an equivalence after suitable localizations and completions.
As a first step toward this, we note that certain completions of localization of $A$ are well behaved in the sense of § 3.3:
Notation 4.4.0.
[…]
The following lemma is a generalization of [EVW16, Proposition 3.4] from groups to racks, whose proof requires no new ideas.
Lemma 4.4.0.
Let $c$ be a finite connected rack, and let $g\in c$. For every $N$ sufficiently large, any tuple of elements $(g_{1},\ldots,g_{N})\in c^{N}$ generating the rack $c$ is equivalent under the braid group action to some tuple of the form $(g,g_{1}^{\prime},\ldots,g_{N-1}^{\prime})$ where $g_{1}^{\prime},\ldots,g_{N-1}^{\prime}$ generates $c$. Similarly, $(g_{1},\ldots,g_{N})$ is equivalent under the braid group action to some tuple of the form $(g_{1}^{\prime\prime},\ldots,g_{N-1}^{\prime\prime},g)$ where $g_{1}^{\prime\prime},\ldots,g_{N-1}^{\prime\prime}$ generates $c$.
Proof.
We summarize the proof, leaving details to the reader, which can also be found in [EVW16, Proposition 3.4] in the case that $c$ is a conjugacy class in a group. We will only prove the first statement regarding equivalence to $(g,g_{1}^{\prime},\ldots,g_{N-1}^{\prime})$. The second statement regarding equivalence to $(g_{1}^{\prime\prime},\ldots,g_{N-1}^{\prime\prime},g)$ has an analogous proof.
Choose $N$ large enough so that some $g^{\prime}\in c$ appears $n+1$ times, where $n$ is such that $\phi_{g^{\prime}}^{n}=\operatorname{id}$. Using the action of the braid group, we can move all of these copies of $g^{\prime}$ to the left to see that $(g_{1},\ldots,g_{N})$ is equivalent under the braid group action to $(\underbrace{g^{\prime},\ldots,g^{\prime}}_{n+1},h_{1},\ldots,h_{N-n-1})$, where $g^{\prime},h_{1},\ldots,h_{N-n-1}$ still generate the rack. Because $\phi_{g^{\prime}}^{n}=\operatorname{id}$, by using the element of the braid group that does a full twist of the last element around the rest, we see that $(\underbrace{g^{\prime},\ldots,g^{\prime}}_{n},h)\in c^{n+1}$ is equivalent to $(\underbrace{h\triangleright g^{\prime},\ldots,h\triangleright g^{\prime}}_{n},h)\in c^{n+1}$, where there are $n$ copies of $g^{\prime}$. Using this along with the fact that $g^{\prime},h_{1},\ldots,h_{N-n-1}$ generate the rack and the rack is connected, shows that we can modify $(\underbrace{g^{\prime},\ldots,g^{\prime}}_{n+1},h_{1},\ldots,h_{N-n-1})$ by an element of the braid group to change the first $n$ copies of $g^{\prime}$ into $n$ copies of $g$, while keeping the fact that the rest of the elements generate the rack. ∎
Lemma 4.4.0.
With notation as in § 4.4, let $c$ be a finite nonempty non-splitting rack and let $c^{\prime}\subset c$ be a subrack. In $A[\alpha_{c^{\prime}}^{-1}]$, any element $\alpha_{\beta}$ with $\beta\in c-c^{\prime}$ satisfies the conditions of § 3.3. Thus for an $A$-module $M$, we can form $M[\alpha_{c^{\prime}}^{-1}]^{\wedge}_{\alpha_{c-c^{\prime}}}$ as in § 3.3.
Proof.
By § 4.3, $\alpha_{\beta}^{\operatorname{ord}(\beta)}$ is $\mathbb{E}_{2}$-central where $\operatorname{ord}(\beta)$ is the order of $\phi_{\beta}$ as an automorphism of $c$. It remains to prove that $y\alpha_{\beta}^{2}=\alpha_{\beta}$ for some $y\in\pi_{*}A[\alpha_{c^{\prime}}^{-1}]$. To do this, choose $\gamma\in c^{\prime}$. Note that because $c$ is non-splitting, we have $\langle c^{\prime},\beta\rangle$ is connected. Applying § 4.4 to the subrack $\langle c^{\prime},\beta\rangle\subset c$ generated by $c^{\prime}$ and $\beta$, , for $N$ sufficiently large, we can write $\alpha_{x_{1}}\cdots\alpha_{x_{N}}\alpha_{\beta}=\alpha_{\gamma}^{N}\alpha_{\beta}$, where $x_{1},\ldots,x_{N}\in c$ together generate the same subrack as $\gamma,\beta$. Applying § 4.4 again, we can write $\alpha_{y_{1}}\cdots\alpha_{y_{N-1}}\alpha_{\beta}=\alpha_{x_{1}}\cdots\alpha_{x_{N}}$. Taking $y:=\alpha_{\gamma}^{-N}\alpha_{y_{1}}\cdots\alpha_{y_{N-1}}$, we therefore have $y\alpha_{\beta}^{2}=\alpha_{\beta}$. This verifies the conditions of § 3.3. ∎
The following reduces one to considering stable homology classes on which $\alpha_{c^{\prime}}$, for a subrack $c^{\prime}\subset c$, act by isomorphisms, and the complementary set act nilpotently.
Lemma 4.4.0.
Let $c$ be a finite non-splitting rack. Using notation from § 4.4 and § 4.2, and letting $v,U$ be as in Theorem 4.2.1, the map $v[U^{-1}]$ is an equivalence if, for every nonempty subrack $c^{\prime}\subset c$, $(v[\alpha_{c^{\prime}}^{-1}])^{\wedge}_{\alpha_{c-c^{\prime}}}$ is an equivalence, where the completion is taken as an $A$-module.
Proof.
By repeatedly applying § 3.3, we learn that it is enough to show that for every subset $c^{\prime}\subset c$,
(4.1) $(A/\alpha^{\operatorname{ord}(\beta_{|c-c^{\prime}|})}_{\beta_{|c-c^{\prime}|}})\otimes_{A}\cdots\otimes_{A}(A/\alpha^{\operatorname{ord}(\beta_{1})}_{\beta_{1}})\otimes_{A}v[U^{-1}][\alpha_{c^{\prime}}^{-1}]$
is an equivalence, where $\coprod_{i=1}^{|c-c^{\prime}|}\beta_{i}=c-c^{\prime}$. Here, we have used § 3.4 to commute the localization with the quotienting, and § 4.3 to make sense of the iterated quotient. Namely, because $\alpha_{\beta_{i}}^{\operatorname{ord}(\beta_{i})}$ is a bimodule map by § 4.3, we can tensor with the $A-A$-bimodule $(A/\alpha^{\operatorname{ord}(\beta_{1})}_{\beta_{1}})$ to make sense of the tensor product above as an $A$-module.
Note that on the source and target of this map, for each $\beta\in c-c^{\prime}$, $\alpha_{\beta}^{\operatorname{ord}(\beta)n_{\beta}}$ acts by 0 for some $n_{\beta}$. This follows from applying § 3.5 to see that after quotienting out by $\alpha_{\beta}^{\operatorname{ord}(\beta)}$, $\alpha_{\beta}^{2\operatorname{ord}(\beta)}$ acts by 00, and using § 3.5 to see that after each further quotient, the smallest power of $\alpha_{\beta}^{\operatorname{ord}(\beta)}$ that acts by 00 gets multiplied by at most $2$.
If $c^{\prime}$ is empty, then we claim that both the source and target of the map (4.1) are 00. We claim that some power of $U$ acts by 00 on the source and target, since a large enough power of $U$, $U^{j}$, is in the two sided ideal of $\pi_{*}R$ generated by $\alpha_{\beta}^{\operatorname{ord}(\beta)n_{\beta}}$. This is because by the pigeonhole principle, one can see that any element of $\pi_{0}\operatorname{Hur}^{G,c}_{n}$, written as a product of elements in $c$ for $n$ large enough contains $\max_{\beta\in c}(\operatorname{ord}(\beta)n_{\beta})$ many copies of a single $\beta\in c$. By pulling these to the left using the braid group action, we see the element is divisible by $\alpha_{\beta}^{\operatorname{ord}_{\beta}n_{\beta}}$. Since $U^{j}$ is in the ideal generated by the smallest powers of $\alpha_{\beta}^{\operatorname{ord}(\beta)}$ that act trivially, and hence it acts trivially. Since $U$ acts both as an isomorphism and nilpotently, the source and target of (4.1) with $c^{\prime}=\emptyset$ are both 00.
We next claim that the source and target of (4.1) vanish unless $c^{\prime}\subset c$ is a subrack. Indeed, observe that $\alpha_{a}\alpha_{b}=\alpha_{b}\alpha_{b\triangleright a}$. Hence, we learn that if $\alpha_{a}$ and $\alpha_{b}$ act invertibly, then so does $\alpha_{b\triangleright a}$. Thus the source and target of (4.1) vanish unless $c^{\prime}$ is closed under the operation $\triangleright$, i.e., $c^{\prime}$ is a subrack.
Suppose the map
(4.2) $(A/\alpha^{\operatorname{ord}(\beta)}_{\beta_{|c-c^{\prime}|}})\otimes_{A}\cdots\otimes_{A}(A/\alpha^{\operatorname{ord}(\beta)}_{\beta_{1}})\otimes_{A}v[\alpha_{c^{\prime}}^{-1}]$
(which is similar to (4.1) but here we have not inverted $U$) is an equivalence for each subrack $c^{\prime}\subset c$. Using § 3.4 to commute the localization and quotienting, we find that if (4.2) is an equivalence for each subrack $c^{\prime}\subset c$, (4.1) will also be an equivalence. Hence, it is enough to show (4.2) is an equivalence for each subrack $c^{\prime}\subset c$. Finally, we note that (4.2) is an equivalence for a fixed $c^{\prime}\subset c$ if and only if $(v[\alpha_{c^{\prime}}^{-1}])^{\wedge}_{\alpha_{c-c^{\prime}}}$ is an equivalence by using § 4.4 and § 3.3. ∎
4.5. Showing the map is an equivalence
In this section we show that the maps in § 4.4 are an equivalence for each subrack $c^{\prime}\subset c$. In the case that $c^{\prime}=c$, the fact that the map in § 4.4 is an equivalence comes from the fact that the homology of the group completion of $\operatorname{Hur}^{c}$ can be identified with the rack homology of the rack $c$, which is computed in [EG03]. We verify this in § 4.5 For $c^{\prime}\subset c$. In general, we show that completion at $\alpha_{c-c^{\prime}}$ (see § 4.4) corresponds to projecting in pointed spaces to the components of $\operatorname{Hur}^{c^{\prime}}$ at the level of homology, allowing us to again apply § 4.5 for the rack $c^{\prime}$ in this case. Identifying the completion in terms of $\operatorname{Hur}^{c^{\prime}}$ is at the core of our proof, and is proven using a homological epimorphism argument whose main input is a topological one identifying two bar constructions in pointed spaces as homotopy equivalent.
Proposition 4.5.0.
Use notation from § 4.2 and § 4.4, let $c$ be a finite connected rack. The map $v[\alpha_{c}^{-1}]$ is an equivalence.
Remark 4.5.0.
§ 4.5 was essentially already known, and can be found in [RW20, Corollary 5.4]. Although this was stated for conjugacy classes of groups, the proof works equally well for racks. We now spell out the details of this argument for general racks.
Proof.
By the group completion theorem, $v[\alpha_{c}^{-1}]$ can be identified with the map $\Omega B\operatorname{Hur}^{c}\to\Omega B\operatorname{Conf}^{c}$ on homology. It is then enough to show that the map $\Omega B\operatorname{Hur}^{c}\to\Omega B\operatorname{Conf}^{c}$ is a rational homology equivalence. Since the map $\pi_{1}B\operatorname{Hur}^{c}\to\pi_{1}B\operatorname{Conf}^{c}$ is an equivalence by construction of $\operatorname{Conf}^{c}$, by the Serre spectral sequence of the path loop fibration, it suffices to check $B\operatorname{Hur}^{c}\to B\operatorname{Conf}^{c}$ is a rational homology equivalence.
We will first show that the map
(4.3) $\psi:\Omega B\operatorname{Conf}^{c}\to\Omega B\operatorname{Conf}\times_{\Omega B\pi_{0}\operatorname{Conf}}\Omega B\pi_{0}\operatorname{Conf}^{c}$
is an equivalence. Note that $\pi_{1}(\psi):\pi_{1}(\Omega B\operatorname{Conf}^{c})\to\pi_{1}(\Omega B\operatorname{Conf}\times_{\Omega B\pi_{0}\operatorname{Conf}}\Omega B\pi_{0}\operatorname{Conf}^{c})$ is an equivalence. Therefore, in order to show $\psi$ is an equivalence, it suffices to show $\psi$ is an equivalence on homology. By the group completion theorem,
(4.4) $H_{i}(\Omega B\operatorname{Conf}^{c};\mathbb{Z})\cong H_{i}(\operatorname{Conf}^{c};\mathbb{Z})[\alpha_{c}^{-1}].$
Since $H_{i}(\operatorname{Conf}^{c}_{n};\mathbb{Z})\cong\oplus_{\pi_{0}\operatorname{Hur}^{c}_{n}}H_{i}(\operatorname{Conf}_{n};\mathbb{Z})$, taking the sum over $n$ and inverting $\alpha_{c}$, and using that $\pi_{0}\operatorname{Hur}^{c}[(\pi_{0}\operatorname{Hur}^{c})^{-1}]\cong\Omega B\pi_{0}\operatorname{Conf}^{c}$, we obtain
(4.5) $H_{i}(\operatorname{Conf}^{c};\mathbb{Z})[\alpha_{c}^{-1}]\cong\oplus_{\Omega B\pi_{0}\operatorname{Conf}^{c}}\operatorname{colim}_{n}H_{i}(\operatorname{Conf}_{n};\mathbb{Z}).$
Note here that the colimit in $n$ is given by the stabilization map $\alpha:\operatorname{Conf}_{n}\to\operatorname{Conf}_{n+1}$, since each $\alpha_{x}$ for $x\in c$ acts through the same map, which can be identified with $\alpha$ on each component of $\operatorname{Conf}^{c}$. Applying the group completion theorem again to $\operatorname{Conf}$, we learn that
(4.6) $H_{i}(\Omega B\operatorname{Conf})\cong H_{i}(\operatorname{Conf};\mathbb{Z})[\alpha^{-1}]\cong\oplus_{\Omega B\pi_{0}\operatorname{Conf}}\operatorname{colim}_{n}H_{i}(\operatorname{Conf}_{n};\mathbb{Z}).$
so combining (4.4),(4.5), and (4.6), we find
$H_{i}(\Omega B\operatorname{Conf}^{c};\mathbb{Z})\cong\oplus_{\Omega B\pi_{0}\operatorname{Conf}^{c}}\operatorname{colim}_{n}H_{i}(\operatorname{Conf}_{n};\mathbb{Z})\cong H_{i}(\Omega B\operatorname{Conf}\times_{\Omega B\pi_{0}\operatorname{Conf}}\Omega B\pi_{0}\operatorname{Conf}^{c};\mathbb{Z}).$
Hence, $\psi$ is an equivalence.
We next claim that $B\operatorname{Conf}^{c}\to B\operatorname{Conf}$ is an equivalence on rational homology. The map $\Omega B\pi_{0}\operatorname{Hur}^{c}\to\Omega B\pi_{0}\operatorname{Conf}$ is surjective with kernel a group we call $G$. Here, $G$ is finite because of the fact that $|\pi_{0}\operatorname{Conf}^{c}_{n}|=|\pi_{0}\operatorname{Hur}^{c}_{n}|$ stabilizes as $n$ grows, which is a consequence of § 4.4 and the fact that a filtered colimit of finite sets along surjections is finite. It follows from this and the fact that $\psi$ from (4.3) is an equivalence that the fiber of the map $B\operatorname{Conf}^{c}\to B\operatorname{Conf}$ is $BG$. We then obtain our claim that $B\operatorname{Conf}^{c}\to B\operatorname{Conf}$ is an equivalence on rational homology using the Serre spectral sequence for this fiber sequence, since the rational homology of $G$ is just $\mathbb{Q}$ in degree 00 with trivial action of $B\operatorname{Conf}$.
It remains then to show that the map $B\operatorname{Hur}^{c}\to B\operatorname{Conf}$ is an equivalence on rational homology. This follows as in [RW20, Corollary 5.4]: This map can be modeled as the map of rack spaces as in [FRS95], from the rack $c$ to the trivial rack *, and the induced map on homology is the map on rack homology induced from this map of racks. It is evident that the map on rack homology is a split surjection (since the associated map of braided vectors spaces has a splitting sending $1$ to $\frac{1}{|c|}\sum_{g\in c}g$), and on the other hand [EG03, Theorem 4.2] shows that the rack cohomology of both sides has rank $1$ in each degree (using that $c$ is connected). Thus the map is an isomorphism. ∎
Remark 4.5.0.
In fact, the proof above (as well as this whole section) doesn’t just work rationally, but also after inverting the order of $G$ showing up in the proof and the order of the ‘reduced structure group’ of the rack $c$ [EG03, Section 2]. In the case $c$ comes from a conjugacy class in a group $\langle c\rangle$, all primes dividing $|G|$ appearing in the proof (which is potentially different from $\langle c\rangle$ and all primes dividing the reduced structure group of the rack are contained in the set of primes dividing $\langle c\rangle$, by [Woo21, Theorem 2.5]. So, we can compute stable homology after one inverts the order of $\langle c\rangle$.
We continue to use notation from § 4.1, § 4.2, and § 4.4.
Notation 4.5.0.
For $c^{\prime}\subset c$ a subrack, there is a map of $\mathbb{E}_{1}$-algebras $\overline{i}_{c^{\prime}}^{c}:\operatorname{Conf}^{c^{\prime}}\to\operatorname{Conf}^{c}$ that is an inclusion of components. Similarly, there is a map $\tilde{i}_{c^{\prime}}^{c}:\operatorname{Hur}^{c^{\prime}}\to\operatorname{Hur}^{c}$. We have a section of the induced map $\operatorname{Conf}^{c^{\prime}}_{+}\to\operatorname{Conf}^{c}_{+}$ of pointed $\mathbb{E}_{1}$-algebras $\overline{r}_{c}^{c^{\prime}}:\operatorname{Conf}^{c}_{+}\to\operatorname{Conf}^{c^{\prime}}_{+}$ obtained by sending all components not in image of $\operatorname{Conf}^{c}$ to the base point. Similarly there is a map $\widetilde{r}_{c}^{c^{\prime}}:\operatorname{Hur}^{c}_{+}\to\operatorname{Hur}^{c^{\prime}}_{+}$ obtained the same way. Let $r_{c}^{c^{\prime}}$ denote the associated $\mathbb{E}_{1}$-algebra map $A\to C_{*}(\operatorname{Hur}^{c^{\prime}})$ obtained by taking reduced chains. In particular, this map of algebras sends all generators $\alpha_{x}$, for $x\in c-c^{\prime}$, to 00.
§ 4.5 is the key to proving the case when $c^{\prime}$ is a proper subrack of $c$. To state it, we need to introduce the following condition on the subrack:
Definition 4.5.0.
For $c$ a rack and $c^{\prime}\subset c$ a subset, we define the normalizer of $c$ in $c^{\prime}$ to be $N(c^{\prime}):=\{x\in c|x\triangleright c^{\prime}=c^{\prime}\}$. We say that a subrack $c^{\prime}\subset c$ is self-normalizing if $c^{\prime}=N(c^{\prime})$.
Proposition 4.5.0.
Let $c$ be a rack. With notation as in § 4.5, let $c^{\prime}$ be a self-normalizing subrack of $c$. Then the map $r_{c}^{c^{\prime}}$ induces an equivalence $A[\alpha_{c^{\prime}}^{-1}]^{\wedge}_{\alpha_{c-c^{\prime}}}\to C_{*}(\operatorname{Hur}^{c^{\prime}};\mathbb{Q})[\alpha_{c^{\prime}}^{-1}]$.
Proof.
We first note that since $\alpha_{c-c^{\prime}}$ acts by 00 on $C_{*}(\operatorname{Hur}^{c^{\prime}};\mathbb{Q})[\alpha_{c^{\prime}}^{-1}]$, completion doesn’t change it, so the map $A[\alpha_{c^{\prime}}^{-1}]^{\wedge}_{\alpha_{c-c^{\prime}}}\to C_{*}(\operatorname{Hur}^{c^{\prime}};\mathbb{Q})[\alpha_{c^{\prime}}^{-1}]$ in the proposition statement is obtained by taking $r_{c}^{c^{\prime}}$, localizing at $[\alpha_{c^{\prime}}^{-1}]$ and completing at $\alpha_{c-c^{\prime}}$.
We will prove the proposition by applying § 3.4 (rigidity of hom. epis) to the triangle
${A}$${C_{*}(\operatorname{Hur}^{c^{\prime}};\mathbb{Q})[\alpha_{c^{\prime}}^{-1}]}$${A[\alpha_{c^{\prime}}^{-1}]^{\wedge}_{\alpha_{c-c^{\prime}}}.}$$\scriptstyle{h}$
The diagram is a triangle of $\mathbb{E}_{1}$-algebras since the completion on the source is obtained by inverting a central idempotent by § 4.4.
Note that all of the objects in the triangle are connective. By § 3.3, $\pi_{0}A[\alpha_{c^{\prime}}^{-1}]^{\wedge}_{\alpha_{c-c^{\prime}}}$ is the quotient of $\pi_{0}A[\alpha_{c^{\prime}}^{-1}]^{\wedge}_{\alpha_{c-c^{\prime}}}$ by the two sided ideal generated by $\alpha_{c-c^{\prime}}$. This ideal is exactly the classes coming from components not in $\operatorname{Hur}^{c^{\prime}}$. Thus the map $h$ is an equivalence on $\pi_{0}$. Moreover, $A\to A[\alpha_{c^{\prime}}^{-1}]^{\wedge}_{\alpha_{c-c^{\prime}}}$ is a hom. epi. by § 4.4 and § 3.3.
It remains to see that the map $A\to C_{*}(\operatorname{Hur}^{c^{\prime}};\mathbb{Q})[\alpha_{c^{\prime}}^{-1}]$ is a hom. epi. Let $X=\pi_{0}\operatorname{Hur}^{c^{\prime}}[\alpha_{c^{\prime}}^{-1}]$, so that $C_{*}(X;\mathbb{Q})\cong\pi_{0}(C_{*}(\operatorname{Hur}^{c^{\prime}};\mathbb{Q})[\alpha_{c^{\prime}}^{-1}])$. By § 3.4, it is enough to check that the map
$C_{*}(X;\mathbb{Q})\otimes_{A}C_{*}(X;\mathbb{Q})\rightarrow C_{*}(X;\mathbb{Q})\otimes_{C_{*}(\operatorname{Hur}^{c^{\prime}};\mathbb{Q})[\alpha_{c^{\prime}}^{-1}]}C_{*}(X;\mathbb{Q})$
is an equivalence. Since $C_{*}(\operatorname{Hur}^{c^{\prime}};\mathbb{Q})\to C_{*}(\operatorname{Hur}^{c^{\prime}};\mathbb{Q})[\alpha_{c^{\prime}}^{-1}]$ is a hom. epi, we may replace the target with $C_{*}(X;\mathbb{Q})\otimes_{C_{*}(\operatorname{Hur}^{c^{\prime}};\mathbb{Q})}C_{*}(X;\mathbb{Q})$. Then this comparison is now induced from taking reduced chains of a map at the level of spaces. In what follows, the tensor product is the two sided bar construction of pointed spaces and we use notation from § 3.1.1:
(4.7) $X_{+}\otimes_{\operatorname{Hur}^{c}_{+}}X_{+}\to X_{+}\otimes_{\operatorname{Hur}^{c^{\prime}}_{+}}X_{+}.$
So, it is enough to verify the map (4.7) is a homology equivalence, which follows from § 4.5 below, using that $c^{\prime}$ is self-normalization so $N(c^{\prime})=c^{\prime}$. ∎
Figure 1. This is a pictorial description of the homotopy from § 4.5. We take $b\in c^{\prime}$ but $g\notin c^{\prime}$. We pass the element $\xi(g)$ from the left to the right. When it meets the right boundary, it acts on the boundary by $b\triangleright(g\triangleright\xi(g)))$. By assumption $g\triangleright\xi(g)\notin c^{\prime}$. Since $b\in c^{\prime}$, $b\triangleright(g\triangleright\xi(g)))$ is also not in $c^{\prime}$. Therefore, the right boundary becomes the base point, and this whole configuration is then identified with the base point.
The final hurdle before proving our main result is the following key result, which was needed for the proof of § 4.5.
Proposition 4.5.0.
Let $c$ be a rack, $c^{\prime}\subset c$ be a subrack with normalizer $N(c^{\prime})$, and $X_{+}$ be $\pi_{0}\operatorname{Hur}^{c^{\prime}}[\alpha_{c^{\prime}}^{-1}]_{+}$, viewed as a $\operatorname{Hur}^{c^{\prime}}_{+}$-bimodule. Then the map
(4.8) $X_{+}\otimes_{\operatorname{Hur}^{c}_{+}}X_{+}\to X_{+}\otimes_{\operatorname{Hur}^{N(c^{\prime})}_{+}}X_{+}$
is a homology equivalence.
Remark 4.5.0.
In fact, one can show that the map (4.8) is not only a homology equivalence but even an equivalence, by verifying that the fundamental groupoids of the two spaces are isomorphic. However, we will not need this stronger statement, so we do not prove it.
Remark 4.5.0.
We note that the idea to define the filtration $F_{\bullet}$ in the following proof was communicated to us by Andrea Bianchi. This allowed us to remove an additional condition we had previously imposed on the subrack $c^{\prime}$, and thereby simplify the following proof. We would like to thank Andrea Bianchi for this very helpful observation.
Proof.
The map of interest $X_{+}\otimes_{\operatorname{Hur}^{c}_{+}}X_{+}\to X_{+}\otimes_{\operatorname{Hur}^{N(c^{\prime})}_{+}}X_{+}$ has a section induced from the inclusion of racks $N(c^{\prime})\to c$. It thus suffices to show that this section induces a homology equivalence.
We use the notation $\overline{Q}^{*}_{\epsilon}[M,\operatorname{Hur}_{+}^{N(c^{\prime})},N]$, as defined in § A.4 and Theorem A.4.1, which we describe below. By Theorem A.4.1, in order to prove this section is a homology equivalence it suffices to show the inclusion
$\iota_{\epsilon}:\overline{Q}^{*}_{\epsilon}[X_{+},\operatorname{Hur}_{+}^{N(c^{\prime})},X_{+}]\to\overline{Q}^{*}_{\epsilon}[X_{+},\operatorname{Hur}_{+}^{c},X_{+}]$
is an ind-weak homology equivalence (see § A.4) as $\epsilon$ approaches 00 where $10<\epsilon<$ 1.
We now recall the description of these spaces, which follows from § A.4. First, ${Q}_{\epsilon}[X_{+},\operatorname{Hur}^{c}_{+},X_{+}]$ is the space whose points consist of tuples $(m,n,x,l:x\to c)$ with the following properties: $m\in X_{+},n\in X_{+}$, $x$ is a finite subset $x$ of $[0,1]\times[\epsilon,1-\epsilon]$ such that if $\pi_{2}$ is the projection onto the second coordinate, then $\pi_{2}$ is injective on $x$, and the points in $\pi_{2}(x)$ are distance at least $\epsilon$ from each other, and $l:x\to c$ is a map of sets, which we colloquially refer to as a labeling function.
Viewing $X_{+}$ as having a left and right action of $\operatorname{Hur}^{c}$, $\overline{Q}^{*}_{\epsilon}[X_{+},\operatorname{Hur}^{c}_{+},X_{+}]$ is the quotient of ${Q}_{\epsilon}[X_{+},\operatorname{Hur}^{c}_{+},X_{+}]$ by the following three relations:
  • $(i)$
    Suppose $x=\{p_{1},\ldots,p_{r},\ldots,p_{k}\}$ with $\pi_{2}(p_{1})>\cdots>\pi_{2}(p_{k})$. If $\pi_{1}(p_{r})=0$, so $p_{r}$ lies on the left boundary, the data ($m,n,x,l:x\to c)$ is equivalent to $(m^{\prime},n,x^{\prime}:=\{p_{1},\ldots,p_{r-1},p_{r+1},\ldots,p_{k}\},l^{\prime}:x^{\prime}\to c)$, where $m^{\prime}=ml(p_{r})$, $l^{\prime}(p_{i}):=l(p_{r})\triangleright l(p_{i})$ for $1\leq i<r$, $l^{\prime}(p_{i}):=l(p_{i})$ for $i>r$.
  • $(ii)$
    Suppose $x=\{p_{1},\ldots,p_{r},\ldots,p_{k}\}$ with $\pi_{2}(p_{1})>\cdots>\pi_{2}(p_{k})$. If $\pi_{1}(p_{r})=1$, so $p_{r}$ lies on the right boundary, the data ($m,n,x,l:x\to c)$ is equivalent to $(m,n^{\prime},x^{\prime}:=\{p_{1},\ldots,p_{r-1},p_{r+1},\ldots,p_{k}\},l^{\prime}:x^{\prime}\to c)$, where $n^{\prime}=((l(p_{r+1})\cdots l(p_{k}))\triangleright l(p_{r}))n$, $l^{\prime}(p_{i}):=l(p_{i})$ for $i\neq r$.
  • $(iii)$
    All points where either $n$ or $m$ are the base point are identified to the base point.
The space $\overline{Q}^{*}_{\epsilon}[X_{+},\operatorname{Hur}^{N(c^{\prime})}_{+},X_{+}]$ is defined similarly, except the the labelling function $l$ must land in $N(c^{\prime})\subset c$.
At this point, it may be useful to refer to the notion of two spaces being ind-weakly equivalent, defined in § A.4. Let $S^{c,c^{\prime}}_{\epsilon}$ denote the quotient of the inclusion $\overline{Q}^{*}_{\epsilon}[X_{+},\operatorname{Hur}^{N(c^{\prime})}_{+},X_{+}]\to\overline{Q}^{*}_{\epsilon}[X_{+},\operatorname{Hur}^{c}_{+},X_{+}]$.
To show that the $\iota_{\epsilon}$ is an ind-weak homology equivalence, using that the inclusion has the homotopy extension property, (§ A.4), it suffices to show the following:
  • $\star^{\prime}$
    The space $S^{c,c^{\prime}}_{\epsilon}$ is ind-weakly homology equivalent to a point.
Note that any point in $S^{c,c^{\prime}}_{\epsilon}$ that is not the base point can be represented by a point in $\overline{Q}^{*}_{\epsilon}[X_{+},\operatorname{Hur}^{c}_{+},X_{+}]$ where some $l(p_{i})$ is not in $N(c^{\prime})$.
It will be easier to prove the desired homology equivalence at the level of the associated graded of a certain filtration, which we define next. We define the filtration $F_{\bullet}S^{c,c^{\prime}}_{\epsilon}$ on $S^{c,c^{\prime}}_{\epsilon}$, where, for $j\geq 0$, $F_{j}S^{c,c^{\prime}}_{\epsilon}$ is the subset of $S^{c,c^{\prime}}_{\epsilon}$ consisting of those points where one of $l(p_{k-j}),\ldots,l(p_{k})$ lies in $N(c^{\prime})-c$, together with the base point. We set $F_{-1}S^{c,c^{\prime}}_{\epsilon}$ to be the basepoint, and use $G_{j}S^{c,c^{\prime}}_{\epsilon}$ to denote the associated graded of the filtration $F_{j}S^{c,c^{\prime}}_{\epsilon}/F_{j-1}S^{c,c^{\prime}}_{\epsilon}$, obtained by identifying $F_{j-1}S^{c,c^{\prime}}_{\epsilon}$ with the base point.
It is easy to see that this is an exhaustive filtration by sub CW-complexes, so that it suffices to prove the following associated graded version of $\star^{\prime}$:
  • $\star$
    For each $j\geq 0$, $G_{j}S^{c,c^{\prime}}_{\epsilon}$ is ind-weakly equivalent to a point.
Fixing $j\geq 0$ and $\epsilon>0,$ we will find a smaller $\epsilon^{\prime}$ such that the map $G_{j}S^{c,c^{\prime}}_{\epsilon}\to G_{j}S^{c,c^{\prime}}_{\epsilon^{\prime}}$ is nullhomotopic.
We may choose a function $\xi:c-N(c^{\prime})\to c^{\prime}$ such that $g\triangleright\xi(g)\notin c^{\prime}$ for each $g\in c-N(c^{\prime})$, since $g\triangleright c^{\prime}\neq c^{\prime}$ from the definition of the normalizer of $c^{\prime}$.
We use ${Q}_{\epsilon}^{c,c^{\prime}}$ as notation to shorten ${Q}_{\epsilon}[X_{+},\operatorname{Hur}^{c}_{+},X_{+}]$. Let $\theta:{Q}_{\epsilon}^{c,c^{\prime}}\to S_{\epsilon}^{c,c^{\prime}}$ be the projection map. We define a filtration on $Q^{c,c^{\prime}}_{\epsilon}$ by $F_{\bullet}Q^{c,c^{\prime}}_{\epsilon}:=\theta^{-1}(F_{\bullet}S^{c,c^{\prime}}_{\epsilon})$ and use $G_{\bullet}Q^{c,c^{\prime}}_{\epsilon}$ to denote the associated graded.
Fix a point $y\in Q^{c,c^{\prime}}_{\epsilon}$ given by the data
  • ()($\dagger$)( † )
    $(m,n,x=\{p_{1},\ldots,p_{k}\},l:x\to c)$ with $\pi_{2}(p_{1})>\cdots>\pi_{2}(p_{k})$.
Suppose $y$ as above lies in $F_{j}Q_{\epsilon}^{c,c^{\prime}}$. Define $v$ to be the largest number so that $l(p_{v})\notin N(c^{\prime})$. By definition of the filtration $F_{j}Q_{\epsilon}^{c,c^{\prime}}$, we have $v\geq k-j$. Let $\epsilon^{\prime}:=\frac{\epsilon}{2}$. We now construct a continuous homotopy $H:F_{j}Q^{c,c^{\prime}}_{\epsilon}\times I\to G_{j}S^{c,c^{\prime}}_{\epsilon^{\prime}}$ as follows: If either $m,n$ is the base point, the point $y$ is equivalent to the base point under relation $(iii)$, and we choose the constant homotopy at the base point. So now assume this is not the case. If $v>k-j$, then the point is in $F_{j-1}Q_{\epsilon}^{c,c^{\prime}}$, so goes to the base point in the associated graded $G_{j}Q_{\epsilon}^{c,c^{\prime}}$. Thus, we may choose the constant homotopy on such points as well. So now we may assume as well that $v=k-j$.
Using relation $(i)$, we can change the label of $m$ to $m\xi(l(p_{v}))^{-1}$, add a point $s$ at $(0,\pi_{2}(p_{v})+\frac{\epsilon}{2})$ labeled $\xi(l(p_{v}))$, and replace the label $l(p_{u})$ by $(\xi(l(p_{v})\triangleright)^{-1}l(p_{u})$ for $1\leq u<v$ to get an equivalent point of $G_{j}S_{\epsilon^{\prime}}^{c,c^{\prime}}$.
Now, define the homotopy $H(y,t)$ by continuously changing the location of the point $s$ to $(t,\pi_{2}(p_{v})+\frac{\epsilon}{2})$ at time $t$. At $t=0$, $H$ is given by the composite $F_{j}Q_{\epsilon}^{c,c^{\prime}}\rightarrow G_{j}S_{\epsilon}^{c,c^{\prime}}\rightarrow G_{j}S_{\epsilon^{\prime}}^{c,c^{\prime}}$. Thus it will suffice to show:
  • (1)
    At $t=1$, $H$ is the constant map to the base point.
  • (2)
    $H$ descends to a continuous map $G_{j}S^{c,c^{\prime}}_{\epsilon}\times I\to G_{j}S^{c,c^{\prime}}_{\epsilon^{\prime}}$.
Indeed, if these are shown, then $H$ descends to a nullhomotopy of the inclusion $G_{j}S^{c,c^{\prime}}_{\epsilon}\to G_{j}S^{c,c^{\prime}}_{\epsilon^{\prime}}$.
We first check $(1)$ straightforwardly. For the point $y$, at $t=1$, the coordinates of $s$ are $(1,\pi_{2}(p_{j})+\frac{\epsilon}{2}))$, and the points with second projection smaller than it in decreasing order are $p_{v},\ldots,p_{k}$. Using relation $(ii)$, (which we used to define $\overline{Q}^{*}_{\epsilon}[X_{+},\operatorname{Hur}^{c}_{+},X_{+}]$ as a quotient of $Q_{\epsilon}[X_{+},\operatorname{Hur}^{c}_{+},X_{+}]$,) this is equivalent to a point where $s$ is removed, and $n$ changes to $n^{\prime}:=[(l(p_{v})\cdots l(p_{k}))\triangleright l(s)]\cdot n$. But $l(p_{v})\triangleright l(s)=l(p_{v})\triangleright\xi(l(p_{v}))\notin c^{\prime}$ by definition of $\xi$. Since the action of $l(p_{u})$ for $u>v$ preserves the property of not being in $c^{\prime}$, it follows that $((l(p_{v})\cdots l(p_{k}))\triangleright l(s))\notin c^{\prime}$. Hence, $n^{\prime}$ is the base point. Thus, by relation $(iii)$, this point is the base point of $G_{j}S^{c,c^{\prime}}_{\epsilon^{\prime}}$, showing $(1)$.
It remains to show $(2)$ for each $j\geq 0$. It is easy to see that the equivalence relation induced by $(i),(ii),(iii)$ when restricted to the subspace $F_{j}Q_{\epsilon}^{c,c^{\prime}}$ is the same as the one induced by the map $F_{j}Q_{\epsilon}^{c,c^{\prime}}\to\overline{Q}^{*}_{\epsilon}[X_{+},\operatorname{Hur}^{c}_{+},X_{+}]$. It then suffices to show that $H$ is compatible with $(i),(ii),(iii)$ as well as the relation sending $F_{j-1}Q_{\epsilon}^{c,c^{\prime}}$ to the base point.
By construction, $H$ sends $F_{j-1}Q_{\epsilon}^{c,c^{\prime}}$ to the constant homotopy at the base point, and is also compatible with the equivalence relation $(iii)$.
We next check $H$ is compatible with the equivalence relation $(ii)$. Indeed, the value $l(s)$ remains unchanged under the identification from $(ii)$, unless $(ii)$ is applied to the point $p_{v}$. In this case, the homotopy is the constant homotopy at the base point. Hence, the value of $l(s)$ is compatible with $(ii)$.
It remains to check that $H$ is compatible with the equivalence relation $(i)$. We first deal with the case that the point $p_{r}$ being removed in $(i)$ has a label $l(p_{r})\notin c^{\prime}$. In this case the homotopy is the constant homotopy at the base point, because for any time $t$, the point $p_{r}$ remains on the left hand side and one can use relations $(i)$ and $(iii)$ to identify it with the base point. Therefore, the nullhomotopy is compatible with the relation in this case.
It only remains to consider then the case that the point $p_{r}$ being removed in $(i)$ has $l(p_{r})\in c^{\prime}$. Note that we may assume $r\neq v$ because $l(p_{v})\notin c^{\prime}$. Because we are working on the subquotient $G_{j}S^{c,c^{\prime}}_{\epsilon^{\prime}}$, if we apply $(i)$ for some point $p_{r}$ with $r>v$, the point becomes identified with the base point, and so $H$ is the constant homotopy on the base point. (This is the reason for introducing the filtration $F_{j}$.) Hence, it only remains to deal with the case that $r<v$ at a point $y$ presented as $(m,n,x,l:x\to c)$, as in ()($\dagger$)( † ) above.
We now verify relation $H$ is compatible with the equivalence relation $(i)$ when $r<v$ and $l(p_{r})\in c^{\prime}$. If we first apply the nullhomotopy before applying $(i)$, $m$ changes to $m\cdot[\xi(l(p_{v}))^{-1}]$, we add a point $s$ at $(t,\pi_{2}(p_{v})+\frac{\varepsilon}{2})$ labeled by $\xi(l(p_{v}))$, and the points $p_{u}$ for $1\leq u<v$ have labels which change from $l(p_{u})$ to $(\xi(l(p_{v}))\triangleright)^{-1}l(p_{u})$. If we now apply $(i)$, $m$ then becomes $m\cdot[\xi(l(p_{v}))]^{-1}\cdot[(\xi(l(p_{v}))\triangleright)^{-1}l(p_{r})]$, and the label on $p_{w}$ for $1\leq w<r$ becomes $((\xi(l(p_{v}))\triangleright)^{-1}l(p_{r}))\triangleright((\xi(l(p_{v}))\triangleright)^{-1}l(p_{w}))$.
If instead we apply $(i)$ before applying the nullhomotopy, $m$ changes to $m\cdot[l(p_{r})]$ and the points $p_{w}$ with $1\leq w<r$ have label which changes from $l(p_{w})$ to $l(p_{r})\triangleright l(p_{w})$. When we next apply the nullhomotopy, we introduce a point $s$ at $(t,\pi_{2}(p_{v})+\frac{\varepsilon}{2})$ labeled by $\xi(l(p_{v}))$, we change $m\cdot[l(p_{r})]$ to $m\cdot[l(p_{r})]\cdot[\xi(l(p_{v}))]^{-1}$, we change the label on $p_{w}$ with $1\leq w<r$ from $l(p_{r})\triangleright l(p_{w})$ to $(\xi(l(p_{v}))\triangleright)^{-1}(l(p_{r})\triangleright l(p_{w}))$, and we change the label on $p_{u}$ with $r<u<v$ from $l(p_{u})$ to $(\xi(l(p_{v}))\triangleright)^{-1}l(p_{u})$.
To conclude, we only need observe that the two points obtained above agree we need to verify the two relations
(4.9) $m\cdot[\xi(l(p_{v}))]^{-1}\cdot[(\xi(l(p_{v}))\triangleright)^{-1}l(p_{r})]$ $=m\cdot[l(p_{r})]\cdot[\xi(l(p_{v}))]^{-1}$
(4.10) $((\xi(l(p_{v}))\triangleright)^{-1}l(p_{r}))\triangleright((\xi(l(p_{v}))\triangleright)^{-1}l(p_{w}))$ $=(\xi(l(p_{v}))\triangleright)^{-1}(l(p_{r})\triangleright l(p_{w})).$
First, recall the defining relation of Hurwitz space from § 4.1 implies that the positive generator of the braid group identifies $[g]\cdot[h]=[h]\cdot[h\triangleright g]$. Therefore, taking $h=\xi(l(p_{v}))$, and using that $[\xi(l(p_{v}))]$ acts invertibly on $M$, the action on $M$ factors through a group where the relation $[\xi(l(p_{v}))]^{-1}\cdot[g]=[\xi(l(p_{v}))\triangleright g]\cdot[\xi(l(p_{v}))]^{-1}$ is satisfied for each $g\in c$. Hence, (4.9) holds by taking $g=(\xi(l(p_{v}))\triangleright)^{-1}l(p_{r})$. Finally, (4.10) is equivalent to property $(2)$ in the definition of rack § 4.1. ∎
Finally, we verify that in the non-splitting case, every non-empty subrack is self-normalizing:
Lemma 4.5.0.
Let $c$ be a non-splitting rack, and $c^{\prime}$ a nonempty subrack. Then $c^{\prime}$ is self-normalizing.
Proof.
Let $S$ be the set of elements $y\in c$ such that $y\triangleright c^{\prime}=c^{\prime}$. Then, $S\subset c$ is a subrack of $c$ containing $c^{\prime}\subset S$ as a subrack. Further, $c^{\prime}$ is closed under the operations $x\triangleright$ for each $x\in S$. The non-splitting hypothesis applied to $S$ implies $S$ is connected, but we also find that $c^{\prime}\subset S$ is a connected subrack. It follows that $S=c^{\prime}$. ∎
4.5.1. Proof of Theorem 4.2.1
By § 4.4, it is enough to show that for each nonempty subrack $c^{\prime}\subset c$, $v[\alpha_{c^{\prime}}^{-1}]^{\wedge}_{\alpha_{c-c^{\prime}}}$ is an equivalence. When $c^{\prime}=c$, the result is § 4.5. When $c^{\prime}\subset c$ is a proper subrack, we have a commutative square
[…]
We wish to show the left vertical map is an equivalence, and so we will show the other three maps in the diagram are equivalences. The upper horizontal map is an equivalence by § 4.5, since the subrack is self-normalizing by § 4.5. The lower horizontal map is an equivalence because any homology class in $A^{\prime}[\alpha_{c^{\prime}}^{-1}]$ supported on a component that does not come from $\operatorname{Conf}^{c^{\prime}}$ is divisible by $\alpha_{c-c^{\prime}}$, and the completion is projection onto the summand where $\alpha_{c-c^{\prime}}$ acts by 00 by § 3.3. Finally, the right vertical map is an equivalence by § 4.5, where $c$ is replaced with $c^{\prime}$. ∎
5. Deducing the moments of the Cohen–Lenstra heuristics
In this section, we explain how we deduce Theorem 1.2.1, and hence Theorem 1.1.1, from our computation of the stable homology of Hurwitz spaces. The structure of this section is as follows. We first give some examples of nonsplitting conjugacy classes relevant to the Cohen–Lenstra heuristics in § 5.1. We then prove § 5.2, which gives a criterion for computing the finite field point counts of a space in terms of its stable homology. In § 5.3, we apply this lemma to Hurwitz spaces, and obtain a point count in terms of the number of geometrically irreducible components of that Hurwitz space. In § 5.3, we count the components of the relevant Hurwitz spaces. In § 5.4, we deduce Theorem 1.1.1 from the finite field point counts of these Hurwitz spaces. Finally, in § 5.5, we compute the moments predicted by the non-abelian Cohen–Lenstra heuristics Theorem 1.3.1.
5.1. Examples of non-splitting conjugacy classes
An important example of a non-splitting $(G,c)$ for our applications to the Cohen–Lenstra heuristics is the following.
Example 5.1.0.
Let $H$ be a finite abelian group of odd order. Write $G=H\rtimes\mathbb{Z}/2\mathbb{Z}$, with $H$ a finite abelian group of odd order and the nontrivial element of $\mathbb{Z}/2\mathbb{Z}$ acting on $H$ via inversion. Let $c\subset G$ denote the conjugacy class of order $2$ elements, which is the preimage of the nontrivial element of $\mathbb{Z}/2\mathbb{Z}$ under the surjection $G\to G/H\simeq\mathbb{Z}/2\mathbb{Z}$. That such $(G,c)$ is non-splitting follows from Sylow’s theorems, since elements of order $2$ in any subgroup $H\subset G$ generate a Sylow subgroup, and hence are all conjugate. This was also observed in [EVW16, Lemma 3.2].
The next example will not be used in this paper, but we mention it to give another example relevant to a generalization of the Cohen–Lenstra heuristics.
Example 5.1.0.
Let $H$ be a finite abelian group of order prime to $6$ with an $S_{3}$ action. Let $G:=H\rtimes S_{3}$. Let $c\subset G$ denote the conjugacy class of elements which map to transpositions in $S_{3}$. A similar argument to § 5.1 by considering $2$-Sylow subgroups shows that $(G,c)$ is again non-splitting.
Remark 5.1.0.
The verification of Theorem 1.4.5 for $(G,c)$ as in § 5.1 is closely connected to the Cohen–Lenstra–Martinet heuristics for torsion in class groups of cubic fields of squarefree discriminant, see [EVW12, §6.1].
5.2. A preliminary lemma on point counting
The next lemma is quite standard, and its proofs essentially appears in many places, such as [EVW16, Theorem 8.8] and also [EL23, Theorem 9.2.3]. The next lemma is a formal consequence of the Grothendieck–Lefschetz trace formula and Deligne’s bounds on eigenvalues of Frobenius.
Remark 5.2.0.
We prove the following lemma in the more general setting of stacks, but we will only apply it in the context of schemes.
In what follows, if $x$ is a complex number, we use $\|x\|$ to denote its absolute value, while if $X$ is a set, we use $|X|$ to denote its cardinality. For $\mathscr{X}$ a Deligne-Mumford stack, we follow the standard convention that $|\mathscr{X}(\mathbb{F}_{q})|$ denotes $\sum_{x\in\mathscr{X}(\mathbb{F}_{q})}\frac{1}{|\operatorname{Aut}(x)|}$, so points are weighted inversely proportionally to their automorphism groups. Of course, when $\mathscr{X}$ is a scheme, this agrees with the usual notion of $|\mathscr{X}(\mathbb{F}_{q})|$.
Lemma 5.2.0.
Fix a prime power $q$ and a sequence $\{Y_{n}\}_{n\in S}$ of nonempty smooth Deligne-Mumford stacks over $\operatorname{Spec}\mathbb{F}_{q}$ with $Y_{n}$ of pure dimension $n$, where $n$ traverses over an infinite sequence of increasing integers $S\subset\mathbb{Z}_{\geq 0}$. Suppose that the following cohomological conditions are satisfied:
  1. (1) Stability:
    Fix $i\geq 0$. There is some prime $\ell$, possibly depending on $n$, and a fixed $\mathbb{Q}_{\ell}$ vector space $V_{i}$ with the action of geometric Frobenius, $\operatorname{Frob}_{q}$, such that $V_{i}$ with its Frobenius action is independent of $n$. Assume there are constants $I$ and $J$, independent of $n$, so that whenever $n>Ii+J$, there is a fixed vector space $V_{i}$ with Frobenius action, so that $\operatorname{tr}(\operatorname{Frob}_{q}|V_{i})=\operatorname{tr}(\operatorname{Frob}_{q}|H^{i}(Y_{n,\overline{\mathbb{F}}_{q}};\mathbb{Q}_{\ell}))$.
  2. (2) Exp. Bound:
    There are constants $C,C^{\prime}$ independent of $n$ so that for any $i$, $\dim H^{i}(Y_{n,\overline{\mathbb{F}}_{q}};\mathbb{Q}_{\ell})\leq C^{\prime}C^{i}$.
Then, if $q>C^{2}$,
$\left\|\frac{\left|Y_{n}(\mathbb{F}_{q})\right|}{q^{n}}-\sum_{i=0}^{\infty}(-1)^{i}\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|V_{i})\right\|\leq\frac{2C^{\prime}}{1-\frac{C}{\sqrt{q}}}\left(\frac{C}{\sqrt{q}}\right)^{\frac{n-J}{I}}.$
Proof.
To ease notation, let $X_{n}:=Y_{n}\times_{\operatorname{Spec}\mathbb{F}_{q}}\operatorname{Spec}\overline{\mathbb{F}}_{q}$. Note that our stability assumption holds whenever $n>Ii+J$, or equivalently when $i<\frac{n-J}{I}$. Using our stability assumption (1) and the Grothendieck–Lefschetz trace formula [Beh93, Theorem 3.1.2], we have
(5.1) $\frac{\left|Y_{n}(\mathbb{F}_{q})\right|}{q^{n}}=$ $\sum_{i=0}^{2n}(-1)^{i}\operatorname{tr}\left(\operatorname{Frob}_{q}^{-1}|H^{i}(X_{n};\mathbb{Q}_{\ell})\right)$
$=$ $\sum_{0\leq i<\frac{n-J}{I}}(-1)^{i}\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|V_{i})+\sum_{i\geq\frac{n-J}{I}}(-1)^{i}\operatorname{tr}\left(\operatorname{Frob}_{q}^{-1}|H^{i}(X_{n};\mathbb{Q}_{\ell}).\right)$
Subtracting $\sum_{i=0}^{\infty}(-1)^{i}\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|V_{i})$ from both sides of (5.1) and taking absolute values, we obtain
(5.2) $\left\|\frac{\left|Y_{n}(\mathbb{F}_{q})\right|}{q^{n}}-\sum_{i=0}^{\infty}(-1)^{i}\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|V_{i})\right\|$
$\leq\sum_{i\geq\frac{n-J}{I}}\left\|\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|V_{i})\right\|+\sum_{i\geq\frac{n-J}{I}}(-1)^{i}\left\|\operatorname{tr}\left(\operatorname{Frob}_{q}^{-1}|H^{i}(X_{n};\mathbb{Q}_{\ell})\right)\right\|$
We claim that $\sum_{i=0}^{\infty}(-1)^{i}\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|V_{i})$ converges absolutely as a function of $q$ for $\sqrt{q}>C$. Indeed, using Sun’s generalization of Deligne’s bounds for algebraic stacks [Sun12, Theorem 1.4] and our exponential bound assumption $(2)$ we obtain $\|\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|V_{i})\|\leq\frac{C^{\prime}C^{i}}{q^{i/2}}$. It follows that
$\sum_{i=0}^{\infty}\|(-1)^{i}\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|V_{i})\|\leq C^{\prime}\sum_{i=0}^{\infty}\left(\frac{C}{\sqrt{q}}\right)^{i}=C^{\prime}\frac{1}{1-\frac{C}{\sqrt{q}}}.$
We similarly find that
(5.3) $\sum_{i=\frac{n-J}{I}}^{\infty}\|(-1)^{i}\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|V_{i})\|\leq C^{\prime}\sum_{i=\frac{n-J}{I}}^{\infty}\left(\frac{C}{\sqrt{q}}\right)^{i}=C^{\prime}\left(\frac{C}{\sqrt{q}}\right)^{\frac{n-J}{I}}\frac{1}{1-\frac{C}{\sqrt{q}}}.$
This bounds the first term on the right hand side of (5.2). It remains to bound the second term. Again, using Sun’s generalizations of Deligne’s bounds and our exponential bound assumption $(2)$, we similarly obtain $\|\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|H^{i}(X_{n};\mathbb{Q}_{\ell}))\|\leq\frac{C^{\prime}C^{2n-j}}{q^{i/2}}$. Hence, so long as $C<\sqrt{q}$, we can bound
$\sum_{i\geq\frac{n-J}{I}}(-1)^{i}\left\|\operatorname{tr}\left(\operatorname{Frob}_{q}^{-1}|H^{i}(X_{n};\mathbb{Q}_{\ell})\right)\right\|$ $\leq C^{\prime}\sum_{i\geq\frac{n-J}{I}}\left(\frac{C}{\sqrt{q}}\right)^{i}$
$=C^{\prime}\left(\frac{C}{\sqrt{q}}\right)^{\frac{n-J}{I}}\frac{1}{1-\frac{C}{\sqrt{q}}}.\qed$
5.3. Applying the lemma to count points on Hurwitz spaces
We next show that the conditions of § 5.2 are satisfied for Hurwitz spaces using our computation of the stable homology. We use § 5.2 to obtain a count of the finite field valued points of relevant Hurwitz spaces in terms of the number of their geometrically irreducible components.
Proposition 5.3.0.
Suppose $G$ is a finite group, $c\subset G$ is a conjugacy class, and $(G,c)$ is non-splitting. Let $q$ be an prime power satisfying $\gcd(q,|G|)=1$. If $Y_{n}\subset\operatorname{CHur}^{G,c}_{n,\mathbb{F}_{q}}$ is a union of connected components which has $r$ geometrically irreducible components for each $n$ sufficiently large, there exist constants $C,I,J$, all independent of $n$ and $q$, but possibly depending on $r$, so that, as long as $q>C^{2}$,
(5.4) $\left\|\frac{\left|Y_{n}(\mathbb{F}_{q})\right|}{q^{n}}-r\left(1-\frac{1}{q}\right)\right\|\leq\frac{2C}{1-\frac{C}{\sqrt{q}}}\left(\frac{C}{\sqrt{q}}\right)^{\frac{n-J}{I}}.$
An important special case of the above is when we take $Y_{n}=\operatorname{CHur}^{G,c,g}_{n,\mathbb{F}_{q}}$ for some $g\in G$. It may be useful to recall the notation $\operatorname{CHur}^{G,c,g}_{n,B}$ for components of $\operatorname{CHur}^{G,c}_{n,B}$ whose boundary monodromy is $g$, as defined in § 2.2
Proof.
This follows from § 5.2, taking the constant $C^{\prime}$ there to be the same as $C$, once we verify its hypotheses. We will take $\ell>\max(n,|G|,q)$. To simplify matters slightly, we will assume all irreducible components of $Y_{n}$ are geometrically irreducible, as components which are not geometrically irreducible have no $\mathbb{F}_{q}$ points, by smoothness of the $Y_{n}$. Let $B$ denote the spectrum of a henselian dvr with residue field $\mathbb{F}_{q}$ and generic point of characteristic 00. In this case, there is a scheme $Z_{n}$ which is a union of $r$ irreducible components of $\operatorname{CHur}^{G,c,g}_{n,B}$, so that $Y_{n}=Z_{n}\times_{B}\operatorname{Spec}\mathbb{F}_{q}$. We let $Y_{n,\mathbb{C}}:=Z_{n}\times_{B}\operatorname{Spec}\mathbb{C}$ and $Y_{n,\overline{\mathbb{F}}_{q}}:=Y_{n}\times_{\operatorname{Spec}\mathbb{F}_{q}}\operatorname{Spec}\overline{\mathbb{F}}_{q}.$
We first verify the stability assumption from § 5.2(1). To do this, we start by constructing the vector spaces $V_{i}$, with the Frobenius action. For $B$ a scheme, we use $\operatorname{Conf}_{n,B}$ denote the configuration space parameterizing $n$ unordered points on $\mathbb{A}^{1}_{B}$. (Previously, we had used $\operatorname{Conf}_{n}$ as notation for this when $B=\operatorname{Spec}\mathbb{C}$.) Now, define $V_{i}:=H^{i}(\operatorname{Conf}_{n,\overline{\mathbb{F}}_{q}};\mathbb{Q}_{\ell})^{\oplus r}$ as a vector space with Frobenius action, where $r$ is the number of geometrically irreducible components of $Y_{n}$ for sufficiently large $n$. Note that the dimension of $V_{i}$ is independent of $n$ and $q$ once $n>2$ as a vector space with Frobenius action, and the Frobenius action is given by and identity matrix of dimension $r$ if $i=0$, $q$ times the identity matrix of dimension $r$ when $i=1$, and the 00 matrix if $i>1$. In particular, we have $\sum_{i=0}^{\infty}(-1)^{i}\operatorname{tr}(\operatorname{Frob}_{q}^{-1}|V_{i})=r(1-1/q)$.
To verify § 5.2(1), we next need to show the trace of Frobenius on $V_{i}$ agrees with the trace of Frobenius on $H^{i}(Y_{n,\overline{\mathbb{F}}_{q}};\mathbb{Q}_{\ell})$ for $n>iI+J$. There is a map
(5.5) $H^{i}(\operatorname{Conf}_{n,\mathbb{C}};\mathbb{Q}_{\ell})^{\oplus r}\to H^{i}(Y_{n,\mathbb{C}};\mathbb{Q}_{\ell}),$
which is an isomorphism for $n>iI+J$ by Theorem 1.4.5. It follows from [EVW16, Proposition 7.7] that there is an isomorphism
(5.6) $H^{i}(Y_{n,\overline{\mathbb{F}}_{q}};\mathbb{Q}_{\ell})\simeq H^{i}(Y_{n,\mathbb{C}};\mathbb{Q}_{\ell})$
once we exhibit a suitable normal crossings compactification of configuration space. This is afforded by [FM94, Theorem 3]. (See also the proof of [EL23, Theorem 9.2.3] for further explanation.) Therefore, (5.6) restricts to an isomorphism
(5.7) $H^{i}(Y_{n,\overline{\mathbb{F}}_{q}};\mathbb{Q}_{\ell})\simeq H^{i}(Y_{n,\mathbb{C}};\mathbb{Q}_{\ell}).$
An analogous argument then shows
(5.8) $H^{i}(\operatorname{Conf}_{n,\overline{\mathbb{F}}_{q}};\mathbb{Q}_{\ell})^{\oplus r}\simeq H^{i}(\operatorname{Conf}_{n,\mathbb{C}};\mathbb{Q}_{\ell})^{\oplus r}.$
Combining (5.5), (5.7), and (5.8), we obtain an Frobenius equivariant isomorphism $U_{i}:H^{i}(\operatorname{Conf}_{n,{\overline{\mathbb{F}}_{q}}};\mathbb{Q}_{\ell})^{\oplus r}\to H^{i}(Y_{n,\overline{\mathbb{F}}_{q}};\mathbb{Q}_{\ell})$. This implies $\operatorname{tr}(\operatorname{Frob}_{q}|V_{i})=\operatorname{tr}(\operatorname{Frob}_{q}|H^{i}(Y_{n,\overline{\mathbb{F}}_{q}};\mathbb{Q}_{\ell}))$.
Finally, we verify the exponential bound assumption, § 5.2(2). This condition with $C=C^{\prime}$ nearly follows from [EVW16, Proposition 7.8], using our assumption that $\ell>\max(n,|G|,q)$ as well as the comparison between étale and Betti cohomology. The slight difference we need is that [EVW16, Proposition 7.8] is proven for Hurwitz stacks, instead of pointed Hurwitz spaces. However, the same proof works for pointed Hurwitz spaces, since the analog of [EVW16, Corollary 6.2] without quotienting by the $G$ conjugation action can be deduced in the same way from [EVW16, Theorem 6.1]. We note that we can take the value of $C$ to be independent of $q$ and $\ell$ satisfying our constraints because $\dim H^{i}(Y_{n},\mathbb{Q}_{\ell})\leq\dim H^{i}(\operatorname{CHur}^{G,c,g}_{n,\mathbb{C}};\mathbb{Q}_{\ell})$ is independent of $q$ and $\ell$. ∎
Remark 5.3.0.
The value of the constant $C^{2}$ in § 5.3 (which is also the value of the constant $C$ in Theorem 1.1.1 and Theorem 1.2.1) can be explicitly computed in terms of the point at which the homology of a certain Hurwitz space stabilizes. The computation of this constant does not use our work, but entirely follows from the proof of [EVW16]. For explicitly determining the constant, it was also useful for us to consult [RW20] and [EL23]. It may be possible to reduce the value of the constant $C$, so what we describe below is really an upper bound for $C^{2}$.
Using [RW20, Proposition 4.4, Proposition 8.1, Theorem 7.1, and Corollary 7.4], we find that the stabilization operator $U$ induces an isomorphism $H_{d}(\operatorname{CHur}^{G,c}_{n})\to H_{d}(\operatorname{CHur}^{G,c}_{n+\deg U})$ once $n\geq(N_{0}+2)(2d+1)+N_{0}$; here $N_{0}$ is the is the constant defined in [RW20, Proposition 4.4], describing the $n$ after which the image of $U$ on $H_{0}(\operatorname{Hur}_{n}^{c})$ stabilizes. Then, tracing through the proof of [EL23, Theorem 4.1.3, Proposition 4.3.3, Theorem 9.2.4], we find it is sufficient to take $q>4((2|c|)^{3N_{0}+\deg U}+1)^{4}$.
For example, in the particular case that $c$ is the conjugacy class of transpositions in $S_{3}$, $|c|=3$, we may take $\deg U=2$, and one may explicitly compute $N_{0}=5$, so we get that we can take $q>4\cdot(6^{17}+1)^{4}$ in § 5.3 for $G=S_{3}$ and in Theorem 1.2.1 for $H=\mathbb{Z}/3\mathbb{Z}$.
§ 5.3 lets us count the points of relevant Hurwitz spaces once we compute their number of geometrically irreducible components. We do so in the next proposition. It will be useful to recall the notion of being split completely over $\infty$ from § 2.3.
Proposition 5.3.0.
Suppose $G,H$, and $c$ are as in § 5.1. Let $q$ be an odd prime power with $\gcd(q,|H|)=1$. Take $h:=\gcd(|H|,q-1)$, and fix $g\in G$. For $n$ sufficiently large, there are $|\wedge^{2}H[h]|$ many geometrically irreducible components of $\operatorname{CHur}^{G,c,g}_{n,\mathbb{F}_{q}}$, parameterizing $H$-covers of hyperelliptic curves which are split completely over $\infty$. There are constants constant $C,I,J$, independent of $n$ and $q$, and depending only on $H$, so that for $q$ satisfying $\gcd(|H|,q)=1$, $q>C^{2}$, and any $g\in G$,
(5.9) $\left\|\frac{\left|\operatorname{CHur}^{G,c,g}_{n,\mathbb{F}_{q}}(\mathbb{F}_{q})\right|}{q^{n}}-|\wedge^{2}H[h]|\left(1-\frac{1}{q}\right)\right\|\leq\frac{2C}{1-\frac{C}{\sqrt{q}}}\left(\frac{C}{\sqrt{q}}\right)^{\frac{n-J}{I}}.$
Proof.
Once we verify the claim about the number of geometrically irreducible components, (5.9) follows from § 5.3. The remainder of the proof is essentially explained in [SW23, Theorem 3.1], and we now spell out the details. In the case that $n$ is even, [LWZB24, Lemma 10.3] yields a bijection between components of $\operatorname{CHur}^{G,c,g}_{n,\overline{\mathbb{F}}_{q}}$ and components of $\operatorname{CHur}^{G,c,g}_{n,\mathbb{C}}$. Then, [LWZB24, Theorem 12.1(2)] and [LWZB24, Corollary 12.6] yield that the number of irreducible components of $\operatorname{CHur}^{G,c,g}_{n,\mathbb{C}}$ can be identified bijectively with a torsor for the set of elements of a certain finite group $H_{2}(G,c)$ which are equal to their $q$th power. There is an identification $H_{2}(G,c)\simeq(\wedge^{2}H)^{\mathbb{Z}/2\mathbb{Z}}$ as is explained in the course of the proof of [SW23, Theorem 3.1]. Moreover, in our case, $\mathbb{Z}/2\mathbb{Z}$ acts on $H$ by inversion, and hence it acts trivially on $\wedge^{2}H$, and so $(\wedge^{2}H)^{\mathbb{Z}/2\mathbb{Z}}\simeq\wedge^{2}H$. In other words, the number of geometrically irreducible components is precisely the $q-1$ torsion subgroup of $\wedge^{2}H$, $\wedge^{2}H[q-1]$. Since $h:=\gcd(|H|,q-1)$, this is also equal to $\wedge^{2}H[h]$, and hence there are $\left|\wedge^{2}H[h]\right|$ geometrically irreducible components of $\operatorname{CHur}^{G,c,g}_{n,{\mathbb{F}}_{q}}$.
We conclude by explaining why the above argument still works when $n$ is odd. The setting of [SW23] and [LWZB24] is slightly different from ours because they make the additional assumption that all $G$-covers are unramified over $\infty\in\mathbb{P}^{1}$. However, with our definition of Hurwitz spaces, § 2.1, the straightforward generalizations of all of the results cited in the above paragraph easily carry through. The main minor modification needed is that instead of working with marked covers $X\to\mathbb{P}^{1}_{S}$ as in [Woo21], one should instead work with marked covers $X\to\mathscr{P}_{S}$, where $\mathscr{P}_{S}$ is a root stack with an order $2$ root over the $\infty$ section, $\infty:S\to\mathbb{P}^{1}_{S}$. Additionally, in [LWZB24, Theorem 12.4], one no longer requires that the inertia elements $g_{i}$ need multiply to $\operatorname{id}$, since the corresponding cover $X\to\mathbb{P}^{1}$ will be branched at $\infty$. With these minor modifications in place, the rest of the argument goes through for $n$ odd and completes the proof.
At this point, the constant $C$ may depend on the values of $h$ and $g$, but since there are at most $|G|^{2}$ many choices for the pair $(h,g)$, we can replace $C$ with the constant from this finite set of values which maximizes the right hand side of (5.9). This concludes the proof. ∎
5.4. Proof of Theorem 1.2.1
We conclude by proving our main result, Theorem 1.2.1, which determines the $H$-moment of the class group of quadratic extensions of $\mathbb{F}_{q}(t)$ for $q$ sufficiently large relative to $H$. It will be useful to recall our notation for the class group of a function field extension of $\mathbb{F}_{q}(t)$ and for the set $\mathcal{MH}_{n,q}$ of monic smooth hyperelliptic curves over $\mathbb{F}_{q}$, defined prior to Theorem 1.1.1. The proof of Theorem 1.2.1 essentially follows immediately from § 5.3 and a little bit of class field theory. In fact, we prove a slight refinement of Theorem 1.2.1 which moreover quantifies the error term in the asymptotic count of torsion in class groups of quadratic fields.
Theorem 5.4.1.
Suppose $H$ is a finite abelian group of odd order. Fix $q$ an odd prime power with $\gcd(|H|,q)=1$ and let $h:=\gcd(|H|,q-1)$. There is are integer constants $I$, $J$, and $C$ depending only on $H$ so that, if $q>C^{2}$,
(5.10) $\left\|\frac{\sum_{K\in\mathcal{MH}_{n,q}}\left|\operatorname{Surj}(\operatorname{Cl}(\mathscr{O}_{K}),H)\right|}{\sum_{K\in\mathcal{MH}_{n,q}}1}-\alpha_{n}\left|\wedge^{2}H[h]\right|\right\|\leq\left(\frac{1}{1-\frac{1}{q}}\right)\frac{2C}{1-\frac{C}{\sqrt{q}}}\left(\frac{C}{\sqrt{q}}\right)^{\frac{n-J}{I}},$
where $\alpha_{n}=1$ if $n$ is odd and $\alpha_{n}=\frac{1}{|H|}$ if $n$ is even.
Proof.
Note that the class group $\operatorname{Cl}(\mathscr{O}_{K})$ is identified with the Galois group of the maximal abelian unramified extension of $K$, completely split completely over $\infty\in\mathbb{P}^{1}_{\mathbb{F}_{q}}$, as follows from class field theory [Hay92, 15.6]. (Also see the paragraph following [Hay92, 15.6] to show that what is called $H_{A}$ there is in fact the maximal unramified abelian extension split completely over $\infty$.) Recall $G=H\rtimes(\mathbb{Z}/2\mathbb{Z})$.
We first deal with the case $n$ is even. When $n$ is even, by § 2.3, every $K\in\mathcal{MH}_{n,q}$ corresponds bijectively to a smooth proper curve $X$ split completely over $\infty$. For $n$ even, we can then identify $\sum_{K\in\mathcal{MH}_{n,q}}\left|\operatorname{Surj}(\operatorname{Cl}(\mathscr{O}_{K}),H)\right|=\frac{\left|\operatorname{CHur}^{G,c,\operatorname{id}}_{n,\mathbb{F}_{q}}(\mathbb{F}_{q})\right|}{\#H}$ by combining [LWZB24, Lemma 9.3 and Lemma 10.2]. Note that the denominator $\sum_{K\in\mathcal{MH}_{n,q}}1=q^{n}-q^{n-1}$ for $n>1$, as it is identified with the $\mathbb{F}_{q}$ points of unordered configuration space of $n$ points in $\mathbb{A}^{1}_{\mathbb{F}_{q}}$. The case that $n$ is even now follows from § 5.3 upon dividing both sides of (5.9) by $1-\frac{1}{q}$.
It remains to deal with the case that $n$ is odd. Recall notation for components with specified boundary monodromy from § 2.2. The case that $n$ is odd is similar to the even case, except that we now claim $\sum_{K\in\mathcal{MH}_{n,q}}\left|\operatorname{Surj}(\operatorname{Cl}(\mathscr{O}_{K}),H)\right|=\frac{\left|\operatorname{CHur}^{G,c}_{n,\mathbb{F}_{q}}(\mathbb{F}_{q})\right|}{\#H},$ with no condition imposed on the boundary monodromy. Indeed, any $G$ cover corresponding to a point of $\operatorname{CHur}^{G,c}_{n,\mathbb{F}_{q}}$ has inertia type $c$ at $\infty$, and so the the $G$ extension corresponds to an $H$ extension $L/K$ followed by a quadratic extension $K/\mathbb{F}_{q}(t)$, where the quadratic extension is ramified over $\infty$ and the $H$ extension is split completely over $\infty$. Here we are using that the quadratic extension $K/\mathbb{F}_{q}(t)$ here is split completely over $\infty$ in the stacky sense of § 2.3, and we are using § 2.3 to identify $\mathcal{MH}_{n,q}\simeq\operatorname{Hur}^{\mathbb{Z}/2\mathbb{Z},\{1\}}_{n,\mathbb{F}_{q}}(\mathbb{F}_{q})$. Hence, via the identification $\operatorname{Hur}^{\mathbb{Z}/2\mathbb{Z},\{1\}}_{n,\mathbb{F}_{q}}$ with the configuration space of $n$ unordered points in $\mathbb{A}^{1}_{\mathbb{F}_{q}}$, we have $\sum_{K\in\mathcal{MH}_{n,q}}1=q^{n}-q^{n-1}$. Altogether, there are $|H|$ possible values for the boundary monodromy at $\infty$ when $n$ is odd, which is why the $H$-moment in the case that $n$ is odd is $|H|$ times the moment in the case that $n$ is even. Thus, the result again follows by summing the result of § 5.3 over those $|H|$ elements $g\in c=G-H$. ∎
5.5. Non-abelian Cohen–Lenstra moments
We next prove a theorem implying Theorem 1.3.1, the non-abelian Cohen–Lenstra moments with roots of unity.
Theorem 5.5.1.
Using notation from § 1.3, suppose $H$ is an admissible group with a $\mathbb{Z}/2\mathbb{Z}$ action. Fix $q$ an odd prime power with $\gcd(|H|,q)=1$. Let $\delta:\hat{\mathbb{Z}}(1)_{(2q)^{\prime}}\to H_{2}(H\rtimes\mathbb{Z}/2\mathbb{Z},\mathbb{Z})_{(2q)^{\prime}}$ be a group homomorphism with $\operatorname{ord}(\operatorname{im}\delta)\mid q-1.$ There is are integer constants $I$, $J$, and $C$ depending only on $H$ so that, if $q>C^{2}$ and $n$ is even,
(5.11) $\left\|\frac{\sum_{K\in\mathcal{MH}_{n,q}}\left|\{\pi\in\operatorname{Surj}_{\mathbb{Z}/2\mathbb{Z}}\left(G^{\sharp}_{\emptyset}(K),H\right):\pi_{*}\circ\omega_{K^{\sharp}/K}=\delta\}\right|}{\sum_{K\in\mathcal{MH}_{n,q}}1}-\frac{1}{[H:H^{\mathbb{Z}/2\mathbb{Z}}]}\right\|$
$\leq\left(\frac{1}{1-\frac{1}{q}}\right)\frac{2C}{1-\frac{C}{\sqrt{q}}}\left(\frac{C}{\sqrt{q}}\right)^{\frac{n-J}{I}},$
where $[H:H^{\mathbb{Z}/2\mathbb{Z}}]$ denotes the index of the $\mathbb{Z}/2\mathbb{Z}$ invariants of $H$ in $H$.
Proof.
Let $G:=H\rtimes\mathbb{Z}/2\mathbb{Z}$ and let $c\subset G$ denote the conjugacy class of order $2$ elements. Note that $(G,c)$ is nonsplitting by Sylow’s theorem, see [EVW16, Lemma 3.2].
We claim there is a unique geometrically irreducible component of $\operatorname{CHur}^{G,c,\operatorname{id}}_{n,\mathbb{F}_{q}}$, whose $\omega$-invariant is $\delta$ precisely when $\operatorname{ord}(\operatorname{im}\delta)\mid q-1$ and $n$ is even, where the $\omega$-invariant of a component was defined (and shown to be well defined) in [Liu22, Proposition 4.3]. To prove this claim, we note that $\operatorname{CHur}^{\mathbb{Z}/2\mathbb{Z},\iota}_{n,\mathbb{F}_{q}}$ has a unique geometrically irreducible component, where $\iota$ is the nontrivial element of $\mathbb{Z}/2\mathbb{Z}$. The claim then follows from [Liu22, Lemma 4.4], [Liu22, Lemma 4.5], and [LWZB24, Proposition 12.7]. Here, we are also using that there is a unique component of the Hurwitz scheme parameterizing quadratic extensions of $\mathbb{F}_{q}(t)$ of reduced discriminant $q^{n}$ precisely when $n$ is even, as follows from Riemann Hurwitz, and so the $M_{a}$ appearing in [Liu22, Lemma 4.4] can all be taken to be $2$ and the $E_{a}$ can be taken to be the singleton 00, using the equality in [Liu22, Lemma 4.5].
The result now follows from (5.4) with $r=1$ and [Liu22, Lemma 4.6], which accounts for the factor of $[H:H^{\mathbb{Z}/2\mathbb{Z}}]$. ∎
Remark 5.5.1.
We now explain why we make the restrictions we do in Theorem 5.5.1, even though conjectures on non-abelian Cohen–Lenstra heuristics have been posed more generally. Before explaining our restrictions, we first explain which aspects of these conjectures have greater generality. First, [Woo19, Conjecture 5.1] is stated more generally for imaginary quadratic fields and for certain subgroups of $H\wr S_{2}$ instead of just for $H\rtimes S_{2}$. Moreover, [Liu22, Conjecture 1.2] (generalizing the closely related [LWZB24, Conjecture 1.3]) is stated more generally for $\Gamma$ extensions $K/\mathbb{F}_{q}(t)$, where $\Gamma$ is an arbitrary finite group, instead of just $\mathbb{Z}/2\mathbb{Z}$ extensions.
The reason we restrict to the case that $H$ has odd order and limit ourselves to counting Galois $H\rtimes\mathbb{Z}/2\mathbb{Z}$ extensions of $\mathbb{F}_{q}(t)$ is that in these cases, we know the conjugacy class of order $2$ elements in $H\rtimes\mathbb{Z}/2\mathbb{Z}$ is non-splitting. Hence, in this case, we are able to apply Theorem 1.4.5.
We now explain why we restrict to the field extension $K/\mathbb{F}_{q}(t)$ being a $\mathbb{Z}/2\mathbb{Z}$ extension, while [Liu22, Conjecture 1.2] and [LWZB24, Conjecture 1.3] are stated for $\Gamma$ extensions $K/\mathbb{F}_{q}(t)$, for $\Gamma$ an arbitrary finite group. For $H$ a finite admissible $\Gamma$ group, let $G:=H\rtimes\Gamma$ and let $c\subset G$ be the set of elements of $G-\operatorname{id}$ which have the same order as their image in $\Gamma$. The reason that we restrict to $\Gamma=\mathbb{Z}/2\mathbb{Z}$ is because this is the only case for which $(G,c)$ is non-splitting, and hence this is the only case to which Theorem 1.4.5 will apply.
Let us now explain why we must have $\Gamma=\mathbb{Z}/2\mathbb{Z}$ for $(G,c)$ to be non-splitting. In order for $(G,c)$ to be non-splitting, it must also be the case that $(\Gamma,\Gamma-\operatorname{id})$ is non-splitting. In particular, $\Gamma-\operatorname{id}$ must consist of a single conjugacy class. If $\Gamma-\operatorname{id}$ is a single conjugacy class, its order must divide $|\Gamma|$, so $|\Gamma|-1$ must divide $|\Gamma|$. Supposing $|\Gamma|>1$, this can only happen when $|\Gamma|=2$.
Remark 5.5.1.
One may wonder why we restrict to quadratic fields $K/\mathbb{F}_{q}(t)$ unramified over infinity in Theorem 5.4.1, as opposed to also dealing with the case that these extensions are ramified over infinity. It seems likely that the same statement could be made for ramified extensions, following the proof of Theorem 5.4.1 in the case $n$ is odd. Another approach to the imaginary quadratic case could be by adapting the argument in the proof of [Woo19, Theorem 1.2]. We believe it would be interesting to address the imaginary quadratic case. This would also verify many additional cases of [Woo19, Conjecture 5.1], and also many cases of a function field version of [BW17, Conjecture 1.6].
Appendix A A topological model for the Hurwitz space bar construction
The goal of this appendix is to provide a good topological model for the two-sided bar construction $M\otimes_{\operatorname{Hur}^{G,c}_{+}}N$, where $M$ is a pointed set with a right $\operatorname{Hur}^{G,c}$-action and $N$ is a pointed set with a left $\operatorname{Hur}^{G,c}$ action. The end result that is used in the main body of the paper is Theorem A.4.1.
A.1. Background on Hurwitz spaces of braided sets
A natural setting in which our constructions go through is that of braided sets, sometimes also called “sets equipped with an invertible solution of the Yang–Baxter equation,” see [Dri90, LYZ00, Sol00]. We recall the definition:
Definition A.1.0.
A braided set is a set $X$ equipped with an automorphism $\sigma:X\times X\to X\times X$ satisfying the braid relation $\sigma_{1}\sigma_{2}\sigma_{1}=\sigma_{2}\sigma_{1}\sigma_{2}$, where $\sigma_{i}$ is the automorphism of $X\times X\times X$ obtained by applying $\sigma$ to factors $i$ and $i+1$ and the identity on the other factor.
Construction A.1.0 (Hurwitz spaces for braided sets).
Let $X$ be a braided set. Then there is a left action of the braid group on $X^{n}$, where the $i$th standard generator $\sigma_{i}$ of the braid group acts by applying $\sigma$ to the $i$ and $i+1$ factors of $X^{n}$ and acts as the identity on all other factors. The collection $(X^{n})_{n\in\mathbb{Z}_{\geq 0}}$ can be viewed as an $\mathbb{E}_{1}$-algebra in the category of local system on $\operatorname{Conf}=\cup_{n\geq 0}BB_{n}$, where the multiplication is given by concatenation. We define $\operatorname{Hur}^{X}$ to be the $\mathbb{E}_{1}$-algebra in $\operatorname{Spc}$ given by the étale cover of $\operatorname{Conf}$ arising from this local system, and use $\operatorname{Hur}^{X}_{n}$ to denote the part lying over $\operatorname{Conf}_{n}=BB_{n}$. This construction is natural in maps of braided sets, and there is a natural $\mathbb{E}_{1}$-algebra map $\operatorname{Hur}^{X}\to\operatorname{Conf}$. Explicitly, the underlying space of $\operatorname{Hur}^{X}_{n}$ is the homotopy quotient $(X^{n})_{hB_{n}}$.
Throughout the appendix, we use $X$ to refer to a braided set.
Example A.1.0.
The reader should keep in mind the following example: If $G$ is a group and $c$ is a union of conjugacy classes, then $c$ forms a braided set, where the twist operation $\sigma$ sends $(x,y)$ to $(y,y^{-1}xy)$. Then the Hurwitz space $\cup_{n}\operatorname{Hur}^{G,c}_{n,\mathbb{C}}$ as defined in § 1.4.2 agrees with $\operatorname{Hur}^{X}$ as defined in § A.1. More generally, if $c$ is a rack, the twist operation on $\operatorname{Hur}^{c}$ is given by sending $(x,y)\mapsto(y,y\triangleright x)$.
The next remark will not be used in what follows, but it may help explain why the category of braided sets is a natural setting to work in.
Remark A.1.0.
The category of $\mathbb{E}_{1}$-algebras in local systems on $\operatorname{Conf}$ with values in sets can be identified with the category of lax-monoidal functors from the free $\mathbb{E}_{2}$-monoidal category on a point into the category of sets. For a braided set $X$, viewing $\operatorname{Hur}^{X}$ as an object in this category via § A.1, gives an identification of the category of braided sets with the full subcategory consisting of monoidal functors.
A.2. A monoid model of Hurwitz spaces
We begin by defining topological spaces that model $\operatorname{Hur}^{X}$ and $\operatorname{Conf}$: First, we define a topological monoid, $\mathrm{conf}$, which is a model for the $\mathbb{E}_{1}$-algebra structure on $\operatorname{Conf}$.
Notation A.2.0.
Define the topological space $\mathrm{conf}$ as the space whose points are given by pairs $(t,x)$ where, $t\in\mathbb{R}_{\geq 0}$, and $x$ is a (possibly empty) configuration of finitely many distinct unordered points $(0,t)\times(0,1)$, not intersecting $[0,\frac{1}{2})\times[0,1]$ or $(t-\frac{1}{2},t]\times[0,1]$. The multiplication is given by concatenation of configurations and adding the elements $t$. We use $\mathrm{conf}$ to denote this topological monoid, and $\mathrm{conf}_{n}$ to denote the component with configuration of $n$ points. There is a map of topological monoids $t:\mathrm{conf}\to\mathbb{R}_{\geq 0}$ given by $(t,x)\mapsto t$. There is also a subset $\mathrm{ord}\subset\mathrm{conf}$ consisting of configurations such that the second coordinate of each point in $x$ is $\frac{1}{2}$. For each $n$, the intersection $\mathrm{ord}\cap\mathrm{conf}_{n}$ is contractible. We will use the inclusion $\mathrm{ord}_{n}:=\mathrm{ord}\cap\mathrm{conf}_{n}\subset\mathrm{conf}_{n}$ to view each component of $\mathrm{conf}$ as being equipped with a ‘base point’, i.e a fixed contractible subspace. We also use $\mathrm{confbig}$ to denote the variant of this topological monoid where we remove the condition that configurations not intersect $[0,\frac{1}{2})\times[0,1]$ and $(t-\frac{1}{2},t]\times[0,1]$, and use $\mathrm{bigord}$ and $\mathrm{bigord}_{n}$ to denote the corresponding subspaces. There is an inclusion of topological monoids $\mathrm{conf}\to\mathrm{confbig}$ which is clearly a homotopy equivalence.
Notation A.2.0.
We identify $\pi_{1}(\mathrm{conf}_{n},\mathrm{ord}_{n})$ with the braid group such that the elementary twist $\sigma_{i}$ (twisting strands $i$ and $i+1$) gets sent to the homotopy class of paths from $\mathrm{ord}_{n}$ that swaps the $i$th and $i+1$th element of the configuration via rotating them clockwise around each other by $\pi$.
Convention A.2.0.
By convention, the composition in the fundamental group/groupoid of a space is such that given two paths $\gamma$, $\gamma^{\prime}$, we have $[\gamma][\gamma^{\prime}]$ is the class of the path which is first $\gamma^{\prime}$ and then $\gamma$ (assuming they are composable). This convention is the one that is compatible with the orientation of the interval in that we view $\gamma^{\prime}$ as a morphism from its starting point to its end point.
We next define a topological monoid $\operatorname{hur}^{X}$ which models $\operatorname{Hur}^{X}$ as an $\mathbb{E}_{1}$-algebra.
Notation A.2.0.
Since $\operatorname{Hur}^{X}$, viewed as an $\mathbb{E}_{1}$-algebra, is obtained as a multiplicative finite cover of $\operatorname{Conf}$, we can construct a model $\operatorname{hur}^{X}$ for $\operatorname{Hur}^{X}$ as the corresponding multiplicative finite cover of $\mathrm{conf}$ in topological spaces. Namely, we first use the base points from the monoid map $\mathbb{N}\to\mathrm{conf}$ constructed in § A.2 to identify the fundamental group of $\mathrm{conf}$ with $B_{n}$, using the convention of § A.2. Then, $\operatorname{hur}^{X}_{n}$, as a cover of $\mathrm{conf}_{n}$, can be constructed as the quotient of $X^{n}\times\widetilde{\mathrm{conf}}_{n}$ by the braid group $B_{n}$, where $\widetilde{\mathrm{conf}}_{n}$ is the universal cover of $\mathrm{conf}_{n}$; More explicitly, a point of $\operatorname{hur}^{X}_{n}$ is given by a $B_{n}$-equivalence class of data $(x,t,\gamma,\alpha_{i})$, where $(x,t)\in\mathrm{conf}_{n}$, $\gamma$ is a homotopy class of paths from this point to $\mathrm{ord}$, and $\alpha_{i}$, for $1\leq i\leq n$, is an $n$ tuple of points in $X$, where $n=|x|$. Here the braid group action is such that $j$th standard generator of $B_{n}$, $\sigma_{j}$, sends $(x,t,\gamma,\alpha_{i})$ to $(x,t,\sigma_{j}\gamma,\sigma_{j}(\alpha_{i}))$ using § A.2 and § A.2 to get a left action of $B_{n}$ on such $\gamma$, and using the left action of the braid group on $X^{n}$ from § A.1.
To make $\operatorname{hur}^{X}:=\cup_{n\geq 0}\operatorname{hur}^{X}_{n}$ into a topological monoid, we use the multiplication of $\mathrm{conf}$, concatenate paths, and concatenate tuples of points in $X$. There is a map of topological monoids $\operatorname{hur}^{X}\to\mathrm{conf}$. Composing the time coordinate $t:\mathrm{conf}\to\mathbb{R}_{\geq 0}$ with this map, we obtain a composite map, (which we also call $t$,) $t:\operatorname{hur}^{X}\to\mathbb{R}_{\geq 0}$. This is a map of topological monoids. We also define a variant $\operatorname{hurbig}^{X}$ of $\operatorname{hur}^{X}$ by doing the same construction with $\mathrm{confbig}$ instead of $\mathrm{conf}$. We note $\operatorname{hurbig}^{X}$ is homotopy equivalent to $\operatorname{hur}^{X}$.
The key construction we use for Hurwitz spaces for braided sets is the following cutting construction. The idea for this construction is that if we start with a configuration in $[0,t^{\prime}]\times[0,1]$ which never contains points in the vertical line $t\times[0,1]$, we can cut it into two configurations by cutting along $t\times[0,1]$.
Construction A.2.0 (Cutting).
Let $t\in[0,\infty)$, and consider the open set $U(t)$ of $\operatorname{hurbig}^{X}$ consisting of $(x,t^{\prime},\gamma,\alpha_{i})$ such that $t^{\prime}>t$ and no points of the configuration $x$ lie on $t\times[0,1]$. Then there is a continuous map $U(t)\to\operatorname{hurbig}^{X}\times\operatorname{hurbig}^{X}$ which we call ‘cutting at $t$’, given as follows: We first cut the configuration on $[0,t^{\prime}]$ into two configurations $x^{\prime},x^{\prime\prime}$ on $[0,t]$ and $[0,t^{\prime}-t]\cong[t,t^{\prime}]$. Choose a homotopy class of paths $\phi$ from $x$ to $\mathrm{ord}$ with a representative that doesn’t cross $t\times[0,1]$. Using the braid group action, the pair $(x,t^{\prime},\gamma,\alpha_{i})$ is equivalent to $(x,t^{\prime},\phi,\beta_{i})$ for some $\beta_{i}$ obtained by acting on the tuple $(x_{i})$ via the element of the braid group given by the class of $\phi\gamma^{-1}$. We can now similarly cut $\phi$ into homotopy classes of paths $\phi^{\prime},\phi^{\prime\prime}$ from the configurations $x^{\prime},x^{\prime\prime}$ to the respective base points, and cut the sequence $\beta_{i}$ into $\alpha_{j}^{\prime},\alpha_{k}^{\prime\prime}$ by unconcatenating it into two sequences corresponding to the number of points in the configuration in $[0,t]\times[0,1]$ and $[t,t^{\prime}]\times[0,1]$ respectively.
Then the cutting map sends $(x,t^{\prime},\gamma,\alpha_{i})$ to $(x^{\prime},t,\phi^{\prime},\alpha^{\prime}_{j})$ and $(x^{\prime\prime},t^{\prime}-t,\phi^{\prime\prime},\alpha^{\prime\prime}_{k})$. Since the concatenation map $X^{a}\times X^{b}\to X^{a+b}$ is $B_{a}\times B_{b}$-equivariant, this doesn’t depend on the choice of $\phi$.
A.3. A scanning map
Next, we produce well behaved topological models for the source and target of § 4.5. The following definition will be the key ingredient in constructing the models:
Notation A.3.0.
Let $M$ be a set with a right action of $\operatorname{Hur}^{X}$ and $N$ be a set with a left action of $\operatorname{Hur}^{X}$. Consider the space $B[M,\operatorname{Hur}^{X},N]$ consisting of points which are of the form
(A.1) $(a,b,y)$
where $a\in M,b\in N$ and $y=(x,1,\gamma,x_{i})\in\operatorname{hurbig}^{X}$, following notation from § A.2. The topology on $B[M,\operatorname{Hur}^{X},N]$ has a basis given as follows. Consider the following data:
  1. (a)
    Numbers $d_{1},d_{2}\in(0,1)$.
  2. (b)
    A finite collection of pairwise disjoint open balls $U_{1},\ldots,U_{n}$ in contained in the interior of $[d_{1},d_{2}]\times[0,1]$.
  3. (c)
    A homotopy class of paths $\phi$ from the center of the balls $U_{i}$, viewed as an element of $\mathrm{confbig}_{n}$, to the base point.
  4. (d)
    Elements $\alpha_{1},\ldots,\alpha_{n}\in X$.
  5. (e)
    Elements $a^{\prime}\in M,b^{\prime}\in N$.
We next define subsets $\mathfrak{B}(d_{1},d_{2},U_{i},\phi,\alpha_{i},a^{\prime},b^{\prime})$ which form a basis of the topology on $B[M,\operatorname{Hur}^{X},N]$. A point of the form (A.1) lies in $\mathfrak{B}(d_{1},d_{2},U_{i},\phi,\alpha_{i},a^{\prime},b^{\prime})$ if the following conditions hold.
  1. (1)
    None of the points in $x$ lie in $[d_{1},d_{2}]\times[0,1]-\cup_{1}^{n}U_{i}$, and there is a unique point from $x$ in each $U_{i}$.
  2. (2)
    Cutting the element of $\operatorname{hurbig}^{X}$ twice to restrict it to the interval $[d_{1},d_{2}]$ yields a point $y^{\prime}\in\operatorname{hurbig}^{X}$ (see § A.2). Then, using the homotopy class of $\phi$, the corresponding tuple of elements in $X$ associated to $y^{\prime}$ is $\alpha_{1},\ldots,\alpha_{n}$.
  3. (3)
    Define $y_{1}\in\operatorname{hurbig}^{X}$ to be the element of $\operatorname{hur}^{X}$ obtained by cutting and restricting to the interval $[0,d_{1}]$, and let $y_{2}$ be the element similarly obtained by using the interval $[d_{2},1]$. We then require that $ay_{1}=a^{\prime}$ and $y_{2}b=b^{\prime}$.
Remark A.3.0.
[…]
We next construct a homotopy equivalence showing that the topological space $B[M,\operatorname{Hur}^{X},N]$ models the two-sided bar construction $M\otimes_{\operatorname{hur}^{X}}N$, which we define next. Our proof is an adaptation of the proof of [BDPW23, Theorem 5.2.3]. Indeed, the case $M=N=X=*$ corresponds to the result there.
Notation A.3.0.
Let $H$ be a topological monoid and let $M$ be a right module for $H$ and let $N$ be a left module for $H$. View $\Delta^{n}$ as the collection of tuples $(y_{0},y_{1},\ldots,y_{n})$ with $\sum_{i=0}^{n}y_{i}=1$. We will describe elements of $M\times N\times\Delta^{n}\times H^{n}$ via the notation
(A.2) $(a,b,(y_{0},\ldots,y_{n}),(x_{1},\ldots,x_{n})).$
We define the two-sided bar construction, $M\otimes_{H}N$, to be the quotient of $\coprod_{n\geq 0}M\times N\times\Delta^{n}\times H^{n}$, by the following equivalence relations:
  1. (1)
    $(a,b,(0,y_{1},\ldots,y_{n}),(x_{1},x_{2},\ldots,x_{n}))\sim(ax_{1},b,(y_{1},\ldots,y_{n}),(x_{2},\ldots,x_{n}))$,
  2. (2)
    $(a,b,(y_{0},\ldots,y_{n-1},0),(x_{1},\ldots,x_{n-1},x_{n}))\sim(a,x_{n}b,(y_{0},\ldots,y_{n-1}),(x_{1},\ldots,x_{n-1}))$,
  3. (3)
    $(a,b,(y_{0},\ldots,y_{i-1},0,y_{i+1},\ldots,y_{n}),(x_{1},\ldots,x_{i-1},x_{i},x_{i+1},\ldots,x_{n}))$
    $\sim(a,b,(y_{0},\ldots,y_{i-1},y_{i+1},\ldots,y_{n}),(x_{1},\ldots,x_{i-1},x_{i}x_{i+1},\ldots,x_{n}))$,
  4. (4)
    $(a,b,(y_{0},\ldots,y_{i-1},y_{i},y_{i+1},\ldots,y_{n}),(x_{1},\ldots,x_{i-1},e,x_{i+1},\ldots,x_{n}))$
    $\sim(a,b,(y_{0},\ldots,y_{i-1}+y_{i},y_{i+1},\ldots,y_{n}),(x_{1},\ldots,x_{i-1},x_{i+1},\ldots,x_{n}))$, where $e$ denotes the identity of the monoid $H$.
Lemma A.3.0.
Let $M$ be a set with a right action of $\operatorname{hur}^{X}$ and let $N$ be a set with a left action of $\operatorname{hur}^{X}$. Then there is a weak homotopy equivalence $\sigma:M\otimes_{\operatorname{hur}^{X}}N\to B[M,\operatorname{Hur}^{X},N]$, natural in $X,M,N$.
Proof.
Recall that $M\otimes_{\operatorname{hur}^{X}}N$ denotes the two-sided bar construction, as defined in § A.3. In particular, since $\operatorname{hur}^{X}$ models $\operatorname{Hur}^{X}$ as an $\mathbb{E}_{1}$-algebra, $M\otimes_{\operatorname{hur}^{X}}N$ models the two-sided bar construction $M\otimes_{\operatorname{Hur}^{X}}N$.
We produce a map $\sigma:M\otimes_{\operatorname{hur}^{X}}N\to B[M,\operatorname{Hur}^{X},N]$, as follows: Given data as in (A.2) for $H=\operatorname{hur}^{X}$, and using notation from § A.2, let $t_{i}:=t(x_{i})\in\mathbb{R}_{>0}$ denote the time associated to $x_{i}\in\operatorname{hur}^{X}$. We multiply $x_{1},\ldots,x_{n}$ to give an element $x\in\operatorname{hur}^{X}$ with $t:=t(x)=\sum_{i=1}^{n}t_{i}$. First, extend the interval on which $x$ is defined from $[0,t]\times[0,1]$ to $[-\frac{1}{2},t+\frac{1}{2}]\times[0,1]$. We then define $t^{\prime}:=\sum_{i=1}^{n}y_{i}\cdot(\sum_{j=1}^{i}t_{j})$. At this point, it will be useful to recall the cutting construction from § A.2. Choose $\epsilon>0$ small enough so that there are no points of the configuration $x$ in $(t^{\prime}-\frac{1}{2},t^{\prime}-\frac{1}{2}+\epsilon]\times[0,1]$ and $[t^{\prime}+\frac{1}{2}-\epsilon,t^{\prime}+\frac{1}{2})\times[0,1]$. We can then cut $x$ into three pieces, $w,x^{\prime\prime},z\in\operatorname{hurbig}^{X}$ coming from the intervals $[-\frac{1}{2},t^{\prime}-\frac{1}{2}+\epsilon]$, $[t^{\prime}-\frac{1}{2}+\epsilon,t^{\prime}+\frac{1}{2}-\epsilon],[t^{\prime}+\frac{1}{2}-\epsilon,t+\frac{1}{2}]$. Extend $x^{\prime\prime}$ to be of length $1$ by extending the length of the interval on both ends by $\epsilon$, and let $x^{\prime}$ denote the resulting element of $\operatorname{hurbig}^{X}$. Note that $x^{\prime}$ doesn’t depend on $\epsilon$, and neither does the class of $w,z$ in $\pi_{0}\operatorname{Hur}^{X}$.
The then define the map $\sigma$ to sends data as in (A.2) to the point $(aw,zb,x^{\prime})$, in $B[M,\operatorname{Hur}^{X},N]$, with notation as in (A.1). To see this is well defined, one must check that this glues along the identifications (1)-(4) in the two sided bar construction given in § A.3. We omit this verification.
To verify that $\sigma$ is continuous, it suffices to do check continuity on $M\times N\times\Delta^{n}\times(\operatorname{hur}^{X})^{n}$. We now carry out this straightforward verification. We choose a basic open set of the form $\mathfrak{B}(d_{1},d_{2},U_{i},\phi,\alpha_{i},a^{\prime},b^{\prime})$. We wish to verify its preimage under $\sigma$ is also open. If $(a,b,(y_{0},\ldots,y_{n}),(x_{1},\ldots,x_{n}))$ is a point in the preimage, then in a small neighborhood, $(1)$ and $(2)$ of § A.3 are clearly still satisfied. The path components of $w$ and $z$ are also unchanged in a small neighborhood, so condition $(3)$ is satisfied as well.
Having shown $\sigma$ is continuous, we next claim that $\sigma$ is surjective on path components. Every point in $B[M,\operatorname{Hur}^{X},N]$ has a path to a point where the configuration on $[0,1]\times[0,1]$ is empty, by linearly pushing the configuration towards $0\times[0,1]$ on the boundary, and multiplying once points collide with the boundary. A point with an empty configuration is determined by $(a,b)\in M\times N$. This point is the image of the point $(a,b,(1),())\in M\otimes_{\operatorname{hur}^{X}}N$, showing the claimed surjectivity on path components. We view these points as above with empty configuration as giving a copy of $M\times N\subset M\otimes_{\operatorname{hur}^{X}}N$.
Recall we are aiming to show $\sigma$ is a homotopy equivalence. To do this, we claim it suffices to show the following two statements.
  1. (i)
    For every $v\geq 1$, given a map of pairs $f:(D^{v},S^{v-1})\to(B[M,\operatorname{hur}^{X},N],M\times N)$, there is a lift, up to homotopy, to a map of pairs $g:(D^{v},S^{v-1})\to(M\otimes_{\operatorname{hur}^{X}}N,M\times N)$ such that $\sigma\circ g$ is homotopic to $f$.
  2. (ii)
    For every $v\geq 1$ and a map of pairs $g:(S^{v},*)\to(M\otimes_{\operatorname{hur}^{X}}N,M\times N)$ such that $\sigma\circ g$ is nullhomotopic, then $g$ is nullhomotopic.
We now explain why $(i)$ and $(ii)$ above imply $\sigma$ is a homotopy equivalence. First, $(i)$ in the case $v=1$, shows in particular that the map $\sigma$ on $\pi_{0}$ is injective, since it shows that the equivalence relation describing $\pi_{0}$ as a quotient of $M\times N$ is the same for the source and target. Since we showed above that $\sigma$ is a surjection on $\pi_{0}$, it will follow that $\sigma$ is an isomorphism on $\pi_{0}$. Then $(i)$ also shows that on each component, the map on $\pi_{v}$ is surjective for each $v\geq 1$, and $(ii)$ shows $\pi_{v}$ is injective.
We turn to proving $(i)$, and assume that we have a map $f$ as given in the statement of $(i)$. We set up notation to define the map $g$ described in $(i)$. Using compactness, we can find a finite open cover $D^{v}=\cup_{\lambda\in\Lambda}U_{\lambda}$ such that there is an $r>0$ and for each $\lambda\in\Lambda$, there is a $t_{\lambda}\in(0,1)$ such that for all $u\in U_{\lambda}$, the configuration associated to $f(u)$ has no points in $(t_{\lambda}-r,t_{\lambda}+r)\times[0,1]$. Choose a partition of unity $w_{\lambda}$ subordinate to the cover $U_{\lambda}$. Let $u$ be a point in $U_{\lambda_{0}}\cap\cdots\cap U_{\lambda_{n}}$, and not in $U_{\lambda}$ for $\lambda\in\Lambda-\{\lambda_{1},\ldots,\lambda_{n}\}$. Assume that $t_{\lambda_{0}}<t_{\lambda_{1}}<\cdots<t_{\lambda_{n}}$ (which can always be achieved by decreasing $r$ and making all $t_{\lambda}$ distinct). Then we consider the $n$-tuple of elements $x_{1},\ldots,x_{n}\in\operatorname{hur}^{X}$, where $x_{i}$ is obtained from the element of $\operatorname{hur}$ associated to $f(u)$ by cutting along $t_{\lambda_{i-1}}\times[0,1]$ and $t_{\lambda_{i}}\times[0,1]$, taking the configuration with points lying in $[t_{\lambda_{i-1}},t_{\lambda_{i}}]\times[0,1]$, and rescaling the intervals by $\frac{1}{2r}$ so that the time function $t$ applied to this new configuration is $(t_{\lambda_{i}}-t_{\lambda_{i-1}})\frac{1}{2r}$. Define $u^{\prime}$ and $u^{\prime\prime}\in\operatorname{hurbig}^{X}$ to be the elements obtained respectively by cutting to restrict to the intervals $[0,t_{\lambda_{0}}]$ and $[t_{\lambda_{n}},1]$ respectively. We define the map $g:D^{v}\to M\otimes_{\operatorname{hur}^{X}}N$ by sending $u$ to the tuple
$(au^{\prime},u^{\prime\prime}b,(w_{\lambda_{0}(u)},\ldots,w_{\lambda_{n}(u)}),(x_{1},\ldots,x_{n})).$
We will show there is a homotopy between $\sigma\circ g$ and $f$, which will verify $(i)$. Choose the open cover $\{U_{\lambda}\}_{\lambda\in\Lambda}$ of $D^{v}$ so that image of each component the boundary $S^{v-1}$ is contained in exactly one $U_{\lambda_{i}}$. (When $v>1$, the boundary is connected, so this is the same as saying the boundary $S^{v-1}$ is contained in exactly one $U_{\lambda_{i}}$.) Then, $g$ is a map of pairs because each component of the boundary is sent to a single point, as $M\times N$ is a discrete set.
We now check continuity of $g$ at a point $u\in D^{v}$. Let $j,k$ be such that $w_{\lambda_{i}}(u)=0$ for $i<j$ and $i>k$ but not for $i=j$ or $i=k$. Any neighborhood of $g(u)$, contains an open set $W$ consisting of data of the form $(a^{\prime},b^{\prime},(y_{0},\ldots,y_{n}),(z_{1},\ldots,z_{n}))$ where $|y_{i}-w_{\lambda_{i}}(u)|<\epsilon$ for some $\epsilon>0$ small enough such that this guarantees that $y_{j},y_{k}\neq 0$, and such that
  • ($\star$)
    $a^{\prime}z_{1}\cdots z_{j-1}=au^{\prime}x_{1}\cdots x_{j-1}$ and $z_{k+1}\cdots z_{n}b^{\prime}=x_{k+1}\cdots x_{n}u^{\prime\prime}b$.
In order to show $g$ is continuous at $u$, it is then enough that for any such open set $W$, $g^{-1}(W)$ contains an open neighborhood of $u$ in $D^{v}$.
Choose a basic open set $\mathfrak{B}(d_{1},d_{2},U_{i},\phi,\alpha_{i},a^{\prime},b^{\prime})$ containing $f(u)$ and a small open neighborhood $V$ containing $u$ with $V\subset f^{-1}(\mathfrak{B}(d_{1},d_{2},U_{i},\phi,\alpha_{i},a^{\prime},b^{\prime}))\cap U_{\lambda_{0}}\cap\cdots\cap U_{\lambda_{n}}$. We may shrink $\mathfrak{B}(d_{1},d_{2},U_{i},\phi,\alpha_{i},a^{\prime},b^{\prime})$, to assume $d_{1}\leq t_{\lambda_{0}},d_{2}\geq t_{\lambda_{n}}$, where, here, shrinking this neighborhood may involve increasing the number of $U_{\lambda}$ so that there is one such open containing each point of the configuration $x$, appearing in the data for $f(u)$, in $(d_{1},d_{2})\times[0,1]$. By further shrinking $V$, we may assume the functions $w_{\lambda}$ for $\lambda\in\Lambda-\{\lambda_{0},\ldots,\lambda_{n}\}$ vanish on $V$. Therefore, using identifications $(1)-(3)$ of the bar construction, for the purposes of proving continuity, we may assume our open cover consists solely of $U_{\lambda_{0}},\ldots,U_{\lambda_{n}}$, i.e, we can ignore the other $U_{\lambda}$. At this point, condition (3) of § A.3 enables us to further shrink $V$ so that ($\star$) holds for all $g(v)$ with $v\in V$. We have therefore found an open set in $D^{v}$ containing $u$ and contained in $g^{-1}(W)$ for $W$ an open set containing $g(u)$. This proves that $g$ is continuous.
For a point $z\in D^{v+1}$, $\sigma(g(z))$ is obtained from $f(z)$ by zooming in on the configuration on the interval around $\sum_{i=0}^{n}w_{\lambda_{i}}t_{\lambda_{i}}$ of radius $r$, rescaling it to be in the interval $[0,1]$, and then multiplying the element of $\operatorname{hurbig}$ obtained by cutting to the left of this interval with $a$ and multiplying the element of $\operatorname{hurbig}$ obtained by cutting to the right with $b$. There is an evident continuous homotopy from $f$ to $\sigma\circ g$, given by linearly rescaling around this interval, and letting points act on the boundary as they reach it. This gives us a homotopy between $f$ and $\sigma\circ g$, which is a homotopy of pairs because of our assumption about $g$ applied to the boundary sphere (each component of which is sent to a single point). Thus we have verified $(i)$.
To prove $(ii)$, given a map of pairs $g:(S^{v},*)\to(M\otimes_{\operatorname{hur}^{X}}N,M\times N)$ and a nullhomotopy of $\sigma\circ g$, the same construction used above for $(i)$ lifts the nullhomotopy of $\sigma\circ g$ to a pointed map $h:D^{v+1}\to M\otimes_{\operatorname{hur}^{X}}N$. It remains to show that $h|_{S^{v}}$ is homotopic to $g$ as a pointed map. Here we follow the last paragraph of the proof of [BDPW23, Theorem 5.2.3]. Namely, given $u\in S^{v}$, suppose $u$ is in $U_{\lambda_{0}}\cap\cdots\cap U_{\lambda_{n}}$ and not in $U_{\lambda}$ for $\lambda\in\Lambda-\{\lambda_{0},\ldots,\lambda_{n}\}$. Associated to $g(u)$ is a sequence of $\kappa$ elements in $\operatorname{hur}^{X}$, for some integer $\kappa$. We can first linearly stretch these $\kappa$ elements each by a factor of $\frac{1}{2r}$, and then use the identifications of (3) of the 2-sided bar construction § A.3 to make $n$ cuts in the places specified by $t_{\lambda_{i}}$ in $\sigma\circ g$. Let $\tilde{g}$ be the function of $u$ obtained this way, so that $\tilde{g}$ is homotopic to $g$. Then, $\tilde{g}$ agrees with $h|_{S^{v}}$ except for the weights defining a point in $\Delta^{n}$. Linearly changing these weights then gives a continuous homotopy between $\tilde{g}$ and $h|_{S^{v}}$. Thus $g$ is homotopic to $h|_{S^{v}}$. ∎
A.4. A simple topological model
We will carry out a scanning argument to obtain a description of the homotopy type of $B[M,\operatorname{Hur}^{X},N]$ in terms of explicit quotient spaces, which we define next. This identification is verified in the unpointed case in § A.4, and in the pointed case in Theorem A.4.1. In the next definition, we have two somewhat complicated relations $(i)$ and $(ii)$. We must impose these relations so that the proof of § A.4 goes through, the key idea being depicted in Figure 3.
Definition A.4.0.
Let $M$ be a set with a right action of $\operatorname{Hur}^{X}$ and let $N$ be a set with a left action of $\operatorname{Hur}^{X}$. Let $10<\epsilon<$ 1. We define $Q_{\epsilon}[M,\operatorname{Hur}^{X},N]$ to be the set consisting of triples $(m,n,x,l:x\to X)$ with $m\in M,n\in N,$ $x=\{p_{1},\ldots,p_{k}\}$ a finite subset of $[0,1]\times[\epsilon,1-\epsilon]$ (with first projection $\pi_{1}:[0,1]\times[0,1]\to[0,1]$ and second projection $\pi_{2}:[0,1]\times[0,1]\to[0,1]$) such that for each $i\neq j$, the distance between $\pi_{2}(p_{i})$ and $\pi_{2}(p_{j})$ is at least $\epsilon$, and $l:x\to X$ is a map of sets.
We define $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ to be the quotient space of $Q_{\epsilon}[M,\operatorname{Hur}^{X},N]$ under the equivalence relation generated by the following two types of equivalences.
  • $(i)$
    Suppose $x=\{p_{1},\ldots,p_{v},\ldots,p_{k}\}$ with $\pi_{2}(p_{1})>\cdots>\pi_{2}(p_{k})$. If $\pi_{1}(p_{v})=0$, so $p_{r}$ lies on the left boundary, the data ($m,n,x,l:x\to X)$ is equivalent to $(m^{\prime},n,x^{\prime}:=\{p_{1},\ldots,p_{r-1},p_{r+1},\ldots,p_{k}\},l^{\prime}:x^{\prime}\to X)$, defined as follows: write $(y_{1},\ldots,y_{r}):=\sigma_{1}\sigma_{2}\cdots\sigma_{r-1}(l(p_{1}),\ldots,l(p_{r})).$ Then $m^{\prime}=my_{1}$, and $l^{\prime}(p_{i}):=y_{i}$ for $1\leq i<r$, $l^{\prime}(p_{i}):=l(p_{i})$ for $i>r$.
  • $(ii)$
    Suppose $x=\{p_{1},\ldots,p_{r},\ldots,p_{k}\}$ with $\pi_{2}(p_{1})>\cdots>\pi_{2}(p_{k})$. If $\pi_{1}(p_{r})=1$, so $p_{r}$ lies on the right boundary, the data ($m,n,x,l:x\to X)$ is equivalent to $(m,n^{\prime},x^{\prime}:=\{p_{1},\ldots,p_{r-1},p_{r+1},\ldots,p_{k}\},l^{\prime}:x^{\prime}\to X)$, defined as follows: write $(y_{r},\ldots,y_{k}):=\sigma_{k-r-1}\cdots\sigma_{1}(l(p_{r}),\ldots,l(p_{k})).$ Then $n^{\prime}=y_{k}n$, and $l^{\prime}(p_{i}):=l(p_{i})$ for $i<r$, $l^{\prime}(p_{i})=y_{i-1}$ for $i>r$.
Remark A.4.0.
In the case that the braided set $X$ comes from a rack $c$, the equivalences in § A.4 can be described more simply as follows.
  • $(i)$
    Suppose $x=\{p_{1},\ldots,p_{r},\ldots,p_{k}\}$ with $\pi_{2}(p_{1})>\cdots>\pi_{2}(p_{k})$. If $\pi_{1}(p_{r})=0$, so $p_{r}$ lies on the left boundary, the data ($m,n,x,l:x\to X)$ is equivalent to $(m^{\prime},n,x^{\prime}:=\{p_{1},\ldots,p_{r-1},p_{r+1},\ldots,p_{k}\},l^{\prime}:x^{\prime}\to X)$, where $m^{\prime}=ml(p_{r})$, $l^{\prime}(p_{i}):=l(p_{r})\triangleright l(p_{i})$ for $1\leq i<r$, $l^{\prime}(p_{i}):=l(p_{i})$ for $i>r$.
  • $(ii)$
    Suppose $x=\{p_{1},\ldots,p_{r},\ldots,p_{k}\}$ with $\pi_{2}(p_{1})>\cdots>\pi_{2}(p_{k})$. If $\pi_{1}(p_{r})=1$, so $p_{r}$ lies on the right boundary, the data ($m,n,x,l:x\to X)$ is equivalent to $(m,n^{\prime},x^{\prime}:=\{p_{1},\ldots,p_{r-1},p_{r+1},\ldots,p_{k}\},l^{\prime}:x^{\prime}\to X)$, where $n^{\prime}=((l(p_{r+1})\cdots l(p_{k}))\triangleright l(p_{r}))n$, $l^{\prime}(p_{i}):=l(p_{i})$ for $i\neq r$.
For technical reasons, we will want to work with inductive systems of topological spaces $(X_{\epsilon})_{\epsilon>0}$, which we view as presenting the colimit of their homotopy types. For example, we will consider $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ as an ind-system as $\epsilon$ approaches 00.
Definition A.4.0.
Let $L:\mathrm{Top}\to\operatorname{Spc}$ and $L:\mathrm{Top}_{*}\to\operatorname{Spc}_{*}$ denote the functors of $\infty$-categories sending a (pointed) topological space to its weak homotopy type. Given a (pointed) map $f:X\to X^{\prime}$ in $\operatorname{Ind}(\mathrm{Top})$ or $\operatorname{Ind}(\mathrm{Top}_{*})$, we say that it is an ind-weak equivalence if $\operatorname{colim}Lf$ is an equivalence. Similarly, we say that $f$ is an ind-weak homology equivalence if $\operatorname{colim}Lf$ is a homology equivalence.
The following is an ind-variant of a well known lemma about weak homotopy equivalences. In what follows, we use the smash product $\wedge$, the definition of which is in § 3.1.1.
Lemma A.4.0.
Let $P$ be a filtered poset, and let $X^{\prime}_{\bullet}\hookrightarrow X_{\bullet},\bullet\in P$ be an injection of $P$-indexed diagrams of path connected pointed topological spaces (not necessarily with the subspace topology) such that for every pointed inclusion of compact spaces $K^{\prime}\hookrightarrow K$, and commutative diagram
[…]
there is a $\beta>\alpha$ and a pointed homotopy $H:K\wedge[0,1]_{+}\to X_{\beta}$ with the following two properties.
  1. (1)
[…]
  2. (2)
    The restriction of $H$ to the point $1\in[0,1]_{+}$ defines a map $K\to X_{\beta}$, which we assume factors continuously through $X^{\prime}_{\beta}$.
Then $X^{\prime}_{\bullet}\to X_{\bullet}$ is an ind-weak homotopy equivalence.
Proof.
Since $\pi_{i}(-)=\pi_{i-1}(\Omega-)$ and homotopy sets and based loop spaces commute with filtered colimits in $\operatorname{Spc}$, it suffices by induction on $i$ to show that $\Omega X^{\prime}_{\bullet}\hookrightarrow{\Omega X_{\bullet}}$ satisfies the same hypotheses $(1)$ and $(2)$ as in the statement that $X^{\prime}_{\bullet}\to X_{\bullet}$ satisfies (except that $\Omega X^{\prime}_{\bullet}$ and $\Omega X_{\bullet}$ need not be connected), and moreover that $\Omega X^{\prime}_{\bullet}\hookrightarrow{\Omega X_{\bullet}}$ is an ind-equivalence on $\pi_{0}$.
To see that the map $\Omega X^{\prime}_{\bullet}\hookrightarrow{\Omega X_{\bullet}}$ satisfies $(1)$ and $(2)$, note that giving a pointed map $K\to\Omega X_{\alpha}$ is the same as giving a map $K\wedge S^{1}\to X_{\alpha}$, and similarly for $K^{\prime}$. Here, we implicitly choose a fixed point in $S^{1}$ and view $S^{1}$ as a pointed space to make sense of the wedge product. Applying the hypotheses to $K\wedge S^{1}$ in place of $K$, we get an pointed map $K\wedge S^{1}\wedge[0,1]_{+}\to X_{\beta}$, which is the same as a map $K\wedge[0,1]_{+}\to\Omega X_{\beta}$ satisfying the desired properties $(1)$ and $(2)$.
We finish the proof by showing that show that $\operatorname{colim}_{\alpha\in P}\pi_{0}\Omega X^{\prime}_{\alpha}\to\operatorname{colim}_{\alpha\in P}\pi_{0}\Omega X_{\alpha}$ is a bijection. We first check it is surjective. Given a point $*\in\Omega X_{\alpha}$, taking $K$ to be the compact space $*_{+}$ and $K^{\prime}\subset K$ the base point, the hypotheses applied to the pointed map $*_{+}\to\Omega X_{\alpha}$ shows surjectivity. For injectivity, if $[0,1]_{+}\to\Omega X_{\alpha}$ is a path between two points with the end points factoring through $\Omega X^{\prime}_{\alpha}$, then applying the hypothesis to this map, we get a map $([0,1]\times[0,1])_{+}\to\Omega X_{\beta}$, and the part of the boundary $\{0,1\}\times[0,1]\cup[0,1]\times\{1\}$ yields a path in $\Omega X^{\prime}_{\beta}$ between the two points. ∎
A useful construction we will use to study the space $B[M,\operatorname{hur}^{X},N]$ is a flow that pushes configurations outwards, which we define next.
Construction A.4.0.
Choose a smooth function $f:[0,1]\to\mathbb{R}$ such that $f([0,\frac{1}{2}))<0$ and $f(1-x)=-f(x)$. In particular, $f(1/2)=0$. Consider the vector field $f(x)\frac{\partial}{\partial x}$ on $[0,1]\times[0,1]$, where $x$ is the first coordinate. We can then form a continuous flow $\phi_{t}:B[M,\operatorname{Hur}^{X},N]\times[0,\infty)\to B[M,\operatorname{Hur}^{X},N]$ by moving points in a configuration along the flow, and having them act on the elements $(m,n)$ in $M\times N$ when they reach the boundary. We can define such a flow (which we give the same name) on $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ via the same procedure.
The following proposition is key in relating $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ to $B[M,\operatorname{Hur}^{X},N]$. This proposition characterizes the image of the natural map from the first to the second.
Proposition A.4.0.
Let $M$ be a set with a right action of $\operatorname{Hur}^{X}$ and $N$ be a set with a left action of $\operatorname{Hur}^{X}$. Let $\pi_{2}:[0,1]\times[0,1]\to[0,1]$ be the second projection. There is a natural continuous injection $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]\to B[M,\operatorname{Hur}^{X},N]$ whose image consists of those points $(a,b,y)$ where $y:=(x:=\{p_{1},\ldots,p_{k}\},1,\gamma,x_{i})$ as in § A.3 with the property that $\epsilon\leq\pi_{i}(p_{j})\leq 1-\epsilon$ for $1\leq j\leq k$ and any two distinct elements in $\{\pi_{2}(p_{1}),\cdots,\pi_{2}(p_{k})\}$ differ by at least $\epsilon$.
Moreover, a continuous map $K\to B[M,\operatorname{Hur}^{X},N]$ lifts to $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ if its image in the image of finitely many connected components of $Q_{\epsilon}[M,\operatorname{Hur}^{X},N]$.
Proof.
To shorten notation, in this proof only, we will use $Q$ to denote $Q_{\epsilon}[M,\operatorname{Hur}^{X},N]$ and $\overline{Q}$ to denote $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$. We will produce a continuous map
$Q\to B[M,\operatorname{Hur}^{X},N]$
which induces a map $\overline{Q}\to B[M,\operatorname{Hur}^{X},N]$ satisfying the claimed properties.
We define the interior $Q^{\circ}$ to be the dense subset of $Q$ on which the configurations are contained in $(0,1)\times[\epsilon,1-\epsilon]$. In other words, a point of $Q^{\circ}$ is given by $m\in M,n\in N$, such a configuration $x$ and a function $l:x\to X$. We define a continuous function $g:Q^{\circ}\to B[M,\operatorname{Hur}^{X},N]$ as we describe next.
We first continuously choose a homotopy class of paths $\gamma$ from the configuration to $\mathrm{bigord}$ as follows: given a configuration, because the projection $\pi_{2}$ to the second coordinate is injective, we can linearly move all points so that their first coordinate is $\frac{1}{2}$. Then, we can rotate the configuration in $[0,1]\times[0,1]$ counterclockwise by $\frac{\pi}{2}$ so that the second coordinate of all the points is $\frac{1}{2}$, so that the path ends in $\mathrm{bigord}$. This ends our description of $\gamma$.
Let $x:=\{p_{1},\ldots,p_{k}\}$ with $\pi_{2}(p_{1})>\cdots>\pi_{2}(p_{k})$ and define $x_{i}:=l(p_{i})$. The map $g$ then sends the above point $(m,n,x,l:x\to X)\in Q$ to the point $(m,n,y)\in B[M,\operatorname{Hur}^{X},N]$, where $y\in\operatorname{hurbig}^{X}$ is given by $(x,1,\gamma,x_{i})$. We omit the straightforward verification that this map bijects $Q^{\circ}$ onto the image stated in the proposition statement.
We claim the above map $Q^{\circ}\to B[M,\operatorname{Hur}^{X},N]$ continuously extends to a map $Q\to B[M,\operatorname{Hur}^{X},N]$. To see this, we can produce a continuous inward flow $\varphi_{-t}$ on $Q$ by pushing points inwards along the vector field $-f(x)\frac{\partial}{\partial x}$ as in § A.4. For positive $t$, this flow lands inside $Q^{\circ}$, so we can continuously extend the function $g$ to a map $\tilde{g}:Q\to B[M,\operatorname{Hur}^{X},N]$ via the formula $\phi_{t}\circ g\circ\varphi_{-t}$ for any $t>0$, with $\phi_{t}$ as in § A.4. It is easy to see that the underlying configuration associated to the map $\tilde{g}$ is similar to the configuration of the original point, with the only difference being that all of the points on the boundary of $[0,1]\times[0,1]$ are removed.
Since the image of $g$ was closed, and $Q^{\circ}$ is dense in $Q$, every element of $Q$ is sent to the same point as some element of $Q^{\circ}$. To show that the equivalence relation induced by $\tilde{g}$ is exactly determined by $(i)$ and $(ii)$ from § A.4, it is then enough to show that $(i)$ and $(ii)$ are satisfied, because one can repeatedly use $(i)$ and $(ii)$ to show that every point is equivalent to a point in the image of $Q^{\circ}$, and the map $g$ is a bijection, so this accounts for the entire equivalence relation.
To show relations $(i)$ and $(ii)$, we can reduce to the case where there is exactly one point on the boundary, since if we prove the relations in this case, by continuity, the general case will follow from continuity by perturbing the other points on the boundary and having them approach the boundary.
We now establish relation $(ii)$ from § A.4 in this case, as $(i)$ follows from a similar argument. Suppose that we have a point $w\in Q$ where $y$ is the only point in the configuration associated to $w$ on the boundary of $[0,1]\times[0,1]$. Moreover, assume $y$ lies on $1\times[0,1]$. We need to compute the limit of $\phi_{s}\circ g\circ\varphi_{-t}(w)$ as $s$ approaches $t$ from below. For the configuration associated to point $\varphi_{-t}(w)$, we can choose some $\delta<1$ so that the point $y^{\prime}$ to which $y$ flows in in $\varphi_{-t}(w)$ has $\pi_{1}(y^{\prime})>\delta$ (where $\pi_{1}(y^{\prime})$ denotes the first coordinate of $y^{\prime}$) and is the only point of the configuration with first coordinate $\geq\delta$. Let $[\gamma]$ be the homotopy class of paths from the configuration in $\varphi_{-t}(w)$ to $\mathrm{bigord}$ that is used in defining the map $g$. We can choose a different homotopy class of paths by first cutting up $[0,1]\times[0,1]$ into $[0,\delta]\times[0,1]$ and $[\delta,1]\times[0,1]$, doing the construction as in the map $g$ on each piece [^5] Really, the rotation used in the construction should be rescaled since $[0,\delta]\times[0,1]$ and $[\delta,1]\times[0,1]$ are not squares., and concatenating to get a new homotopy class of paths $[\gamma^{\prime}]$. One checks that the class $[\gamma^{\prime}][\gamma]^{-1}$ in the braid group is given by $\sigma_{k-1}\cdots\sigma_{i}$, where there are $k$ points of the configuration, and where $y$ has the $i$th highest second coordinate of all elements in the configuration. We have illustrated checking this in Figure 3. Thus $g\circ\varphi_{-t}(w)$ is equivalent to the point where the path $[\gamma^{\prime}]$ is used instead of $[\gamma]$, and the element of $X^{k}$ is changed according to the action of $\sigma_{k-1}\cdots\sigma_{i}$. Letting $s$ approach $t$, we see that this is exactly the image of the point as described in relation $(ii)$.
We turn to proving the last statement of the proposition. Let $\overline{Q}^{\prime}$ denote a subspace of $\overline{Q}$ that is the image of finitely many components of $Q$ so that the image of $K$ factors through $\overline{Q}^{\prime}$. Each component of $Q$ is compact, and so $\overline{Q}^{\prime}$ is compact, since it is a quotient of finitely many such components. Since $B[M,\operatorname{Hur}^{X},N]$ is Hausdorff, and a continuous bijection from a compact space to a Hausdorff space is a homeomorphism, the inclusion of $\overline{Q}^{\prime}\to\overline{Q}$ is a homeomorphism onto its image. Thus, the unique lift of the map from $K$ to $\overline{Q}^{\prime}$ (and hence to $\overline{Q}$) is continuous. ∎
The following scanning argument is a ind-version of a quite standard scanning argument (see for example [RW20, Lemma 5.2] or [BDPW23, Proposition 5.2.7]), and is illustrated in Figure 2.
Figure 2. This is a picture of the scanning argument of § A.4. In the diagrams, there vertical distance at least $\varepsilon$ between any two red points, and all red points have second coordinate lie in $[\epsilon,1-\epsilon]$.
Figure 3. This is a picture depicting the compatibility of the map $\tilde{g}$ in the proof of § A.4 with relations $(i)$ and $(ii)$ from § A.4. The first two columns depict the paths $\gamma$ and $\gamma^{\prime}$ the points take during the map $\tilde{g}$ under relation $(i)$, and the composite $\gamma^{\prime}\gamma^{-1}$ is depicted underneath. The last two columns depict the paths $\gamma$ and $\gamma^{\prime}$ which the points take under relation $(ii)$, and the composite $\gamma^{\prime}\gamma^{-1}$ is again depicted underneath. In the diagrams, there are $4$ points, $p_{1}$ which is purple, $p_{2}$ which is red, $p_{3}$ which is green, and $p_{4}$ which is blue. So $r=2$ and $k=4$ in the notation of § A.4. The point $p_{2}$ is on the boundary. When $p_{2}$ is on the left boundary, the path is $\sigma_{1}\cdots\sigma_{r-1}=\sigma_{1}$ since $r=2$. When $p_{2}$ is on the right boundary, the path is $\sigma_{r}\cdots\sigma_{k-1}=\sigma_{2}\sigma_{3}$.
Lemma A.4.0.
The inclusions $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]\to B[M,\operatorname{Hur}^{X},N]$ of § A.4 indexed over the poset of real numbers $01>\epsilon>$ 0 are an ind-weak equivalence.
Proof.
It is easy to see that this map is always an equivalence on path components for each $\frac{1}{2}\geq\epsilon>0$, and that every element is in the component of a point given by $(m,n)\in M\times N$ and an empty configuration of points. It suffices to prove the statement for each connected path component, so we now fix a path component $Z$, restrict to $\epsilon\leq\frac{1}{2}$, and fix some $(m,n)$ on this component with empty configuration as the base point for all $\epsilon$. We will now apply § A.4 to the path component $Z$ of this base point. Consider a diagram
[…]
with $K,K^{\prime}$ compact. Choose a $\delta,\epsilon>0$ such that for all points in the image of $f:K\to B[M,\operatorname{Hur}^{X},N]$, the part of the configuration in $[\frac{1}{2}-\delta,\frac{1}{2}+\delta]\times[0,1]$ projects injectively under the second projection and lands in $[\epsilon,1-\epsilon]$, and each two points in the image of this projection have distance at least $\epsilon$ from each other. Since the flow in § A.4 sends $\frac{1}{2}\pm\delta$ to the boundary in a finite amount of time, we can use the flow to produce a continuous homotopy starting from $f$ and ending with a map landing in the image of $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$. First, we verify condition (2) from § A.4. Indeed, since $K$ is compact, at time $t=1$ of the homotopy, the map satisfies the criterion of the last sentence of § A.4, so factors through $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$. Finally, we verify condition $(1)$ of § A.4 by using the flow on $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$. Indeed, the homotopy also restricts on $K^{\prime}$ to a homotopy of pointed maps to $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$. This verifies the conditions of § A.4, as desired. ∎
We turn to proving a pointed version of § A.4. As a first step, we prove a lemma which shows that certain pushouts of our topological models are homotopy pushouts.
Lemma A.4.0.
Let $M$ be a right $\operatorname{Hur}^{X}$ module and $N$ be a left $\operatorname{Hur}^{X}$ module. Suppose that $Y\subset M\times N$ is a subset closed under the right and left actions of $\operatorname{Hur}^{X}$, and that $X^{\prime}\to X$ is an inclusion of braided sets. Let $A\subset\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ denote the subset of points in the image of points in $Q_{\epsilon}[M,\operatorname{Hur}^{X^{\prime}},N]$ whose projection to $M\times N$ lies in $Y$. Then $A\subset\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ has the homotopy extension property.
Proof.
It is easy to see that for each $01>\epsilon>$ 0, $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ admits the structure of a CW-complex of dimension $<\frac{2}{\epsilon}$. This is because each component of $Q_{\epsilon}[M,\operatorname{Hur}^{X},N]$ is a union of disks, and the gluing relations to form $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ are along subcomplexes. It is finite dimensional since $\epsilon$ puts a bound on the number of points in the configurations that are allowed. For a suitably chosen CW structure on $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$, the inclusion $A\subset\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ is that of a subcomplex, showing the lemma. ∎
We are finally prepared to prove a pointed version of § A.4, which is the main result of this section:
Theorem A.4.1.
Let $M$ be a pointed set with a right $\operatorname{Hur}^{X}$ action and $N$ be a pointed set with a left $\operatorname{Hur}^{X}$ action. There is a natural identification of $M\otimes_{\operatorname{Hur}^{X}_{+}}N$ with the ind-weak homotopy type of the quotient of the pointed spaces
$\overline{Q}^{*}_{\epsilon}[M,\operatorname{Hur}_{+}^{X},N]:=\overline{Q}_{\epsilon}[M,\operatorname{Hur}_{+}^{X},N]/S_{M,X,N},$
as $\epsilon$ approaches 00 (for $10<\epsilon<$ 1) where $S_{M,X,N}\subset\overline{Q}_{\epsilon}[M,\operatorname{Hur}_{+}^{X},N]$ is the subspace consisting of all points whose projection to either $M$ or $N$ is the base point.
Proof.
The pointed bar construction $M\otimes_{\operatorname{Hur}^{X}_{+}}N$ is the total cofiber of the square
(A.4) ${*\otimes_{\operatorname{Hur}^{X}}*}$${M\otimes_{\operatorname{Hur}^{X}}*}$${*\otimes_{\operatorname{Hur}^{X}}N}$${M\otimes_{\operatorname{Hur}^{X}}N.}$
In other words, $M\otimes_{\operatorname{Hur}^{X}_{+}}N$ is the cofiber of the map from the pushout of the diagram
$M\otimes_{\operatorname{Hur}^{X}}*\leftarrow*\otimes_{\operatorname{Hur}^{X}}*\rightarrow*\otimes_{\operatorname{Hur}^{X}}N$
to $M\otimes_{\operatorname{Hur}^{X}}N$.
By § A.3 and § A.4, $M^{\prime}\otimes_{\operatorname{Hur}^{X}}N^{\prime}$ can be modeled as the colimit over $\epsilon>0$ of the homotopy types of $Q_{\epsilon}[M^{\prime},\operatorname{Hur}^{X},N^{\prime}]$ for any unpointed right module $M^{\prime}$ and unpointed left module $N^{\prime}$. It then follows by applying § A.4, that the maps
$\overline{Q}_{\epsilon}[*,\operatorname{Hur}^{X},*]\to\overline{Q}_{\epsilon}[*,\operatorname{Hur}^{X},N]$
$\overline{Q}_{\epsilon}[*,\operatorname{Hur}^{X},*]\to\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},*]$
$\overline{Q}_{\epsilon}[*,\operatorname{Hur}^{X},N]\cup_{\overline{Q}_{\epsilon}[*,\operatorname{Hur}^{X},*]}\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},*]\to\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$
are cofibrations, i.e., they both have the homotopy extension property. Hence, the total cofiber of the square
(A.5) ${\overline{Q}_{\epsilon}[*,\operatorname{Hur}^{X},*]}$${\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},*]}$${\overline{Q}_{\epsilon}[*,\operatorname{Hur}^{X},N]}$${\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]}$
can be modeled as the quotient of $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ by $S_{M,X,N}$. The total cofiber of (A.4) is the colimit over $\epsilon$ of the total cofibers of the squares (A.5) for $\varepsilon>0$, since the total cofiber is a colimit, so commutes with the filtered colimit over $\varepsilon$. Thus we obtain the desired result. ∎